### v1

# Finite solvable groups of square-free Euler characteristic

Abstract.
A characterization is completed for finite groups which act regularly on maps with square-free Euler characteristic, associated with infinite families of square-free Euler characteristic regular maps presented. This is based on a classification of finite groups of which each Sylow subgroup has a cyclic or dihedral subgroup of prime index.
Keywords: solvable group, regular map, square-free, Euler characteristic
The project was partially supported by the NNSF of China (11931005).
November 18, 2025
1. Introduction
Edge-transitive maps are categorized into fourteen classes according to local structures and local actions in [9, 15]. The problem of constructing and classifying important classes of such maps with specific Euler characteristic has attracted considerable attention; refer to [4] - [7] for the Euler characteristic to be a negative prime or product of two primes. We aim towards to a classification of regular maps with square-free Euler characteristic.
Assume that $G$ is a finite group acting edge-transitively on a map with square-free Euler characteristic, which is said to be of square-free Euler characteristic for convenience. Then $G$ has the following restricted property (Lemma 2.1):
each Sylow subgroup has a cyclic or dihedral subgroup of prime index.
It leads us to studying groups of square-free Euler characteristic. This project extends the characterization of almost-Sylow cyclic groups, namely, each Sylow subgroup of odd order is cyclic, and the Sylow $2$-subgroup either is trivial or contains a cyclic subgroup of index 2; refer to Zassenhaus [18], Suzuki [16] and Wong [17].
In previous work [11], we characterized non-solvable groups of square-free Euler characteristic, and in [10], we determined $2$-groups of square-free Euler characteristic. As the third paper of this series, the purpose of this paper is to present a classification (characterization) of such groups in the solvable case.
As usual, for a group $G$ and a subset $\pi$ of prime divisors of $|G|$, denote by $G_{\pi}$ a Hall $\pi$-subgroup of $G$, which is a Sylow subgroup if $\pi$ contains a single prime $p$, and denoted by $G_{p}$. For a positive integer $n$, denote by ${\sf Z}_{n}$ a cyclic group of order $n$, by ${\rm D}_{2n}$ or ${\rm Q}_{2n}$ a dihedral or generalized quaternion group of order $2n$, and by ${\rm A}_{n}$ or ${\rm S}_{n}$ an alternating or symmetric group of degree $n$, respectively. A finite group $G$ is a semidirect product of $A$ by $B$ if $A\lhd G$ and $B<G$ are such that $G/A\cong B$, denoted by G=A:BG=A{:}B. Our first main result is the following theorem, which characterizes finite solvable groups of which each Sylow subgroup has a cyclic or dihedral subgroup of prime index.
Theorem 1.1.
Let $G$ be a finite solvable group such that each Sylow subgroup of $G$ has a cyclic or dihedral subgroup of prime index. Then:
(1) The Sylow subgroups of $G$ are characterized in Proposition 4.1 (for $G_{2}$) and Lemma 4.3 (for $G_{p}$, $p$ odd prime).
(2) G=(A:B):KG=(A{:}B){:}K, where $\gcd(|A{:}B|,|K|)=1$, such that
  • (i)
    AA is abelian, $B$ is nilpotent, $A\cap{\bf Z}(A{:}B)=1;$
  • (ii)
    $K$ is one of the following groups:
    • (a)
      $K=G_{2}$ ($K$ might be $1$ in case of $G$ of odd order);
    • (b)
      $K=G_{\{2,3\}}$, explicitly given in Lemma 5.6 (ii) - (iii);
    • (c)
      $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}G_{7}$, or $K=G_{\{2,3,7\}}={\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3})$.
The next theorem shows that automorphism groups of regular maps with square-free Euler characteristic are more restricted.
Theorem 1.2.
Let $\mathcal{M}$} be a regular map with square-free Euler characteristic, and let $G={\sf Aut}{\mathcal{M}}$ be solvable. Then G=(A:B):KG=(A{:}B){:}K, where $A$ is abelian, $B$ is nilpotent, $A\cap{\bf Z}(A{:}B)=1$, and $K=G_{2}$ or $G_{\{2,3\}}$ is one of the following groups:
  1. (i)
    $K={\sf Z}_{2}$, ${\rm D}_{2^{\ell+1}}$, ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$ or ${\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}$, where $\ell\geqslant 1;$
  2. (ii)
    $K=X$ or $X\times{\sf Z}_{2}$, where $X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\};$
  3. (iii)
    $K={\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$.
The paper is organized as follows. After the Introduction, we collect some preliminaries in Section 2, and construct several infinite families of regular maps with square-free Euler characteristic in Section 3. Then, we prove our two main results in Section 4 - 5 and Section 6, respectively.
2. Preliminaries
Let $G$ be a group. Denote by $\Phi(G)$ and $G^{\prime}$ the Frattini subgroup and the commutator subgroup of $G$. If $G$ is a $p$-group, then $^{\prime}\leqslant\Phi(G), G/G$$G^{\prime}$ is abelian, and $G/\Phi(G)$ is elementary abelian. For a subgroup $H\leqslant G$, denote by ${\bf N}_{G}(H)$ and ${\bf C}_{G}(H)$ the normalizer and the centralizer of $H$ in $G$, respectively.
Let $\it\Gamma$} be a graph and $\lambda$ be a positive integer. Denote by ${\it\Gamma}^{(\lambda)}$ the graph with each edge of $\it\Gamma$} replaced by $\lambda$ edges.
It is known (refer to [12]) that an arc-transitive (multi-) graph $\it\Gamma$} can be expressed as a coset graph :
${\it\Gamma}={\sf Cos}(G,H,J),\ \mbox{or}\ {\sf Cos}(G,H,HgH)\ \mbox{(for simple graphs)},$
which has vertex set $[G:H]$ and edge set $[G:J]$, such that
    $Hg_{1}$ is incident with $Jg_{0}$ if and only if $g_{1}g_{0}^{-1}\in HJ$; or equivalently,
    $Hg_{1}$, $Hg_{2}$ are adjacent if and only if $g_{2}g_{1}^{-1}\in HgH$ in the simple graph case.
A map $\mathcal{M}$} is a $2$-cell embedding of a graph into a closed surface. Let ${\mathcal{M}}=(V,E,F)$ be a map with vertex set $V$, edge set $E$ and face set $F$, and let $G={\sf Aut}{\mathcal{M}}$. The Euler characteristic of $\mathcal{M}$} is given by Euler formula:
$\chi({\mathcal{M}})=|V|-|E|+|F|.$
A flag of $\mathcal{M}$} is an incidence triple $(\alpha,e,f)$ with $\alpha\in V$, $e\in E$ and $f\in F$. Given a flag $(\alpha,e,f)$, it is well-known that each of the stabilizers $G_{\alpha}$ and $G_{f}$ is cyclic or dihedral, and meanwhile, $G_{e}=1$, ${\sf Z}_{2}$ or ${\rm D}_{4}$.
Further, if $\mathcal{M}$} is a flag-regular map (also called a regular map), i.e., $G$ acts regularly on flags, then there exists a triple $(x,y,z)$ of involutions of $G$ which determines the map $\mathcal{M}$} as follows:
$G_{\alpha}=\langle x,y\rangle$, $G_{e}=\langle x,z\rangle\cong{\rm D}_{4}$, $G_{f}=\langle y,z\rangle$.
Meanwhile, the Euler characteristic has the form
$\mbox{$\chi({\mathcal{M}})={|G|\over|G_{\alpha}|}-{|G|\over 4}+{|G|\over|G_{f}|}$}.$
Conversely, if a group $G$ has a triple of involutions $(x,y,z)$ such that
$G=\langle x,y,z\rangle$ and $xz=zx$ $(x\neq z)$,
then there exists a unique regular map ${\mathcal{M}}=(V,E,F)$ with
$V=[G:\langle x,y\rangle],\ E=[G:\langle x,z\rangle],\ F=[G:\langle y,z\rangle],$
and the incidence relation defined by non-empty intersection, such that $G={\sf Aut}{\mathcal{M}}$ acts regularly on flags. Hence, such a triple of involutions is called a regular triple for $G$, and the induced regular map is denoted by ${\mathcal{M}}(G,x,y,z)$. Moreover, the underlying graph of ${\mathcal{M}}(G,x,y,z)$ can be expressed as a coset graph:
${\sf Cos}(G,\langle x,y\rangle,\langle x,z\rangle)$, or ${\sf Cos}(G,H,HzH)$ where $H=\langle x,y\rangle$ (for simple graphs).
In conclusion, each of the following two things determines the other:
    the regular maps arisen from a group $G$;
    the regular triples for $G$.
The following lemma established in [11] characterizes Sylow subgroups of $G={\sf Aut}({\mathcal{M}})$ in the case where $\mathcal{M}$} is with square-free characteristic. We note that if $\chi({\mathcal{M}})$ is square-free, then of course so is $\gcd(\chi({\mathcal{M}}),|{\sf Aut}{\mathcal{M}}|)$.
Lemma 2.1.
([11, Lem. 2.2]) Let $\mathcal{M}$} be a map, and let $G={\sf Aut}{\mathcal{M}}$. Assume that $\gcd(\chi({\mathcal{M}}),|G|)$ is square-free. Then each Sylow $p$-subgroup of $G$ has a cyclic or dihedral subgroup of index $p$.
We thus study finite groups satisfying the following hypothesis.
Hypothesis 2.2.
Let $G$ be a finite group such that each Sylow subgroup of $G$ has a cyclic or dihedral subgroup of prime index.
The next lemma plays a key role in the analysis of groups satisfying Hypothesis 2.2.
Lemma 2.3.
([11, Lem. 2.5]) Let a finite group $G$ satisfy Hypothesis 2.2. Then so does each subgroup and each factor group of $G$.
3. Examples of square-free Euler characteristic regular maps
We construct several infinite families of regular maps with square-free Euler characteristic in this section. The constructions are based on the following groups:
$\begin{array}[]{lllll}X&=&(\langle a,s\rangle\times\langle b,t\rangle){:}\langle\sigma\rangle\cong{\rm D}_{2n}\wr{\rm S}_{2}&&\\ &=&\langle a,b\rangle{:}\langle s,\sigma\rangle\cong{\sf Z}_{n}^{2}{:}{\rm D}_{8},&&(\ast)\end{array}$
where $\langle a,s\rangle=\langle a\rangle{:}\langle s\rangle={\rm D}_{2n}$, $\langle b,t\rangle=\langle b\rangle{:}\langle t\rangle={\rm D}_{2n}$, $(a,s)^{\sigma}=(b,t)$ and $|\sigma|=2$.
The first family is given now.
Construction 3.1.
Let $G=\langle a,s\rangle\times\langle b,t\rangle={\rm D}_{2n}\times{\rm D}_{2n}\leqslant X$ as given in ($\ast$), with $n$ odd. Let
x=s,y=abst,z=st.x=s,\ y=abst,\ z=st.
Then $(x,y,z)$ is a triple of involutions with $xz=zx$. Further,
$(yx)^{2}=(abt)^{2}=a^{2}$ and $(yxz)^{2}=(abs)^{2}=b^{2}$.
As $n$ is odd, $\langle a\rangle=\langle a^{2}\rangle$ and $\langle b\rangle=\langle b^{2}\rangle$, so that $G=\langle x,y,z\rangle$. Thus, $(x,y,z)$ is a regular triple for $G$.
Lemma 3.2.
Let ${\mathcal{M}}_{n}={\mathcal{M}}(G,x,y,z)$ be a regular map defined in Construction 3.1. Then the underlying graph of ${\mathcal{M}}_{n}$ is a multiple cycle ${\bf C}_{n}^{(n)}$, and $\chi({\mathcal{M}}_{n})=-n(n-3)$. In particular, $\chi({\mathcal{M}}_{n})$ is square-free for infinitely many odd values of $n$.
Proof.
The vertex stabilizer $H$ is
$H=\langle x,y\rangle=\langle s,abst\rangle=\langle abt\rangle{:}\langle s\rangle=\langle a,s\rangle\times\langle bt\rangle={\rm D}_{4n}.$
Meanwhile, the face stabilizer is
$\langle y,z\rangle=\langle abst,st\rangle=\langle ab\rangle{:}\langle st\rangle={\rm D}_{2n}.$
Thus the map ${\mathcal{M}}_{n}$ has valency $2n$ and face length $n$. Further, $H\cap H^{z}=\langle a,s\rangle\cong{\rm D}_{2n}$ is of index $2$ in $H$, so the underlying graph ${\it\Gamma}={\sf Cos}(G,H,HzH)\cong{\bf C}_{n}^{(n)}$.
The Euler characteristic
$\chi({\mathcal{M}}_{n})=|G|\left({1\over|\langle x,y\rangle|}-{1\over 4}+{1\over|\langle y,z\rangle|}\right)=4n^{2}\left({1\over 4n}-{1\over 4}+{1\over 2n}\right)=-n(n-3).$
Let $n=4x+1$. Then $\chi({\mathcal{M}}_{n})=-2(4x+1)(2x-1)$. By a result given by G. Ricci in 1933 (which is improved by $\rm P.Erd\ddot{o}s$} in [8]), there are infinitely many positive integers $x$ for which f(x)=(4x+1)(2x1)f(x)=(4x+1)(2x-1) is square-free, so there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. $\Box$
Here is some data such that n(n3)-n(n-3) is square-free, with $n$ odd.
$\begin{array}[]{c|ccccccc}\hline\cr n&5&13&17&29&37&41&\cdots\\ -n(n-3)&-2.5&-2.5.13&-2.7.17&-2.13.29&-2.17.37&-2.19.41$$\cdots\\ \hline\cr\end{array}$
Next we construct the second family.
Construction 3.3.
Let $G=\langle a,b\rangle{:}\langle st,\sigma\rangle={\sf Z}_{n}^{2}{:}{\sf Z}_{2}^{2}\leqslant X$ as given in ($\ast$). Let
$x=\sigma,\ y=ast,\ z=abst.$
Then $(x,y,z)$ is a triple of involutions with $xz=zx$. Further,
$yz=b^{-1}$ and $(yz)^{x}=a^{-1}$,
so that $G=\langle x,y,z\rangle$. Thus, $(x,y,z)$ is a regular triple for $G$.
Lemma 3.4.
Let ${\mathcal{M}}_{n}={\mathcal{M}}(G,x,y,z)$ be a regular map defined in Construction 3.3. Then the underlying graph of ${\mathcal{M}}_{n}$ is ${\bf C}_{n}^{(n)}$, and $\chi({\mathcal{M}}_{n})=-n(n-3)$. In particular, $\chi({\mathcal{M}}_{n})$ is square-free for infinitely many values of $n$.
Proof.
The vertex stabilizer $H$ is
$H=\langle x,y\rangle=\langle ast\sigma\rangle{:}\langle\sigma\rangle={\rm D}_{4n}.$
Meanwhile, the face stabilizer is
$\langle y,z\rangle=\langle abst,ast\rangle=\langle b\rangle{:}\langle ast\rangle={\rm D}_{2n}.$
Thus the map ${\mathcal{M}}_{n}$ has valency $2n$ and face length $n$. Further, $H\cap H^{z}=\langle a^{-1}b\rangle{:}\langle\sigma\rangle={\rm D}_{2n}$ is of index $2$ in $H$, so the underlying graph ${\it\Gamma}={\sf Cos}(G,H,HzH)\cong{\bf C}_{n}^{(n)}$.
The Euler characteristic
$\chi({\mathcal{M}})=|G|\left({1\over|\langle x,y\rangle|}-{1\over 4}+{1\over|\langle y,z\rangle|}\right)=4n^{2}\left({1\over 4n}-{1\over 4}+{1\over 2n}\right)=-n(n-3).$
As Lemma 3.2 shows, there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. Further, let $n=4x+2$. Then $\chi({\mathcal{M}}_{n})=-2(2x+1)(4x-1)$. According to [8], there are infinitely many positive integers $x$ for which f(x)=(2x+1)(4x1)f(x)=(2x+1)(4x-1) is square-free, so there are infinitely many even values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. $\Box$
Here is some data such that n(n3)-n(n-3) is square-free, with $n$ even.
$\begin{array}[]{c|ccccccc}\hline\cr n&2&10&14&22&26&34&\cdots\\ -n(n-3)&2&-2.5.7&-2.7.11&-2.11.19&-2.13.23&-2.17.31$$\cdots\\ \hline\cr\end{array}$
The third family is given below.
Construction 3.5.
Let $G=\langle a,b\rangle{:}\langle s,\sigma\rangle={\sf Z}_{n}^{2}{:}{\rm D}_{8}=X$ as given in ($\ast$), with $n$ odd. Let
$x=\sigma,\ y=s,\ z=abst.$
Then $(x,y,z)$ is a triple of involutions with $xz=zx$. Further,
$(zy)^{2}=(abt)^{2}=a^{2}$ and $((zy)^{2})^{x}=(a^{2})^{x}=b^{2}$.
As $n$ is odd, $\langle a\rangle=\langle a^{2}\rangle$ and $\langle b\rangle=\langle b^{2}\rangle$, so that $G=\langle x,y,z\rangle$. Thus, $(x,y,z)$ is a regular triple for $G$.
Lemma 3.6.
Let ${\mathcal{M}}_{n}={\mathcal{M}}(G,x,y,z)$ be a regular map defined in Construction 3.5. Then the underlying graph of ${\mathcal{M}}_{n}$ is ${\bf C}_{n}\square{\bf C}_{n}$, and $\chi({\mathcal{M}}_{n})=-n(n-2)$. In particular, $\chi({\mathcal{M}}_{n})$ is square-free for infinitely many prime values of $n$.
Proof.
The vertex stabilizer $H$ is
$H=\langle x,y\rangle=\langle\sigma,s\rangle={\rm D}_{8}.$
Meanwhile, the face stabilizer
$\langle y,z\rangle=\langle s,abst\rangle=\langle abt\rangle{:}\langle s\rangle={\rm D}_{4n}.$
Thus the map ${\mathcal{M}}_{n}$ has valency $4$ and face length $2n$. Let ${\it\Gamma}={\sf Cos}(G,H,HzH)$ be the underlying graph of ${\mathcal{M}}_{n}$. Then $G=\langle a,b\rangle{:}H$, and so $\langle a,b\rangle$ is regular on the vertex set of $\it\Gamma$}. Hence $\it\Gamma$} is a (connected and undirected) Cayley graph ${\sf Cay}(K,S)$ of valency 4, where $K=\langle a,b\rangle={\sf Z}_{n}^{2}$. It follows that $S=\{x,x^{-1},y,y^{-1}\}$ with $\langle x,y\rangle=\langle a,b\rangle$, and so ${\it\Gamma}={\bf C}_{n}\square{\bf C}_{n}$.
The Euler characteristic
$\chi({\mathcal{M}}_{n})=|G|\left({1\over|\langle x,y\rangle|}-{1\over 4}+{1\over|\langle y,z\rangle|}\right)=8n^{2}\left({1\over 8}-{1\over 4}+{1\over 4n}\right)={-n(n-2).}$
According to [8], there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free. Further, by a theorem of Chen [3], there indeed exist infinitely many primes $p$ such that p2p-2 is square-free, then so is p(p2)-p(p-2). $\Box$
Here is some data such that p(p2)-p(p-2) is square-free, with $p$ prime.
$\begin{array}[]{c|cccccccc}\hline\cr p&3&5&7&13&17&19&23&\cdots\\ -p(p-2)&-3&-3.5&-5.7&-11.13&-3.5.17&-17.19&-3.7.23$$\cdots\\ \hline\cr\end{array}$
4. $p$-Groups
In this section, we study $p$-groups satisfying Hypothesis 2.2, i.e., $p$-groups with a cyclic or dihedral subgroup of index $p$. Note that the case for $p=2$ is dealt with in the previous work [10], which is introduced below.
Let $A,B$ be finite groups whose centers ${\bf Z}(A),{\bf Z}(B)$ have isomorphic subgroups $C_{1},C_{2}$, respectively. Let $\varphi$ be an isomorphism from $C_{1}$ to $C_{2}$, and let $C=\{(c,c^{\varphi})\mid c\in C_{1}\}$. Then $C$ is a normal subgroup of $A\times B$. The factor group $(A\times B)/C$ is called a central product of $A$ and $B$, denoted by $A\circ_{C}B$, and sometimes simply by $A\circ B$ if $C$ is equal to ${\bf Z}(A)$ or ${\bf Z}(B)$.
Let ${\rm Q}_{2^{\ell+1}}=\langle u,v\rangle$ be a generalized quaternion group of order $2^{\ell+1}$, where
$|u|=2^{\ell}\geqslant 2^{2}$, $u^{2^{\ell-1}}=v^{2}$, and $u^{v}=u^{-1}$.
Then ${\rm Q}_{2^{\ell+1}}$ has $u^{2^{\ell-1}}=v^{2}$ as the only involution. Let
$G={\rm Q}_{2^{\ell+1}}\circ{\sf Z}_{4}=\langle u,v\rangle\circ\langle b\rangle=(\langle u,v\rangle\times\langle b\rangle)/\langle b^{2}v^{2}\rangle.$
Then $G$ contains normal subgroups $\langle u,v\rangle={\rm Q}_{2^{\ell+1}}$ and $\langle b\rangle={\sf Z}_{4}$, and the only involutions contained in $\langle u,v\rangle$ and $\langle b\rangle$ are $u^{2^{\ell-1}}=v^{2}$ and $b^{2}$, respectively, which are actually identical in $G$ as $b^{2}v^{2}=1$.
Proposition 4.1.
([10], Thm. 1.1) Let $G$ be a $2$-group which has a cyclic or dihedral subgroup of index $2$. Then $G$ is one of the following groups. Further, either ${\sf Aut}(G)$ is a $2$-group, or $(G,{\sf Aut}(G))$ is one of $({\sf Z}_{2}^{2},{\rm S}_{3}),\,({\sf Z}_{2}^{3},{\rm GL}_{3}(2)),\,({\rm Q}_{8},{\rm S}_{4}),\,({\rm Q}_{8}\circ{\sf Z}_{4},{\rm S}_{4}\times{\sf Z}_{2})$.
  • (i)
    $G$ is one of the following groups, with a cyclic subgroup of index $2$, $\ell\geqslant 1$:
    • (i.1)
      $G={\sf Z}_{2^{\ell}}$ or ${\sf Z}_{2^{\ell}}\times{\sf Z}_{2};$
    • (i.2)
      $G={\rm D}_{2^{\ell+2}}$ or ${\rm Q}_{2^{\ell+2}};$
    • (i.3)
      $G={\sf Z}_{2^{\ell+2}}{:}{\sf Z}_{2}=\langle a\rangle{:}\langle b\rangle$ with $a^{b}=a^{2^{\ell+1}\pm 1};$
  • (ii)
    $G$ is one of the following groups, with a dihedral subgroup of index $2$, $\ell\geqslant 1$:
    • (ii.1)
      $G={\rm D}_{2^{\ell+1}}\times{\sf Z}_{2};$
    • (ii.2)
      $G={\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}=\langle a,b\rangle{:}\langle c\rangle$, where $\langle a\rangle{:}\langle b\rangle={\rm D}_{2^{\ell+3}}$, $(a,b)^{c}=(a^{2^{\ell+1}+1},b);$
    • (ii.3)
      $G={\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4}$.
Proposition 4.2.
([10], Cor. 4.9) Let $G$ be a $2$-group which has a cyclic or dihedral subgroup of index $2$. If $G$ has a regular triple, then it is one of the following groups:
  • (i)
    $G={\rm D}_{2^{\ell+1}}$ or ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$, $\ell\geqslant 1;$
  • (ii)
    $G={\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}=\langle a,b\rangle{:}\langle c\rangle$, where $\langle a\rangle{:}\langle b\rangle={\rm D}_{2^{\ell+3}}$, $(a,b)^{c}=(a^{2^{\ell+1}+1},b)$, $\ell\geqslant 1$.
For the odd primes $p$, we have the following characterization.
Lemma 4.3.
Let $p>2$ be an odd prime, and let $G$ be a $p$-group which has a cyclic subgroup of index $p$. Then one of the following holds. In particular, each prime divisor of $|{\sf Aut}(G)|$ is less than or equal to $p$.
  • (i)
    $G={\sf Z}_{p^{\ell}}$ is cyclic, and ${\sf Aut}(G)={\sf Z}_{(p-1)p^{\ell-1}}$.
  • (ii)
    $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$ is abelian, and
    1. (ii.1)
      if $\ell>1$, then ${\sf Aut}(G)$ is solvable, and letting $G=\langle a,b\rangle$, then any Hall $p^{\prime}$-subgroup of ${\sf Aut}(G)$ is conjugate to $({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{p^{\prime}}={\sf Z}_{p-1}\times{\sf Z}_{p-1};$
    2. (ii.2)
      if $\ell=1$, then ${\sf Aut}(G)={\rm GL}(2,p)$, and further, any subgroup $M\leqslant{\rm GL}(2,p)$ of order coprime to $2p$ is an abelian group.
  • (iii)
    $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ is non-abelian with $a^{b}=a^{p^{\ell-1}+1}$, $\ell\geqslant 2$, and ${\sf Aut}(G)$ is solvable such that any Hall $p^{\prime}$-subgroups of ${\sf Aut}(G)$ is conjugate to $\langle\rho\rangle={\sf Z}_{p-1}$, where $a^{\rho}=a^{\xi}\not=a$ with $\xi\,\,\big|\,\,p-1$, and $b^{\rho}=b$.
Proof.
According to [13, 5.3.4, pp.136], $G$ is one of the groups stated in this lemma. We only need to verify the statements regarding the automorphism groups of $G$.
(1) The automorphism groups of cyclic groups are well-known; refer to [14, Thm. 7.3].
(2) Suppose $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$ with $\ell>1$. Then by [1, Thm. 3.2], ${\sf Aut}(G)$ is of order
$\begin{array}[]{ccl}|{\sf Aut}(G)|&=&|{\sf Aut}({\sf Z}_{p^{\ell}})|\cdot|{\sf Aut}({\sf Z}_{p})|\cdot|{\sf Hom}({\sf Z}_{p^{\ell}},{\sf Z}_{p})|\cdot|{\sf Hom}({\sf Z}_{p},{\sf Z}_{p^{\ell}})|\\ &=&(p-1)p^{\ell-1}\cdot(p-1)\cdot p\cdot p=p^{\ell+1}(p-1)^{2}.\end{array}$
More explicitly, let $G=\langle a\rangle\times\langle b\rangle$ with $|a|=p^{\ell}$, $|b|=p$. Then $\Phi(G)=\langle a^{p}\rangle$. Let $\overline{G}=G/\Phi(G)$, and consider the homomorphism
$\sigma:{\sf Aut}(G)\rightarrow{\sf Aut}(\overline{G})={\rm GL}(2,p).$
On the one hand, the kernel of $\sigma$ consists of automorphisms $g\in{\sf Aut}(G)$ which stabilizes each of the $\Phi(G)$-cosets, and in particular, stabilizes $\Phi(G)\langle a\rangle$ and $\Phi(G)\langle b\rangle$. Thus, by direct calculation we have:
$a^{g}\in\{a,a^{p+1},a^{2p+1},\cdots,a^{(p^{\ell-1}-1)p+1}\}\mbox{\ and\ }b^{g}\in\{b,a^{p^{\ell-1}}b,a^{2p^{\ell-1}}b,\cdots,a^{(p-1)p^{\ell-1}}b\}.$
Consequently, ${\rm Ker}(\sigma)=\langle x\rangle\times\langle y\rangle\cong{\sf Z}_{p^{\ell-1}}\times{\sf Z}_{p},$ where
$x:a\rightarrow a^{p+1},b\rightarrow b\ \mbox{\,and\,}\ y:a\rightarrow a,b\rightarrow a^{p^{\ell-1}}b.$
On the other hand, the image of $\sigma$ falls into the (solvable) parabolic subgroup $P$ of ${\sf Aut}(\overline{G})={\rm GL}(2,p)$ which stabilizes $\langle\overline{b}\rangle$, because
$\langle\overline{b}\rangle=\{\overline{1},\overline{b},\cdots,\overline{b}^{p-1}\}=\{\Phi(G),b\Phi(G),\cdots,b^{p-1}\Phi(G)\}$
exactly consists of those elements of $G$ of order less than $p^{\ell}$. Moreover, it follows from ${\sf Aut}(G)/{\rm Ker}(\sigma)\cong{\rm Im}(\sigma)$ that $|{\rm Im}(\sigma)|=p(p-1)^{2}=|P|$, and so ${\rm Im}(\sigma)=P$.
Then ${\sf Aut}(G)$ is solvable, since both ${\rm Ker}(\sigma)$ and ${\rm Im}(\sigma)$ are solvable. Note that the following subgroup
$X=({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{p^{\prime}}={\sf Z}_{p-1}\times{\sf Z}_{p-1}$
is a Hall $p^{\prime}$-subgroup of $G$. Then any other Hall $p^{\prime}$-subgroup is conjugate to $X$.
(3) Suppose $G={\sf Z}_{p}\times{\sf Z}_{p}$. Then ${\sf Aut}(G)={\rm GL}(2,p)$, of order $p(p+1)(p-1)^{2}$ with prime divisors less than or equal to $p$. Further, let $M\leqslant{\rm GL}(2,p)$ be a subgroup of order coprime to $2p$, and let $N=M/{\bf Z}(M)$. Then
$N=M/{\bf Z}(M)\leqslant M/({\bf Z}({\rm GL}(2,p))\cap M)\lesssim{\rm PGL}(2,p).$
Further, since ${\rm PSL}(2,p)N\leqslant{\rm PGL}(2,p)$, we have
$|N/({\rm PSL}(2,p)\cap N)|\leqslant|{\rm PGL}(2,p)|/|{\rm PSL}(2,p)|=2.$
It then follows by $2\nmid|N|$ that $N$ is indeed embedded into ${\rm PSL}(2,p)$. Note that the subgroups of ${\rm PSL}(2,p)$ are all explicitly known, by check which a subgroup of order coprime to $2p$ must be a cyclic group. Then $M$ is an abelian group, since $N=M/{\bf Z}(M)$ is cyclic.
(4) Suppose that $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ with $a^{b}=a^{p^{\ell-1}+1}$ is non-abelian, $\ell\geqslant 2$. The automorphism group ${\sf Aut}(G)$ is described in [2, Thm. 3.1 & 3.2], by which we have: $|{\sf Aut}(G)|=(p-1)p^{\ell+1}$, ${\sf Aut}(G)$ is solvable with the Sylow $p$-subgroup normal, and $X=\langle\rho\rangle={\sf Z}_{p-1}$ is a Hall $p^{\prime}$-subgroup of $G$, where $\rho$ stabilizes $\langle a\rangle$ and centralizes $\langle b\rangle$, that is, $a^{\rho}=a^{\xi}\not=a$ with $\xi\,\,\big|\,\,p-1$, and $b^{\rho}=b$. $\Box$
5. Proof of Theorem 1.1
In this section, we study the way how the $p$-groups described in Section 4 compose a larger solvable group, and give a proof of Theorem 1.1
Let $G$ be a finite group, with $\{p_{1},\ldots,p_{s}\}$ the prime divisors of $|G|$. For $p=p_{i}$, denote by ${\bf O}_{p}(G)$ the maximal normal $p$-subgroup of $G$, or equivalently, the intersection of all Sylow $p$-subgroups of $G$. Let $F={\rm Fit}(G)$ be the Fitting subgroup of $G$, which is the largest nilpotent normal subgroup. Then
$F=F_{p_{1}}\times\cdots\times F_{p_{s}},\mbox{\ with\ each\ }F_{p_{i}}={\bf O}_{p_{i}}(G).$
Note that both ${\bf O}_{p}(G)$ and $F$ are characteristic in $G$. Further, if $G$ is solvable, then ${\bf C}_{G}(F)\leqslant F$, and so
$G/F={\bf N}_{G}(F)/\left(F{\bf C}_{G}(F)\right)\lesssim{\sf Out}(F).$
We first consider solvable groups of odd order.
Lemma 5.1.
Let $G$ be a solvable group of odd order satisfying Hypothesis 2.2. Let $p$ be the largest prime divisor of $|G|$. Then a Sylow $p$-subgroup $G_{p}$ is normal in $G$, and so $G=G_{p}{:}G_{p^{\prime}}$.
Proof.
Let $\overline{G}=G/{\bf O}_{p}(G)$, and let $F={\rm Fit}(\overline{G})$. Notice that ${\bf O}_{p}(\overline{G})=1$. We have $F=F_{p_{1}}\times\cdots\times F_{p_{s}}$ with all $p_{i}<p$. Now,
$\overline{G}/F\lesssim{\sf Out}(F)={\sf Out}(F_{p_{1}})\times\cdots\times{\sf Out}(F_{p_{s}}),$
and so $p\nmid|\overline{G}/F|$ by Lemma 4.3. Thus, $p\nmid|\overline{G}|$, so that, $G_{p}={\bf O}_{p}(G)\lhd G$. $\Box$
Lemma 5.2.
Let $G$ be a $\{p,q\}$-group satisfying Hypotheses 2.2, where $p>q$ are both odd primes, and $q\nmid p^{2}-1$. Then $G=G_{p}\times G_{q}$.
Proof.
It is proved in a similar way as the previous lemma. Note that by Lemma 4.3 any prime divisor of $|{\sf Aut}(G_{p})|$ divides p(p1)(p+1)p(p-1)(p+1). $\Box$
Before giving a general characterization of solvable groups of odd order, we need the following technical lemma.
Lemma 5.3.
Let $H={\sf Z}_{p^{\ell}}$ be a cyclic group and $x$ be an automorphism of $H$ with order coprime to $p$. Let $X=H{:}\langle x\rangle$. Then, (i) $H\leqslant X^{\prime};$ (ii) $H\cap{\bf Z}(X)=1$.
Proof.
By Lemma 4.3 (i), the order of $x$ divides p1p-1. Let $H=\langle a\rangle$ and suppose
$a^{x}=a^{m}$, where $1<m\leqslant p^{\ell}-1$,  $p\nmid m$  and  $p^{\ell}\mid m^{p-1}-1$.
If $p\mid m-1$, then by LTE (Lift the Exponent Lemma) we have $p^{\ell}\mid m-1$, a contradiction. Hence $p\nmid m-1$. Thus, (i) the commutator $a,x]=a^{-1}a^{x}=a^{m-1}$ can generate $H$, so that, $H\leqslant X^{\prime}$; (ii) for any $1\leqslant k<p^{\ell}$, $(a^{k})^{x}=a^{km}\neq a^{k}$, so that, $H\cap{\bf Z}(X)=1$. $\Box$
Lemma 5.4.
Let $G$ be a solvable group of odd order satisfying Hypothesis 2.2. Then G=C:DG=C{:}D, where $C$ is abelian, $D$ is nilpotent, and $C\cap{\bf Z}(G)=1$.
Proof.
Let $n$ be the number of prime divisors of $|G|$. The proof proceeds by induction on $n$. For $n=1$, the lemma holds with $C=1$, $D=G$. Suppose $n\geqslant 2$, and assume inductively that the lemma holds for groups of order with less prime divisors.
Let $p$ be the largest prime divisor of $|G|$. By Lemma 5.1 and the induction assumption, we have
$G=G_{p}{:}G_{p^{\prime}}=G_{p}{:}(A{:}B),$
with $A$ abelian, $B$ nilpotent, and $A\cap{\bf Z}(A{:}B)=1$, so that, $B$ does not centralize any Sylow subgroup of $A$. Further, if $G=G_{p}\times G_{p^{\prime}}$, then
$G=G_{p}{\times}(A{:}B)=(G_{p}{\times}A){:}B=A{:}(G_{p}\times B),$
which gives an expected factorization of $G$. We thus assume that $G\not=G_{p}\times G_{p^{\prime}}$, namely, $G_{p^{\prime}}$ acts by conjugation on $G_{p}$ non-trivially. We shall confirm
Claim: $G_{p}$ is centralized by each Sylow $r$-subgroup $A_{r}$ of $A$, so that, $G_{p}$ is centralized by $A$.
By Lemma 4.3, we have two cases to treat: $A_{r}={\sf Z}_{r^{k}}$ or ${\sf Z}_{r^{k}}\times{\sf Z}_{r}$.
Case 1. Let $A_{r}={\sf Z}_{r^{k}}$. If $A_{r},B_{r^{\prime}}]=1$, then by Lemma 4.3, $(A{:}B)_{r}=A_{r}{:}B_{r}={\sf Z}_{r^{k}}{:}{\sf Z}_{r}$ is non-abelian, and
$1\not=A_{r}\cap{\bf Z}(A_{r}{:}B_{r})=A_{r}\cap{\bf Z}(A{:}B)\leqslant A\cap{\bf Z}(A{:}B),$
a contradiction. Thus there exists a Sylow $q$-subgroup $B_{q}$ with $q\not=r$ which does not centralize $A_{r}$. Take $x\in B_{q}$ to be an element acting non-trivially on $A_{r}$ and let $X=A_{r}{:}\langle x\rangle$. Then $X$ acts on $G_{p}$ and induces an abelian automorphism group of $G_{p}$ by Lemma 4.3. It then follows by Lemma 5.3 (i) that $A_{r}\leqslant X^{\prime}\leqslant{\bf C}_{X}(G_{p})$.
Case 2. Let $A_{r}={\sf Z}_{r^{k}}\times{\sf Z}_{r}$, where $k\geqslant 1$. Then $r$ does not divide $|B|$. By Lemma 4.3 (ii), it implies that either
  • (i)
    $A_{r}={\sf Z}_{r}^{2}$ and $B$ is irreducible on $A_{r}$, or
  • (ii)
    there exists $a,b\in A_{r}$ such that $A_{r}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$ and $\langle b\rangle$ are normalized by $B$.
For the former, $A_{r}$ centralizes $G_{p}$ since $A_{r}{:}B$ induces an abelian automorphism group of $G_{p}$ by Lemma 4.3. For the latter, there exist elements $u,v\in B$ such that $u$ acts on $\langle a\rangle$ non-trivially, and $v$ acts on $\langle b\rangle$ non-trivially. Then similar to Case 1 we have $G_{p},a]=[G_{p},b]=1$, so that, $G_{p},A_{r}]=[G_{p},\langle a,b\rangle]=1$.
We have therefore confirmed the claim. Consequently we have
$G=G_{p}{:}(A{:}B)=(G_{p}\times A){:}B.$
Meanwhile, by Lemma 4.3, $G_{p}$ is one of the following groups:
$G_{p}={\sf Z}_{p^{\ell}}$,  ${\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$  or  ${\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$.
Note that by assumption, $G_{p^{\prime}}=A{:}B$ acts by conjugation on $G_{p}$ non-trivially, which implies that $B$ acts on $G_{p}$ non-trivially.
Suppose $G_{p}={\sf Z}_{p^{\ell}}$. As $B$ acts on $G_{p}$ non-trivially, it follows from Lemma 5.3 (ii) that $G_{p}\cap{\bf Z}(G)=1$. Thus G=C:DG=C{:}D satisfies the theorem, with $C=G_{p}\times A$, $D=B$.
Suppose $G_{p}={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}=\langle a\rangle{:}\langle b\rangle$. By Lemma 4.3 (iii), $B$ acts on $\langle a\rangle$ non-trivially and acts on $\langle b\rangle$ trivially. Rewrite $G$ as
$G=(A\times(\langle a\rangle{:}\langle b\rangle)){:}B=(A\times\langle a\rangle){:}(B\times\langle b\rangle).$
Then by Lemma 5.3 (ii), G=C:DG=C{:}D satisfies the theorem, with $C=A\times\langle a\rangle$, $D=B\times\langle b\rangle$.
Suppose $G_{p}={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$, where $\ell\geqslant 1$. By Lemma 4.3 (ii), it implies that either (i) $G_{p}={\sf Z}_{p}^{2}$ and $B$ is irreducible on $G_{p}$; or (ii) there exists $a,b\in G_{p}$ such that $G_{p}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$ and $\langle b\rangle$ are normalized by $B$. For the former, G=C:DG=C{:}D satisfies the theorem, with $C=G_{p}\times A$, $D=B$. For the former, if $B$ acts non-trivially on both $\langle a\rangle$ and $\langle b\rangle$, then also let $C=G_{p}\times A$, $D=B$. Otherwise, $B$ acting non-trivially on $G_{p}$ implies $B$ acting non-trivially on some one of $\langle a\rangle$ and $\langle b\rangle$, say $\langle a\rangle$. Then similar to the last paragraph, let $C=A\times\langle a\rangle$, $D=B\times\langle b\rangle$. $\Box$
We next deal with solvable groups of even order.
Lemma 5.5.
Let $G$ be a solvable group satisfying Hypothesis 2.2. If $F:={\rm Fit}(G)={\bf O}_{2}(G)$, then either $G$ is a $2$-group, or $G$ is one of the following:
  • (i)
    $G={\sf Z}_{2}^{3}{:}{\sf Z}_{7}$ or ${\sf Z}_{2}^{3}{:}{\sf Z}_{7}{:}{\sf Z}_{3};$
  • (ii)
    $G={\sf Z}_{2}^{2}{:}H$, ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}H)$, ${\rm Q}_{8}{:}H$, or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}H)$, where $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$, ${\sf Z}_{2}^{2}{:}H\cong{\rm A}_{4}$ or ${\rm S}_{4}$, and ${\rm Q}_{8}{:}H\cong{\rm SL}(2,3)$ or ${\rm GL}(2,3)$.
Proof.
Assume that $G$ is not a 2-group. Then $G/F\leqslant{\sf Out}(F)$ is not a 2-group either. Thus by Lemma 4.1, $F$ is one of the following 2-groups:
$F\in\{{\sf Z}_{2}^{2},{\sf Z}_{2}^{3},{\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8}\}.$
Moreover, ${\bf O}_{2}(G/F)$ is trivial, namely, $G/F$ has no normal 2-subgroups. Then by checking ${\sf Out}(F)$, either $|G/F|$ is divisible by 7, and $F={\sf Z}_{2}^{3}$, $G/F={\sf Z}_{7}$ or ${\sf Z}_{7}{:}{\sf Z}_{3}$; or $G/F={\sf Z}_{3}$ or ${\rm S}_{3}$. It follows that $G$ is one of the groups listed in the lemma. In particular, if $G=F{:}{\rm S}_{3}$, then $G_{2}=F{:}{\sf Z}_{2}>F$, and respectively for $F={\sf Z}_{2}^{2}$, ${\sf Z}_{2}^{3}$, ${\rm Q}_{8}$ or ${\sf Z}_{4}\circ{\rm Q}_{8}$, it is easy to verify that $G_{2}$ has a subgroup $K$ of index $2$ with $K={\sf Z}_{4}$, ${\rm D}_{8}$, ${\sf Z}_{8}$ or ${\rm D}_{16}$. That is, $G=F{:}{\rm S}_{3}$ surely satisfies Hypothesis 2.2. $\Box$
Lemma 5.6.
Let $K$ be a $\{2,p\}$-group with $p$ odd prime satisfying Hypotheses 2.2. If $K\neq K_{p}{:}K_{2}$, then one of the following holds.
  1. (i)
    $p=7$, $K={\sf Z}_{2}^{3}{:}K_{7};$
  2. (ii)
    $p=3$, $K=X{:}K_{3}$ with $X\in\{{\sf Z}_{2}^{2},\,{\sf Z}_{2}^{3},\,{\rm Q}_{8},\,{\sf Z}_{4}\circ{\rm Q}_{8}\};$
  3. (iii)
    $p=3$, $K$ is one of the groups in Tables 1 - 2
Proof.
Suppose that $K_{2}$ is normal in $K$, that is, $K=K_{2}{:}K_{p}$. Then $K_{2}$ has an automorphism of order $p$. By Lemma 4.1, $K_{2}$ is one of the following groups:
${\sf Z}_{2}^{2},\ {\rm Q}_{8},\ {\sf Z}_{2}^{3}\ \mbox{and}\ {\sf Z}_{4}\circ{\rm Q}_{8},$
and either $p=3$, $K=K_{2}{:}K_{3}$; or $K_{2}={\sf Z}_{2}^{3}$, $p=7$ and $K={\sf Z}_{2}^{3}{:}K_{7}$, as in part (i).
We then assume that neither $K_{p}$ nor $K_{2}$ is normal in $K$. Let
$F={\rm Fit}(K)=F_{p}\times F_{2}$, with $F_{p}<K_{p}$, $F_{2}<K_{2}$.
Consider the quotient group $\overline{K}=K/F_{p}$, whose Fitting group is a 2-group, namely ${\rm Fit}(\overline{K})={\bf O}_{2}(\overline{K})$. Thus, by Lemma 5.5, $\overline{K}={\bf O}_{2}(\overline{K}){:}H$, where either
(1) ${\bf O}_{2}(\overline{K})\in\{{\sf Z}_{2}^{2},{\sf Z}_{2}^{3},{\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8}\}$, $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$; or (2) ${\bf O}_{2}(\overline{K})={\sf Z}_{2}^{3}$, $H={\sf Z}_{7}$ or ${\sf Z}_{3}$.
Further, if (2) holds, $\overline{K}={\sf Z}_{2}^{3}{:}{\sf Z}_{p}$ with $p=3$ or $7$, then ${\sf Z}_{2}^{3}$ is a minimal normal subgroup of $\overline{K}$, and so $\overline{F_{2}}=F_{2}$ as a normal subgroup of $\overline{K}$ must be identity. Consequently, we have $F=F_{p}$, ${\sf Z}_{2}^{3}{:}{\sf Z}_{p}=K/F_{p}\lesssim{\sf Out}(F_{p})$, which does not hold for the $p$-groups appearing in Lemma 4.3.
Therefore, we have that (1) holds, and since $\overline{K}={\bf O}_{2}(\overline{K}){:}H$ is known, $F_{2}=\overline{F_{2}}$ as a normal 2-subgroup of $\overline{K}$ is also known. We then analyze all the possibilities.
Case 1. $F_{2}={\bf O}_{2}(\overline{K})$. In this case, we have $H={\rm S}_{3}$ since $F_{2}<K_{2}$. Then
$K=F_{3}.(F_{2}{:}{\rm S}_{3})=(F_{3}\times F_{2}).{\rm S}_{3}=F_{2}{:}(F_{3}.{\rm S}_{3})=F_{2}{:}(K_{3}{:}{\sf Z}_{2}).$
Note that $K_{3}{:}{\sf Z}_{2}/F_{3}\cong{\rm S}_{3}$. By checking the $3$-groups appearing in Lemma 4.3, we conclude that there are seven possibilities for the triple $T:=(K_{3},\,K_{3}{:}{\sf Z}_{2},\,F_{3})$ as follows (also note that ${\sf Aut}({\sf Z}_{3}\times{\sf Z}_{3})={\rm GL}_{2}(3)$ is solvable):
[…]
Consequently, Case 1 gives the groups listed in Table 1.
Table 1. Case: $F_{2}={\bf O}_{2}(\overline{K})$
case $F_{2}$ ${\sf Z}_{2}^{2}$ ${\sf Z}_{2}^{3}$ ${\rm Q}_{8}$ ${\sf Z}_{4}\circ{\rm Q}_{8}$
(1.1) ${\rm S}_{4}$ ${\sf Z}_{2}\times{\rm S}_{4}$ ${\rm GL}_{2}(3)$ ${\sf Z}_{4}\circ{\rm GL}_{2}(3)$
(1.2) ${\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}}$ ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}})$ ${\rm Q}_{8}{:}{\rm D}_{2.3^{\ell+1}}$ ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell+1}})$
(1.3) ${\sf Z}_{3^{\ell}}\times{\rm S}_{4},$ ${\sf Z}_{2.3^{\ell}}\times{\rm S}_{4}$ ${\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\rm S}_{3})$ ${\sf Z}_{4.3^{\ell}}\circ({\rm Q}_{8}{:}{\rm S}_{3})$
(1.4) ${\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$, ${\sf Z}_{6}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{3}\times({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{12}\circ({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$
(1.5) $({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$ $({\sf Z}_{2.3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$ $({\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\rm S}_{3})){:}{\sf Z}_{2}$ $({\sf Z}_{4.3^{\ell}}\circ({\rm Q}_{8}{:}{\rm S}_{3})){:}{\sf Z}_{2}$
(1.6) $({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$ $({\sf Z}_{6}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$ $({\sf Z}_{3}\times({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$ $({\sf Z}_{12}\circ({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$
(1.7) ${\sf Z}_{2}^{2}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3})$ ${\sf Z}_{2}\times{\sf Z}_{2}^{2}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3})$ ${\rm Q}_{8}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3})$ ${\sf Z}_{4}\circ({\rm Q}_{8}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3}))$
Case 2. $(F_{2},{\bf O}_{2}(\overline{K}))=({\sf Z}_{2}^{2},{\sf Z}_{2}^{3})$ or $({\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8})$. Then
$K=F_{3}.({\bf O}_{2}(\overline{K}){:}H)=(F_{3}\times F_{2}).({\sf Z}_{2}\times H)=F_{2}{:}(F_{3}.({\sf Z}_{2}\times H)).$
Further, if $H={\sf Z}_{3}$, then $K/F_{2}=K_{3}{:}{\sf Z}_{2}$. Note that ${\sf Z}_{2}$ acts non-trivially on $K_{3}$, and meanwhile it acts trivially on $K_{3}/F_{3}=H$. Then by checking the $3$-groups appearing in Lemma 4.3, we conclude that there are three possibilities:
[…]
Otherwise, $H={\rm S}_{3}$, and so $K/F_{2}=K_{3}{:}({\sf Z}_{2}\times{\sf Z}_{2})$. By Lemma 4.3, ${\sf Z}_{2}\times{\sf Z}_{2}\leqslant{\sf Aut}(K_{3})$ implies $K_{3}={\sf Z}_{3^{\ell}}\times{\sf Z}_{3}$ and $K_{3}{:}({\sf Z}_{2}\times{\sf Z}_{2})={\rm D}_{2.3^{\ell}}\times{\rm D}_{6}$. We then get two possibilities:
[…]
Consequently, Case 2 gives the groups listed in Table 2.
Table 2. Case: $F_{2}<{\bf O}_{2}(\overline{K})$
case $(F_{2},{\bf O}_{2}(\overline{K}))$ $({\sf Z}_{2}^{2},{\sf Z}_{2}^{3})$ $({\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8})$
(2.1) ${\rm D}_{2.3^{\ell}}\times{\rm A}_{4}$ $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\sf Z}_{3})$
(2.2) $({\sf Z}_{2}^{2}{:}{{\sf Z}_{3^{\ell}}})\times{\rm D}_{6}$ $({\rm Q}_{8}{:}{{\sf Z}_{3^{\ell}}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$
(2.3) $({\rm D}_{2.3^{\ell}}\times{\sf Z}_{2}^{2}){:}{\sf Z}_{3}$ $(({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ{\rm Q}_{8}){:}{\sf Z}_{3}$
(2.4) ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\rm S}_{3})$
(2.5) $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$ $({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$
Case 3. Recall that $F_{2}$ is a normal 2-subgroup of ${\bf O}_{2}(\overline{K})$. Notice that ${\rm Q}_{8}{:}{\sf Z}_{3}$ or ${\rm Q}_{8}{:}{\rm S}_{3}$ only has the center ${\sf Z}_{2}$ and ${\rm Q}_{8}$ as normal $2$-subgroups. Then all the remaining cases are collected as follows :
[…]
For case (3.1), $(F_{2},{\bf O}_{2}(\overline{K}))=(1,{\rm Q}_{8})$, we have
$F=F_{3}\mbox{\ and\ }K/F_{3}={\rm Q}_{8}{:}H\lesssim{\sf Aut}(F_{3}).$
By lemma 4.3, it implies that $F_{3}={\sf Z}_{3}\times{\sf Z}_{3}$ and $K_{3}={\sf Z}_{9}{:}{\sf Z}_{3}$ is non-abelian. Let
$K_{3}=\langle a\rangle{:}\langle b\rangle={\sf Z}_{9}{:}{\sf Z}_{3}$, with $a^{b}=a^{4}$, $b^{a}=a^{6}b$,
where $F_{3}=\langle a^{3}\rangle\times\langle b\rangle$. Note that $K_{2}=({\rm Q}_{8}{:}H)_{2}={\rm Q}_{8}$ or ${\rm Q}_{8}{:}{\sf Z}_{2}$. Consider the central involution $c$ of $K_{2}$. Then $\overline{c}$ is the central involution of $\overline{K}$. On the one hand, $\overline{c}$ acts on $F_{3}$ by mapping each $x$ to $x^{-1}$. That is,
$x^{c}=x^{-1}$, so that, $c^{x}=x^{-2}c$ for any $x\in F_{3}$.
On the other hand, $\overline{c}^{\overline{a}}=\overline{c}$, and so $c^{a}=cy$ where $y\in F_{3}$. Let $y=a^{3i}b^{j}$. Then
$c^{a^{3}}=(ca^{3i}b^{j})^{a^{2}}=(ca^{(6i+6j)}b^{2j})^{a}=ca^{(9i+18j)}b^{3j}=c,\mbox{\ where\ }a^{3}\in F_{3},$
a contradiction. Hence case (3.1) cannot occur.
For case (3.2), $(F_{2},{\bf O}_{2}(\overline{K}))=(1,{\sf Z}_{4}\circ{\rm Q}_{8})$, we have ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\sf Z}_{3})\lesssim{\sf Aut}(F_{3})$, which is impossible by checking the 3-groups appearing in Lemma 4.3.
For case (3.3), since ${\bf O}_{2}(\overline{K})$ acts on $F_{3}$ with the centralizer ${\bf C}_{{\bf O}_{2}(\overline{K})}(F_{3})=F_{2}$, we deduce that in this case ${\sf Aut}(F_{3})$ contains a subgroup isomorphic to ${\rm A}_{4}\cong{\sf Z}_{2}^{2}{:}{\sf Z}_{3}$. By Lemma 4.3, $|{\sf Aut}(F_{3})|$ being divisible by $4$ implies $F_{3}={\sf Z}_{3^{\ell}}\times{\sf Z}_{3}$. For $\ell=1$, it is clear that ${\sf Aut}({\sf Z}_{3}^{2})={\rm GL}(3,2)\cong{\rm Q}_{8}{:}{\rm S}_{3}$ contains no subgroups isomorphic to ${\rm A}_{4}$. We thus assume $\ell>1$ and let $F_{3}=\langle a\rangle\times\langle b\rangle$ with $|a|=3^{\ell}$, $|b|=3$. By Lemma 4.3, ${\sf Aut}(F_{3})$ is solvable, and it contains a Hall $3^{\prime}$-subgroup
$H_{0}=({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{3^{\prime}}\cong{\sf Z}_{2}^{2},$
which stabilizes subgroups $\langle a\rangle$ and $\langle b\rangle$. Meanwhile, by simple calculation, $H_{0}$ cannot stabilize
    any of the other two subgroups of order $3^{\ell}$, i.e., $\langle ab\rangle$, $\langle ab^{-1}\rangle$, and
    any of the other two non-characteristic subgroups of $3$, i.e., $\langle a^{3^{\ell-1}}b\rangle$, $\langle a^{-3^{\ell-1}}b\rangle$.
Consider the normalizer $N={\bf N}_{{\sf Aut}(F_{3})}(H_{0})$. For any $g\in N$, $H_{0}=H_{0}^{g}$ stabilizes subgroups $\langle a\rangle^{g}$ and $\langle b\rangle^{g}$, and so $\langle a\rangle^{g}=\langle a\rangle$, $\langle b\rangle^{g}=\langle b\rangle$, which implies that $N\leqslant{\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle)$ is abelian. Note that any subgroup $H\leqslant{\sf Aut}(F_{3})$ with $H\cong{\sf Z}_{2}^{2}$ is a Hall $3^{\prime}$-subgroup, then conjugate to $H_{0}$. Then the normalizer ${\bf N}_{{\sf Aut}(F_{3})}(H)$, which is conjugate to $N$, is also abelian. Thus, ${\sf Aut}(F_{3})$ has no subgroup isomorphic to ${\rm A}_{4}$.
Finally, with Case 3 excluded, the proof is complete. $\Box$
Lemma 5.7.
Let $K$ be a $\{2,3,7\}$-group satisfying Hypotheses 2.2, with neither $K_{3}$ nor $K_{7}$ normal in $K$. Then $K=K_{2}{:}(K_{7}{:}K_{3})={\sf Z}_{2}^{3}{:}(K_{7}{:}K_{3})$.
Proof.
By Lemma 5.1, $K_{\{3,7\}}=K_{7}{:}K_{3}$. Thus $K_{\{2,7\}}\neq K_{7}{:}K_{2}$, and by Lemma 5.6,
$K_{\{2,7\}}=K_{2}{:}K_{7}={\sf Z}_{2}^{3}{:}K_{7},$
Let $F={\rm Fit}(K)=F_{2}\times F_{3}\times F_{7}$. As $K_{2}$ is normalized by $K_{7}$, and meanwhile it normalizes $F_{7}$, so $K_{2}$ centralizes $F_{7}$. Moreover, $K_{2}$ also centralizes $F_{3}$. If not, consider the subgroup $F_{3}{:}(K_{2}{:}K_{7})\leqslant K$. Then ${\sf Z}_{2}^{3}$ as a minimal normal subgroup of $K_{2}{:}K_{7}$ is embedded into ${\sf Aut}(F_{3})$, which is impossible due to Lemma 4.3 and the fact ${\sf Z}_{2}^{3}\nleqslant{\rm GL}(2,3)\cong{\rm Q}_{8}{:}{\rm S}_{3}$. Hence $K_{2}$ centralizes $F$, so $K_{2}\leqslant C_{K}(F)\leqslant F$, and then $K_{2}=F_{2}\lhd K$, and $K=K_{2}{:}K_{\{3,7\}}=K_{2}{:}(K_{7}{:}K_{3})$. $\Box$
We are now ready to prove the first main theorem of this paper.
Proof of Theorem 1.1:
Let $G$ be a finite solvable group of which each Sylow subgroup has a cyclic or dihedral subgroup of prime index. Let $H$ be the largest normal Hall subgroup of odd order, with $\pi=\pi(H)$. Let $K$ be a Hall $\pi^{\prime}$-subgroup of $G$, where $\pi^{\prime}=\pi(G)\setminus\pi(H)$. Then we have factorization
G=H:K.G=H{:}K.
Since $H$ is of odd order, by Lemma 5.4, H=A:BH=A{:}B such that $A$ is abelian, $B$ is nilpotent, and $A\cap{\bf Z}(H)=1$.
We then consider the subgroup $K$. Let $p$ be the largest prime divisor of $|G|$. If $p>7$, then by Lemmas 5.1 and 5.6, $G_{p}$ is normal in both $G_{2^{\prime}}$ and $G_{\{2,p\}}$, and so $G_{p}$ is normal in $G$, and $p\in\pi$. Then, by inductively considering $G/G_{p}$ we deduce that any prime divisor $r$ of $|G|$ with $r>7$ lies in $\pi$. That is, $\pi^{\prime}\subseteq\{2,3,5,7\}$. Further, by Lemmas 5.1, 5.2 and 5.6, $G_{5}$ is normalized by $G_{3}$, $G_{7}$ and $G_{2}$, respectively. Hence $5\notin\pi^{\prime}$. Thus, either
  • (i)
    $\pi^{\prime}=\emptyset$, $G=H$ is of odd order; or
  • (ii)
    $\pi^{\prime}=\{2\}$, $\{2,3\}$, $\{2,7\}$ or $\{2,3,7\}$.
At last, respectively for $\pi^{\prime}=\{2\}$, $\pi^{\prime}=\{2,3\}$ or $\{2,7\}$, and $\pi^{\prime}=\{2,3,7\}$, the group $K=G_{\pi^{\prime}}$ is described in Proposition 4.1, Lemma 5.6 and Lemma 5.7. $\Box$
6. Proof of Theorem 1.2
In this section we focus on the automorphism groups of regular maps.
The following lemma shows that the class of groups with regular triples is closed under taking quotients.
Lemma 6.1.
Let $G$ be a finite group which has a regular triple, and let $\overline{G}$ be a non-identity factor group of $G$. Then either $\overline{G}$ has a regular triple, or $\overline{G}$ is ${\sf Z}_{2}$ or a dihedral group. In particular, $\overline{G}$ is of even order.
Proof.
Let $(x,y,z)$ be a regular triple for $G$, and let $\overline{G}=G/N$ for a proper normal subgroup $N\lhd G$. If $\overline{x},\overline{y},\overline{z}$ are all involutions of $\overline{G}$ with $\overline{x}\neq\overline{z}$, then they form a regular triple for $\overline{G}$, since $\overline{G}=\langle\overline{x},\overline{y},\overline{z}\rangle$ and $\overline{x}\,\overline{z}=\overline{z}\,\overline{x}$; otherwise, either some of $\overline{x},\overline{y},\overline{z}$ equals the identity-element in $\overline{G}$, or $\overline{x}=\overline{z}$, while in both cases, $\overline{G}$ is generated by one or two involutions, namely, $G$ is ${\sf Z}_{2}$ or dihedral, respectively. $\Box$
Lemma 6.2.
Let $G={\rm Q}_{8}{:}{\rm S}_{3}\cong{\rm GL}(2,3)$. Then neither $G$ nor ${\sf Z}_{4}\circ G$ has a regular triple.
Proof.
Suppose that $G$ has a regular triple $(x,y,z)$. Since $xz=zx$, we have $\langle x,z\rangle\cong{\sf Z}_{2}^{2}$, and so $\{x,z,xz\}\cap{\rm Q}_{8}\not$$\emptyset$. That is, one of $x,z,xz$ has to be the central involution 1-1 of $G$. It follows that $\overline{G}=G/\langle-1\rangle=\langle\overline{x},\overline{y}\rangle$ or $\langle\overline{y},\overline{z}\rangle$ is a dihedral group, which is impossible as $\overline{G}={\sf Z}_{2}^{2}{:}{\rm S}_{3}$.
Suppose that $K={\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ has a regular triple $(x,y,z)$. Write $K$ as
$K=\langle u\rangle\circ({\rm Q}_{8}{:}(\langle\sigma\rangle{:}\langle\tau\rangle)),$
where $|u|=4$, ${\rm Q}_{8}=\pm\{1,i,j,k\}$ with $i^{2}=j^{2}=k^{2}=ijk=-1$, and $\sigma,\tau\in{\sf Aut}({\rm Q}_{8})$ such that
$\sigma:i\rightarrow j\rightarrow k\,(\rightarrow i),\mbox{\,\ and\ \,}\tau:i\leftrightarrow-j,j\leftrightarrow-i,k\leftrightarrow-k.$
Then the subgroup $\langle\sigma\rangle{:}\langle\tau\rangle\cong{\rm S}_{3}$ contains three involutions, i.e., $\tau_{k}=\tau$,
$\tau_{i}:j\leftrightarrow-k,k\leftrightarrow-j,i\leftrightarrow-i,\mbox{\ and\ }\tau_{j}:i\leftrightarrow-k,k\leftrightarrow-i,j\leftrightarrow-j.$
By calculation, we conclude that $K$ contains the following involutions:
$\{-1\}\cup\{\pm u\lambda\}\cup\{\pm\tau_{\lambda},\pm\lambda\tau_{\lambda}\},\mbox{\ where\ }\lambda\in\{i,j,k\}.$
We shall confirm the following claim:
Claim. The only involutions commutes with $u\lambda$ (or $u\lambda$) are $\pm u\lambda$ and 1-1.
In fact, for any $\lambda^{\prime}\in\{i,j,k\}$ with $\lambda\neq\lambda^{\prime}$, we have:
$\lambda^{\lambda^{\prime}}=\lambda^{\tau_{\lambda}}=\lambda^{\lambda\tau_{\lambda}}=-\lambda$, and $\lambda^{\tau_{\lambda^{\prime}}}=-\lambda^{\prime\prime},\ \lambda^{\lambda^{\prime}\tau_{\lambda^{\prime}}}=-\lambda^{\tau_{\lambda^{\prime}}}=\lambda^{\prime\prime}$,
where $\lambda^{\prime\prime}\in\{i,j,k\}\setminus\{\lambda,\lambda^{\prime}\}.$ As $u$ and 1-1 lie in the center of $K$, we then have:
$(\pm u\lambda)^{\pm u\lambda^{\prime}}=\pm u(\lambda^{\pm u\lambda^{\prime}})=\pm u(\lambda^{\lambda^{\prime}})=\mp u\lambda,$
$(\pm u\lambda)^{\pm\tau_{\lambda}}=\pm u(\lambda^{\pm\tau_{\lambda}})=\mp u\lambda,\ (\pm u\lambda)^{\pm\lambda\tau_{\lambda}}=\pm u(\lambda^{\pm\lambda\tau_{\lambda}})=\mp u\lambda,$
and
$(\pm u\lambda)^{\pm\tau_{\lambda^{\prime}}}=\pm u(\lambda^{\pm\tau_{\lambda^{\prime}}})=\mp u\lambda^{\prime\prime},\ (\pm u\lambda)^{\pm\lambda^{\prime}\tau_{\lambda^{\prime}}}=\pm u(\lambda^{\pm\lambda^{\prime}\tau_{\lambda^{\prime}}})=\pm u\lambda^{\prime\prime}.$
We have therefore confirmed the claim.
Let $H=\langle u,i,j,k\rangle\leqslant K$. Then $H\cong{\sf Z}_{4}\circ{\rm Q}_{8}$ is normal in $K$, and $K/H\cong{\rm S}_{3}$. Further, the involutions contained in $H$ are: $\{-1\}\cup\{\pm u\lambda\}$. Recall that $(x,y,z)$ is a regular triple for $G$. Thus, $\langle x,z\rangle\cong{\sf Z}_{2}^{2}$, and so $\langle x,z\rangle\cap H>1$, that is,
$\{x,z,xz\}\cap(\{-1\}\cup\{\pm u\lambda\})\not$$\emptyset$.
By the claim given above, one of $x,z,xz$ has to be 1-1. Thus, $\overline{K}=K/\langle-1\rangle=\langle\overline{x},\overline{y}\rangle$ or $\langle\overline{y},\overline{z}\rangle$ is a dihedral group, which is impossible as $\overline{K}=K/\langle-1\rangle={\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\rm S}_{3})$. $\Box$
Now we are ready to prove Theorem 1.2.
Proof of Theorem 1.2:
Let $\mathcal{M}$} be a regular map with square-free Euler characteristic, and let $G={\sf Aut}{\mathcal{M}}$. By Lemma 2.1, each Sylow subgroup of $G$ has a cyclic or dihedral subgroup of prime index, and so G=(A:B):KG=(A{:}B){:}K as described in Theorem 1.1.
Note that $G$ has a regular triple. Then it is of even order, and so $K\neq 1$. Further, if $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}G_{7}$ or $K=G_{\{2,3,7\}}={\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3})$, then $K$ has $G_{2}={\sf Z}_{2}^{3}$ as a minimal normal subgroup. By checking the $p$-groups appearing in Lemma 4.3, meanwhile noticing that ${\sf Z}_{2}^{3}$ cannot be embedded into ${\rm GL}(2,p)$ for all primes $p$ (as a faithful representation of ${\sf Z}_{2}^{3}$ on $\mathbb{F}_{p}$ is of degree at least $3$), we deduce that $G_{2}={\sf Z}_{2}^{3}$ centralizes each Sylow subgroup of A:BA{:}B. Hence $G_{2}\lhd G$, and so $G$ has an odd factor group $G/G_{2}$, leading to a contradiction by Lemma 6.1. Therefore, $K=G_{2}$ or $G_{\{2,3\}}$.
Suppose $K=G_{2}$. Since K=G/(A:B)K=G/(A{:}B) is a factor group of $G$, by Lemma 6.1, either $K$ has a regular triple and then it is one of the groups given in Proposition 4.2; or $K$ is ${\sf Z}_{2}$ or a dihedral group. In conclusion, $K$ is ${\sf Z}_{2}$, ${\rm D}_{2^{\ell+1}}$, ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$ or ${\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}$, where $\ell\geqslant 1$.
Suppose $K=G_{\{2,3\}}$, which is one of the groups given in Lemma 5.6 (ii) - (iii). Since K=G/(A:B)K=G/(A{:}B) is a factor group of $G$, by Lemma 6.1, either $K$ has a regular triple, or $K$ is a dihedral group. Further, note that a factor group of a dihedral group is also dihedral (or ${\sf Z}_{2}$). Thus, $K$ has no odd factor groups, and has no factor groups isomorphic to ${\rm Q}_{8}{:}{\rm S}_{3}$ or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ either, because by Lemma 6.2 neither ${\rm Q}_{8}{:}{\rm S}_{3}$ nor ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ has a regular triple, and meanwhile, neither of them is a dihedral group. Then, by checking the groups appearing in Lemma 5.6 (ii) - (iii), we find that the only candidates for $K=G_{\{2,3\}}$ are the following groups in Table 1: $K=X$ or $X\times{\sf Z}_{2}$, where $X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\};$ and the two groups in Table 2: $K={\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$. $\Box$
Declarations
Funding and/or Conflicts of interests/Competing interests
The author(s) declares that there is no any financial or personal relationship with other people or organizations not mentioned that can inappropriately influence the work.

### v2

@@
-Finite solvable groups of square-free Euler characteristic
+Finite groups and arc-transitive maps of square-free Euler characteristic
 Abstract.
-A characterization is completed for finite groups which act regularly on maps with square-free Euler characteristic, associated with infinite families of square-free Euler characteristic regular maps presented.
+A characterization is completed for finite groups acting arc-transitively on maps with square-free Euler characteristic, associated with infinite families of regular maps of square-free Euler characteristic presented.
 This is based on a classification of finite groups of which each Sylow subgroup has a cyclic or dihedral subgroup of prime index.
-Keywords: solvable group, regular map, square-free, Euler characteristic
+Keywords: solvable group, arc-transitive map, regular map, Euler characteristic
 The project was partially supported by the NNSF of China (11931005).
-November 18, 2025
+December 1, 2025
 1. Introduction
-Edge-transitive maps are categorized into fourteen classes according to local structures and local actions in [9, 15].
+Edge-transitive maps are categorized into fourteen classes according to local structures and local actions in [9, 16], among which five are arc-transitive.
 The problem of constructing and classifying important classes of such maps with specific Euler characteristic has attracted considerable attention; refer to [4] - [7] for the Euler characteristic to be a negative prime or product of two primes.
-We aim towards to a classification of regular maps with square-free Euler characteristic.
+We aim towards to a classification of arc-transitive maps with square-free Euler characteristic.
 Assume that $G$ is a finite group acting edge-transitively on a map with square-free Euler characteristic, which is said to be of square-free Euler characteristic for convenience.
@@
 each Sylow subgroup has a cyclic or dihedral subgroup of prime index.
-It leads us to studying groups of square-free Euler characteristic.
-This project extends the characterization of almost-Sylow cyclic groups, namely, each Sylow subgroup of odd order is cyclic, and the Sylow $2$-subgroup either is trivial or contains a cyclic subgroup of index 2; refer to Zassenhaus [18], Suzuki [16] and Wong [17].
+It leads us to studying groups of square-free Euler characteristic, which in fact extends the characterization of almost-Sylow cyclic groups, namely, each Sylow subgroup of odd order is cyclic, and the Sylow $2$-subgroup either is trivial or contains a cyclic subgroup of index 2; refer to Zassenhaus [19], Suzuki [17] and Wong [18].
 In previous work [11], we characterized non-solvable groups of square-free Euler characteristic, and in [10], we determined $2$-groups of square-free Euler characteristic.
@@
 A finite group $G$ is a semidirect product of $A$ by $B$ if $A\lhd G$ and $B<G$ are such that $G/A\cong B$, denoted by G=A:BG=A{:}B.
-Our first main result is the following theorem, which characterizes finite solvable groups of which each Sylow subgroup has a cyclic or dihedral subgroup of prime index.
+Our first main result is the following theorem.
 Theorem 1.1.
@@
     • (b)
-      $K=G_{\{2,3\}}$, explicitly given in Lemma 5.6 (ii) - (iii);
+      $K=G_{\{2,3\}}$, explicitly given in Lemma 5.6 (2) - (3);
     • (c)
       $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}G_{7}$, or $K=G_{\{2,3,7\}}={\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3})$.
-The next theorem shows that automorphism groups of regular maps with square-free Euler characteristic are more restricted.
+The next theorem shows that arc-transitive automorphism groups of maps with square-free Euler characteristic are more restricted.
 Theorem 1.2.
-Let $\mathcal{M}$} be a regular map with square-free Euler characteristic, and let $G={\sf Aut}{\mathcal{M}}$ be solvable.
-Then G=(A:B):KG=(A{:}B){:}K, where $A$ is abelian, $B$ is nilpotent, $A\cap{\bf Z}(A{:}B)=1$, and $K=G_{2}$ or $G_{\{2,3\}}$ is one of the following groups:
-  1. (i)
-    $K={\sf Z}_{2}$, ${\rm D}_{2^{\ell+1}}$, ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$ or ${\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}$, where $\ell\geqslant 1;$
-  2. (ii)
-    $K=X$ or $X\times{\sf Z}_{2}$, where $X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\};$
-  3. (iii)
-    $K={\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$.
-The paper is organized as follows.
-After the Introduction, we collect some preliminaries in Section 2, and construct several infinite families of regular maps with square-free Euler characteristic in Section 3.
-Then, we prove our two main results in Section 4 - 5 and Section 6, respectively.
+Let $\mathcal{M}$} be a map with square-free Euler characteristic, and let $G\leqslant{\sf Aut}{\mathcal{M}}$ be solvable and arc-transitive on $\mathcal{M}$}.
+Then G=(A:B):KG=(A{:}B){:}K such that $A$ is abelian with $A\cap{\bf Z}(A{:}B)=1$, $B$ is nilpotent, and either $K=G_{2}$ is ${\sf Z}_{2}$ or one of the groups given in Proposition 4.2, or one of the following holds:
+  • (1)
+    $\mathcal{M}$} is $G$-regular, $K=G_{\{2,3\}}$ is one of $X$, $X\times{\sf Z}_{2}$, ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$, where $X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\};$
+  • (2)
+    $\mathcal{M}$} is $G$-vertex-reversing, $K=G_{\{2,3\}}$ is one of the groups in Table 1 and Table 2, corresponding to cases (1.1), (1.2), (1.5), (1.6), (2.4), (2.5);
+  • (3)
+    $\mathcal{M}$} is $G$-vertex-rotary, $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}{\sf Z}_{7^{\ell}}$, or $K=G_{\{2,3\}}$ is one of ${\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}}$, ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})$, ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})$, or one of the groups in Table 1 and Table 2, excluding cases (1.5), (1.6).
+The proofs of the above two results are given in Section 4 - 5 and Section 6.
+Before that, we construct several interesting examples of regular maps with square-free Euler characteristic in Section 3, from which the following result follows.
+Theorem 1.3.
+There are infinite families of regular maps $\mathcal{M}$} with underlying graphs ${\bf C}_{n}^{(n)}$ or ${\bf C}_{n}\square{\bf C}_{n}$, and with Euler characteristic $\chi({\mathcal{M}})=-n(n-3)$ or n(n2)-n(n-2), respectively.
+In particular, there are infinitely many of them with $\chi({\mathcal{M}})$ square-free.
 2. Preliminaries
@@
 For a subgroup $H\leqslant G$, denote by ${\bf N}_{G}(H)$ and ${\bf C}_{G}(H)$ the normalizer and the centralizer of $H$ in $G$, respectively.
-Let $\it\Gamma$} be a graph and $\lambda$ be a positive integer. Denote by ${\it\Gamma}^{(\lambda)}$ the graph with each edge of $\it\Gamma$} replaced by $\lambda$ edges.
-It is known (refer to [12]) that an arc-transitive (multi-) graph $\it\Gamma$} can be expressed as a coset graph :
-${\it\Gamma}={\sf Cos}(G,H,J),\ \mbox{or}\ {\sf Cos}(G,H,HgH)\ \mbox{(for simple graphs)},$
-which has vertex set $[G:H]$ and edge set $[G:J]$, such that
-    $Hg_{1}$ is incident with $Jg_{0}$ if and only if $g_{1}g_{0}^{-1}\in HJ$; or equivalently,
-    $Hg_{1}$, $Hg_{2}$ are adjacent if and only if $g_{2}g_{1}^{-1}\in HgH$ in the simple graph case.
-A map $\mathcal{M}$} is a $2$-cell embedding of a graph into a closed surface.
-Let ${\mathcal{M}}=(V,E,F)$ be a map with vertex set $V$, edge set $E$ and face set $F$, and let $G={\sf Aut}{\mathcal{M}}$.
+Let ${\it\Gamma},\Sigma$ be two graphs and $\lambda$ be a positive integer.
+Denote by ${\it\Gamma}^{(\lambda)}$ the graph with each edge of $\it\Gamma$} replaced by $\lambda$ edges.
+Denote by ${\it\Gamma}\square\Sigma$ the Cartesian product of $\it\Gamma$} and $\Sigma$, which has vertex set $V({\it\Gamma})\times V(\Sigma)$, where $(x_{1},y_{1})$ is adjacent to $(x_{2},y_{2})$ if and only if $x_{1}=y_{1}$ and $x_{2}$ is adjacent to $y_{2}$, or $x_{1}$ is adjacent to $y_{1}$ and $x_{2}=y_{2}$.
+A map ${\mathcal{M}}=(V,E,F)$ is a $2$-cell embedding of a graph into a closed surface, with vertex set $V$, edge set $E$ and face set $F$.
 The Euler characteristic of $\mathcal{M}$} is given by Euler formula:
@@
 A flag of $\mathcal{M}$} is an incidence triple $(\alpha,e,f)$ with $\alpha\in V$, $e\in E$ and $f\in F$.
-Given a flag $(\alpha,e,f)$, it is well-known that each of the stabilizers $G_{\alpha}$ and $G_{f}$ is cyclic or dihedral, and meanwhile, $G_{e}=1$, ${\sf Z}_{2}$ or ${\rm D}_{4}$.
-Further, if $\mathcal{M}$} is a flag-regular map (also called a regular map), i.e., $G$ acts regularly on flags, then there exists a triple $(x,y,z)$ of involutions of $G$ which determines the map $\mathcal{M}$} as follows:
-$G_{\alpha}=\langle x,y\rangle$, $G_{e}=\langle x,z\rangle\cong{\rm D}_{4}$, $G_{f}=\langle y,z\rangle$.
-Meanwhile, the Euler characteristic has the form
-$\mbox{$\chi({\mathcal{M}})={|G|\over|G_{\alpha}|}-{|G|\over 4}+{|G|\over|G_{f}|}$}.$
-Conversely, if a group $G$ has a triple of involutions $(x,y,z)$ such that
-$G=\langle x,y,z\rangle$ and $xz=zx$ $(x\neq z)$,
-then there exists a unique regular map ${\mathcal{M}}=(V,E,F)$ with
-$V=[G:\langle x,y\rangle],\ E=[G:\langle x,z\rangle],\ F=[G:\langle y,z\rangle],$
-and the incidence relation defined by non-empty intersection, such that $G={\sf Aut}{\mathcal{M}}$ acts regularly on flags.
-Hence, such a triple of involutions is called a regular triple for $G$, and the induced regular map is denoted by ${\mathcal{M}}(G,x,y,z)$.
-Moreover, the underlying graph of ${\mathcal{M}}(G,x,y,z)$ can be expressed as a coset graph:
-${\sf Cos}(G,\langle x,y\rangle,\langle x,z\rangle)$, or ${\sf Cos}(G,H,HzH)$ where $H=\langle x,y\rangle$ (for simple graphs).
-In conclusion, each of the following two things determines the other:
-    the regular maps arisen from a group $G$;
-    the regular triples for $G$.
-The following lemma established in [11] characterizes Sylow subgroups of $G={\sf Aut}({\mathcal{M}})$ in the case where $\mathcal{M}$} is with square-free characteristic.
+Let $G\leqslant{\sf Aut}({\mathcal{M}})$, and let $(\alpha,e,f)$ be a flag.
+It is well-known that each of the stabilizers $G_{\alpha}$ and $G_{f}$ is cyclic or dihedral, and meanwhile, $G_{e}=1$, ${\sf Z}_{2}$ or ${\rm D}_{4}$.
+A map $\mathcal{M}$} is called a $G$-arc-transitive map, if $G$ acts transitively on arcs (directed edges) of $\mathcal{M}$}.
+Arc-transitive maps are divided into five types in [9, 16], for which recent papers [12, 13] provide a new description: letting $\mathcal{M}$} be a $G$-arc-transitive map, then one of the following holds:
+(1) $\mathcal{M}$} is a $G$-regular map (type 1), which is determined a regular triple $(x,y,z)$ for $G$ satisfying:
+$\mbox{$G=\langle x,y,z\rangle$, $|x|=|y|=|z|=2$ and $\langle x,z\rangle\cong{\sf Z}_{2}^{2}$}.$
+In this case, we have:
+$G={\sf Aut}({\mathcal{M}}),\,G_{\alpha}=\langle x,y\rangle,\,G_{e}=\langle x,z\rangle,\,G_{f}=\langle y,z\rangle,$
+and the underlying graph of $\mathcal{M}$} can be expressed as a coset graph ${\sf Cos}(G,H,HzH)$, where $H=\langle x,y\rangle$ (refer to [12].)
+(2) $\mathcal{M}$} is a $G$-vertex-reversing map (type $2^{\ast}$ or $2^{P}$), which is determined by a reversing triple $(x,y,z)$ for $G$ satisfying: $G=\langle x,y,z\rangle$ and $|x|=|y|=|z|=2$.
+(3) $\mathcal{M}$} is a $G$-vertex-rotary map (type $2^{*}{\rm ex}$ or $2^{P}{\rm ex}$), which is determined by a rotary pair $(\alpha,z)$ for $G$ satisfying: $G=\langle\alpha,z\rangle$ and $|z|=2$.
+The following lemma established in [11] characterizes Sylow subgroups of ${\sf Aut}({\mathcal{M}})$ in the case where $\mathcal{M}$} is with square-free characteristic.
 We note that if $\chi({\mathcal{M}})$ is square-free, then of course so is $\gcd(\chi({\mathcal{M}}),|{\sf Aut}{\mathcal{M}}|)$.
@@
 ([11, Lem.
-2.2]) Let $\mathcal{M}$} be a map, and let $G={\sf Aut}{\mathcal{M}}$.
-Assume that $\gcd(\chi({\mathcal{M}}),|G|)$ is square-free.
-Then each Sylow $p$-subgroup of $G$ has a cyclic or dihedral subgroup of index $p$.
+2.2]) Let $\mathcal{M}$} be a map such that $\gcd(\chi({\mathcal{M}}),|{\sf Aut}({\mathcal{M}})|)$ is square-free.
+Then each Sylow $p$-subgroup of ${\sf Aut}({\mathcal{M}})$ has a cyclic or dihedral subgroup of index $p$.
 We thus study finite groups satisfying the following hypothesis.
@@
 $\chi({\mathcal{M}})=|G|\left({1\over|\langle x,y\rangle|}-{1\over 4}+{1\over|\langle y,z\rangle|}\right)=4n^{2}\left({1\over 4n}-{1\over 4}+{1\over 2n}\right)=-n(n-3).$
-As Lemma 3.2 shows, there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free.
+By Lemma 3.2, there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free.
 Further, let $n=4x+2$.
-Then $\chi({\mathcal{M}}_{n})=-2(2x+1)(4x-1)$.
-According to [8], there are infinitely many positive integers $x$ for which f(x)=(2x+1)(4x1)f(x)=(2x+1)(4x-1) is square-free, so there are infinitely many even values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free.
+Then according to [8], there are infinitely many positive integers $x$ for which f(x)=(2x+1)(4x1)f(x)=(2x+1)(4x-1) is square-free, so there are infinitely many even values of $n$ such that $\chi({\mathcal{M}}_{n})=-2(2x+1)(4x-1)$ is square-free.
 $\Box$
@@
 Further, either ${\sf Aut}(G)$ is a $2$-group, or $(G,{\sf Aut}(G))$ is one of $({\sf Z}_{2}^{2},{\rm S}_{3}),\,({\sf Z}_{2}^{3},{\rm GL}_{3}(2)),\,({\rm Q}_{8},{\rm S}_{4}),\,({\rm Q}_{8}\circ{\sf Z}_{4},{\rm S}_{4}\times{\sf Z}_{2})$.
-  • (i)
+  • (1)
     $G$ is one of the following groups, with a cyclic subgroup of index $2$, $\ell\geqslant 1$:
-    • (i.1)
-      $G={\sf Z}_{2^{\ell}}$ or ${\sf Z}_{2^{\ell}}\times{\sf Z}_{2};$
-    • (i.2)
+    • (1.1)
+      $G={\sf Z}_{2^{\ell+1}}$ or ${\sf Z}_{2^{\ell}}\times{\sf Z}_{2};$
+    • (1.2)
       $G={\rm D}_{2^{\ell+2}}$ or ${\rm Q}_{2^{\ell+2}};$
-    • (i.3)
+    • (1.3)
       $G={\sf Z}_{2^{\ell+2}}{:}{\sf Z}_{2}=\langle a\rangle{:}\langle b\rangle$ with $a^{b}=a^{2^{\ell+1}\pm 1};$
-  • (ii)
+  • (2)
     $G$ is one of the following groups, with a dihedral subgroup of index $2$, $\ell\geqslant 1$:
-    • (ii.1)
+    • (2.1)
       $G={\rm D}_{2^{\ell+1}}\times{\sf Z}_{2};$
-    • (ii.2)
+    • (2.2)
       $G={\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}=\langle a,b\rangle{:}\langle c\rangle$, where $\langle a\rangle{:}\langle b\rangle={\rm D}_{2^{\ell+3}}$, $(a,b)^{c}=(a^{2^{\ell+1}+1},b);$
-    • (ii.3)
+    • (2.3)
       $G={\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4}$.
 Proposition 4.2.
-([10], Cor. 4.9) Let $G$ be a $2$-group which has a cyclic or dihedral subgroup of index $2$. If $G$ has a regular triple, then it is one of the following groups:
-  • (i)
-    $G={\rm D}_{2^{\ell+1}}$ or ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$, $\ell\geqslant 1;$
-  • (ii)
-    $G={\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}=\langle a,b\rangle{:}\langle c\rangle$, where $\langle a\rangle{:}\langle b\rangle={\rm D}_{2^{\ell+3}}$, $(a,b)^{c}=(a^{2^{\ell+1}+1},b)$, $\ell\geqslant 1$.
+([10], Thm. 1.1) Let $G$ be a $2$-group which has a cyclic or dihedral subgroup of index $2$. Then:
+  • (1)
+    If $G$ has a reversing triple, then $G={\rm D}_{2^{\ell+1}}$, ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$, ${\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}$ or ${\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4}$, where $\ell\geqslant 1;$ further, if $G$ has a regular triple, then $G\neq{\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4};$
+  • (2)
+    If $G$ has a rotary pair, then $G={\sf Z}_{2^{\ell+1}}$, ${\sf Z}_{2^{\ell}}\times{\sf Z}_{2}$, ${\rm D}_{2^{\ell+2}}$ or ${\sf Z}_{2^{\ell+2}}{:}{\sf Z}_{2}$, $\ell\geqslant 1$.
 For the odd primes $p$, we have the following characterization.
@@
 In particular, each prime divisor of $|{\sf Aut}(G)|$ is less than or equal to $p$.
-  • (i)
+  • (1)
     $G={\sf Z}_{p^{\ell}}$ is cyclic, and ${\sf Aut}(G)={\sf Z}_{(p-1)p^{\ell-1}}$.
-  • (ii)
+  • (2)
     $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$ is abelian, and
-    1. (ii.1)
+    1. (2.1)
       if $\ell>1$, then ${\sf Aut}(G)$ is solvable, and letting $G=\langle a,b\rangle$, then any Hall $p^{\prime}$-subgroup of ${\sf Aut}(G)$ is conjugate to $({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{p^{\prime}}={\sf Z}_{p-1}\times{\sf Z}_{p-1};$
-    2. (ii.2)
+    2. (2.2)
       if $\ell=1$, then ${\sf Aut}(G)={\rm GL}(2,p)$, and further, any subgroup $M\leqslant{\rm GL}(2,p)$ of order coprime to $2p$ is an abelian group.
-  • (iii)
+  • (3)
     $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ is non-abelian with $a^{b}=a^{p^{\ell-1}+1}$, $\ell\geqslant 2$, and ${\sf Aut}(G)$ is solvable such that any Hall $p^{\prime}$-subgroups of ${\sf Aut}(G)$ is conjugate to $\langle\rho\rangle={\sf Z}_{p-1}$, where $a^{\rho}=a^{\xi}\not=a$ with $\xi\,\,\big|\,\,p-1$, and $b^{\rho}=b$.
 Proof.
-According to [13, 5.3.4, pp.136], $G$ is one of the groups stated in this lemma. We only need to verify the statements regarding the automorphism groups of $G$.
-(1) The automorphism groups of cyclic groups are well-known; refer to [14, Thm. 7.3].
-(2) Suppose $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$ with $\ell>1$. Then by [1, Thm. 3.2], ${\sf Aut}(G)$ is of order
+According to [14, 5.3.4, pp.136], $G$ is one of the groups stated in this lemma. We only need to verify the statements regarding the automorphism groups of $G$.
+(i) The automorphism groups of cyclic groups are well-known; refer to [15, Thm. 7.3].
+(ii) Suppose $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$ with $\ell>1$. Then by [1, Thm. 3.2], ${\sf Aut}(G)$ is of order
 $\begin{array}[]{ccl}|{\sf Aut}(G)|&=&|{\sf Aut}({\sf Z}_{p^{\ell}})|\cdot|{\sf Aut}({\sf Z}_{p})|\cdot|{\sf Hom}({\sf Z}_{p^{\ell}},{\sf Z}_{p})|\cdot|{\sf Hom}({\sf Z}_{p},{\sf Z}_{p^{\ell}})|\\ &=&(p-1)p^{\ell-1}\cdot(p-1)\cdot p\cdot p=p^{\ell+1}(p-1)^{2}.\end{array}$
@@
 Moreover, it follows from ${\sf Aut}(G)/{\rm Ker}(\sigma)\cong{\rm Im}(\sigma)$ that $|{\rm Im}(\sigma)|=p(p-1)^{2}=|P|$, and so ${\rm Im}(\sigma)=P$.
-Then ${\sf Aut}(G)$ is solvable, since both ${\rm Ker}(\sigma)$ and ${\rm Im}(\sigma)$ are solvable. Note that the following subgroup
+Now both ${\rm Ker}(\sigma)$ and ${\rm Im}(\sigma)$ are solvable, so is ${\sf Aut}(G)$. Note that the following subgroup
 $X=({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{p^{\prime}}={\sf Z}_{p-1}\times{\sf Z}_{p-1}$
 is a Hall $p^{\prime}$-subgroup of $G$. Then any other Hall $p^{\prime}$-subgroup is conjugate to $X$.
-(3) Suppose $G={\sf Z}_{p}\times{\sf Z}_{p}$.
-Then ${\sf Aut}(G)={\rm GL}(2,p)$, of order $p(p+1)(p-1)^{2}$ with prime divisors less than or equal to $p$.
+(iii) Suppose $G={\sf Z}_{p}\times{\sf Z}_{p}$.
+Then ${\sf Aut}(G)={\rm GL}(2,p)$ is of order $p(p+1)(p-1)^{2}$, with prime divisors less than or equal to $p$.
 Further, let $M\leqslant{\rm GL}(2,p)$ be a subgroup of order coprime to $2p$, and let $N=M/{\bf Z}(M)$.
@@
 Then $M$ is an abelian group, since $N=M/{\bf Z}(M)$ is cyclic.
-(4) Suppose that $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ with $a^{b}=a^{p^{\ell-1}+1}$ is non-abelian, $\ell\geqslant 2$.
+(iv) Suppose that $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ with $a^{b}=a^{p^{\ell-1}+1}$ is non-abelian, $\ell\geqslant 2$.
 The automorphism group ${\sf Aut}(G)$ is described in [2, Thm.
@@
 Proof.
-It is proved in a similar way as the previous lemma. Note that by Lemma 4.3 any prime divisor of $|{\sf Aut}(G_{p})|$ divides p(p1)(p+1)p(p-1)(p+1). $\Box$
+By Lemma 5.1, $G=G_{p}{:}G_{q}$.
+Since $G$ is solvable, we obtain that $G_{q}/{\bf C}_{G_{q}}(G_{p})\cong G/(G_{p}{\bf C}_{G_{q}}(G_{p}))=G/(G_{p}{\bf C}_{G}(G_{p}))\leq{\sf Out}(G_{p})$.
+By Lemma 4.3, any prime divisor of $|{\sf Out}(G_{p})|$ divides $p(p^{2}-1)$.
+However, $q\nmid p(p^{2}-1)$ by our assumption, and we conclude that $G_{q}/{\bf C}_{G_{q}}(G_{p})=1$, so that $G_{q}={\bf C}_{G_{q}}(G_{p})$ and $G=G_{p}\times G_{q}$.
+$\Box$
 Before giving a general characterization of solvable groups of odd order, we need the following technical lemma.
@@
 Proof.
-By Lemma 4.3 (i), the order of $x$ divides p1p-1. Let $H=\langle a\rangle$ and suppose
+By Lemma 4.3 (1), the order of $x$ divides p1p-1. Let $H=\langle a\rangle$ and suppose
 $a^{x}=a^{m}$, where $1<m\leqslant p^{\ell}-1$,  $p\nmid m$  and  $p^{\ell}\mid m^{p-1}-1$.
@@
 It then follows by Lemma 5.3 (i) that $A_{r}\leqslant X^{\prime}\leqslant{\bf C}_{X}(G_{p})$.
-Case 2. Let $A_{r}={\sf Z}_{r^{k}}\times{\sf Z}_{r}$, where $k\geqslant 1$. Then $r$ does not divide $|B|$. By Lemma 4.3 (ii), it implies that either
+Case 2. Let $A_{r}={\sf Z}_{r^{k}}\times{\sf Z}_{r}$, where $k\geqslant 1$. Then $r$ does not divide $|B|$. By Lemma 4.3 (2), it implies that either
   • (i)
@@
 Suppose $G_{p}={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}=\langle a\rangle{:}\langle b\rangle$.
-By Lemma 4.3 (iii), $B$ acts on $\langle a\rangle$ non-trivially and acts on $\langle b\rangle$ trivially.
+By Lemma 4.3 (3), $B$ acts on $\langle a\rangle$ non-trivially and acts on $\langle b\rangle$ trivially.
 Rewrite $G$ as
@@
 Suppose $G_{p}={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$, where $\ell\geqslant 1$.
-By Lemma 4.3 (ii), it implies that either (i) $G_{p}={\sf Z}_{p}^{2}$ and $B$ is irreducible on $G_{p}$; or (ii) there exists $a,b\in G_{p}$ such that $G_{p}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$ and $\langle b\rangle$ are normalized by $B$.
+By Lemma 4.3 (2), it implies that either (i) $G_{p}={\sf Z}_{p}^{2}$ and $B$ is irreducible on $G_{p}$; or (ii) there exists $a,b\in G_{p}$ such that $G_{p}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$ and $\langle b\rangle$ are normalized by $B$.
 For the former, G=C:DG=C{:}D satisfies the theorem, with $C=G_{p}\times A$, $D=B$.
-For the former, if $B$ acts non-trivially on both $\langle a\rangle$ and $\langle b\rangle$, then also let $C=G_{p}\times A$, $D=B$.
+For the latter, if $B$ acts non-trivially on both $\langle a\rangle$ and $\langle b\rangle$, then also let $C=G_{p}\times A$, $D=B$.
 Otherwise, $B$ acting non-trivially on $G_{p}$ implies $B$ acting non-trivially on some one of $\langle a\rangle$ and $\langle b\rangle$, say $\langle a\rangle$.
@@
 Let $G$ be a solvable group satisfying Hypothesis 2.2. If $F:={\rm Fit}(G)={\bf O}_{2}(G)$, then either $G$ is a $2$-group, or $G$ is one of the following:
-  • (i)
+  • (1)
     $G={\sf Z}_{2}^{3}{:}{\sf Z}_{7}$ or ${\sf Z}_{2}^{3}{:}{\sf Z}_{7}{:}{\sf Z}_{3};$
-  • (ii)
-    $G={\sf Z}_{2}^{2}{:}H$, ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}H)$, ${\rm Q}_{8}{:}H$, or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}H)$, where $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$, ${\sf Z}_{2}^{2}{:}H\cong{\rm A}_{4}$ or ${\rm S}_{4}$, and ${\rm Q}_{8}{:}H\cong{\rm SL}(2,3)$ or ${\rm GL}(2,3)$.
+  • (2)
+    $G={\sf Z}_{2}^{2}{:}H$, ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}H)$, ${\rm Q}_{8}{:}H$ or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}H)$, where $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$, ${\sf Z}_{2}^{2}{:}H\cong{\rm A}_{4}$ or ${\rm S}_{4}$, and ${\rm Q}_{8}{:}H\cong{\rm SL}(2,3)$ or ${\rm GL}(2,3)$.
 Proof.
@@
 Let $K$ be a $\{2,p\}$-group with $p$ odd prime satisfying Hypotheses 2.2. If $K\neq K_{p}{:}K_{2}$, then one of the following holds.
-  1. (i)
+  1. (1)
     $p=7$, $K={\sf Z}_{2}^{3}{:}K_{7};$
-  2. (ii)
+  2. (2)
     $p=3$, $K=X{:}K_{3}$ with $X\in\{{\sf Z}_{2}^{2},\,{\sf Z}_{2}^{3},\,{\rm Q}_{8},\,{\sf Z}_{4}\circ{\rm Q}_{8}\};$
-  3. (iii)
+  3. (3)
     $p=3$, $K$ is one of the groups in Tables 1 - 2
@@
 ${\sf Z}_{2}^{2},\ {\rm Q}_{8},\ {\sf Z}_{2}^{3}\ \mbox{and}\ {\sf Z}_{4}\circ{\rm Q}_{8},$
-and either $p=3$, $K=K_{2}{:}K_{3}$; or $K_{2}={\sf Z}_{2}^{3}$, $p=7$ and $K={\sf Z}_{2}^{3}{:}K_{7}$, as in part (i).
+and either $p=3$, $K=K_{2}{:}K_{3}$; or $K_{2}={\sf Z}_{2}^{3}$, $p=7$ and $K={\sf Z}_{2}^{3}{:}K_{7}$, as in (1) - (2).
 We then assume that neither $K_{p}$ nor $K_{2}$ is normal in $K$. Let
@@
 Thus, by Lemma 5.5, $\overline{K}={\bf O}_{2}(\overline{K}){:}H$, where either
-(1) ${\bf O}_{2}(\overline{K})\in\{{\sf Z}_{2}^{2},{\sf Z}_{2}^{3},{\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8}\}$, $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$; or (2) ${\bf O}_{2}(\overline{K})={\sf Z}_{2}^{3}$, $H={\sf Z}_{7}$ or ${\sf Z}_{3}$.
-Further, if (2) holds, $\overline{K}={\sf Z}_{2}^{3}{:}{\sf Z}_{p}$ with $p=3$ or $7$, then ${\sf Z}_{2}^{3}$ is a minimal normal subgroup of $\overline{K}$, and so $\overline{F_{2}}=F_{2}$ as a normal subgroup of $\overline{K}$ must be identity.
+(i) ${\bf O}_{2}(\overline{K})\in\{{\sf Z}_{2}^{2},{\sf Z}_{2}^{3},{\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8}\}$, $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$; or (ii) ${\bf O}_{2}(\overline{K})={\sf Z}_{2}^{3}$, $H={\sf Z}_{7}$ or ${\sf Z}_{3}$.
+Further, if (ii) holds, $\overline{K}={\sf Z}_{2}^{3}{:}{\sf Z}_{p}$ with $p=3$ or $7$, then ${\sf Z}_{2}^{3}$ is a minimal normal subgroup of $\overline{K}$, and so $\overline{F_{2}}=F_{2}$ as a normal subgroup of $\overline{K}$ must be identity.
 Consequently, we have $F=F_{p}$, ${\sf Z}_{2}^{3}{:}{\sf Z}_{p}=K/F_{p}\lesssim{\sf Out}(F_{p})$, which does not hold for the $p$-groups appearing in Lemma 4.3.
-Therefore, we have that (1) holds, and since $\overline{K}={\bf O}_{2}(\overline{K}){:}H$ is known, $F_{2}=\overline{F_{2}}$ as a normal 2-subgroup of $\overline{K}$ is also known.
-We then analyze all the possibilities.
+Therefore, $\overline{K}={\bf O}_{2}(\overline{K}){:}H$ is given in (i). We then analyze all the possibilities for $F_{2}=\overline{F_{2}}$, which a normal 2-subgroup of $\overline{K}$.
 Case 1. $F_{2}={\bf O}_{2}(\overline{K})$. In this case, we have $H={\rm S}_{3}$ since $F_{2}<K_{2}$. Then
@@
 Consequently, Case 1 gives the groups listed in Table 1.
-Table 1. Case: $F_{2}={\bf O}_{2}(\overline{K})$
+Table 1. Case 1: $F_{2}={\bf O}_{2}(\overline{K})$
 case $F_{2}$ ${\sf Z}_{2}^{2}$ ${\sf Z}_{2}^{3}$ ${\rm Q}_{8}$ ${\sf Z}_{4}\circ{\rm Q}_{8}$
@@
 (1.2) ${\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}}$ ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}})$ ${\rm Q}_{8}{:}{\rm D}_{2.3^{\ell+1}}$ ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell+1}})$
-(1.3) ${\sf Z}_{3^{\ell}}\times{\rm S}_{4},$ ${\sf Z}_{2.3^{\ell}}\times{\rm S}_{4}$ ${\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\rm S}_{3})$ ${\sf Z}_{4.3^{\ell}}\circ({\rm Q}_{8}{:}{\rm S}_{3})$
-(1.4) ${\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$, ${\sf Z}_{6}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{3}\times({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{12}\circ({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$
+(1.3) ${\sf Z}_{3^{\ell}}\times{\rm S}_{4}$ ${\sf Z}_{2.3^{\ell}}\times{\rm S}_{4}$ ${\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\rm S}_{3})$ ${\sf Z}_{4.3^{\ell}}\circ({\rm Q}_{8}{:}{\rm S}_{3})$
+(1.4) ${\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{6}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{3}\times({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{12}\circ({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$
 (1.5) $({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$ $({\sf Z}_{2.3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$ $({\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\rm S}_{3})){:}{\sf Z}_{2}$ $({\sf Z}_{4.3^{\ell}}\circ({\rm Q}_{8}{:}{\rm S}_{3})){:}{\sf Z}_{2}$
@@
 Consequently, Case 2 gives the groups listed in Table 2.
-Table 2. Case: $F_{2}<{\bf O}_{2}(\overline{K})$
+Table 2. Case 2: $F_{2}<{\bf O}_{2}(\overline{K})$
 case $(F_{2},{\bf O}_{2}(\overline{K}))$ $({\sf Z}_{2}^{2},{\sf Z}_{2}^{3})$ $({\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8})$
@@
 $\Box$
-We are now ready to prove the first main theorem of this paper.
+Now we are ready to prove Theorem 1.1.
 Proof of Theorem 1.1:
@@
 6. Proof of Theorem 1.2
-In this section we focus on the automorphism groups of regular maps.
-The following lemma shows that the class of groups with regular triples is closed under taking quotients.
+In this section we focus on the arc-transitive automorphism groups of maps. The following lemma shows that the class of such groups is closed under taking quotients.
 Lemma 6.1.
-Let $G$ be a finite group which has a regular triple, and let $\overline{G}$ be a non-identity factor group of $G$.
-Then either $\overline{G}$ has a regular triple, or $\overline{G}$ is ${\sf Z}_{2}$ or a dihedral group.
+Let $G$ be a finite group, with $\overline{G}$ a non-identity factor group. Then:
+  • (1)
+    If $G$ has a regular triple (resp.
+reversing triple), then either $\overline{G}$ has a regular triple (resp.
+reversing triple), or $\overline{G}$ is ${\sf Z}_{2}$ or a dihedral group.
 In particular, $\overline{G}$ is of even order.
+  • (2)
+    If $G$ has a rotary pair, then either $\overline{G}$ has a rotary pair, or $\overline{G}$ is cyclic. In particular, if $\overline{G}$ is of odd order, then it is cyclic.
 Proof.
-Let $(x,y,z)$ be a regular triple for $G$, and let $\overline{G}=G/N$ for a proper normal subgroup $N\lhd G$.
-If $\overline{x},\overline{y},\overline{z}$ are all involutions of $\overline{G}$ with $\overline{x}\neq\overline{z}$, then they form a regular triple for $\overline{G}$, since $\overline{G}=\langle\overline{x},\overline{y},\overline{z}\rangle$ and $\overline{x}\,\overline{z}=\overline{z}\,\overline{x}$; otherwise, either some of $\overline{x},\overline{y},\overline{z}$ equals the identity-element in $\overline{G}$, or $\overline{x}=\overline{z}$, while in both cases, $\overline{G}$ is generated by one or two involutions, namely, $G$ is ${\sf Z}_{2}$ or dihedral, respectively.
-$\Box$
+Let $(x,y,z)$ be a regular triple for $G$.
+If $\overline{x},\overline{y},\overline{z}\in\overline{G}$ are all involutions, and $\overline{x}\neq\overline{z}$, then they form a regular triple for $\overline{G}$, since $\overline{G}=\langle\overline{x},\overline{y},\overline{z}\rangle$ and $\overline{x}\,\overline{z}=\overline{z}\,\overline{x}$.
+Otherwise, either some of $\overline{x},\overline{y},\overline{z}$ equals the identity-element in $\overline{G}$, or $\overline{x}=\overline{z}$.
+In both cases, $\overline{G}$ is generated by one or two involutions and then it is ${\sf Z}_{2}$ or dihedral.
+It is similar to prove the remaining part of this lemma. $\Box$
 Lemma 6.2.
@@
 $(\pm u\lambda)^{\pm\tau_{\lambda}}=\pm u(\lambda^{\pm\tau_{\lambda}})=\mp u\lambda,\ (\pm u\lambda)^{\pm\lambda\tau_{\lambda}}=\pm u(\lambda^{\pm\lambda\tau_{\lambda}})=\mp u\lambda,$
-and
 $(\pm u\lambda)^{\pm\tau_{\lambda^{\prime}}}=\pm u(\lambda^{\pm\tau_{\lambda^{\prime}}})=\mp u\lambda^{\prime\prime},\ (\pm u\lambda)^{\pm\lambda^{\prime}\tau_{\lambda^{\prime}}}=\pm u(\lambda^{\pm\lambda^{\prime}\tau_{\lambda^{\prime}}})=\pm u\lambda^{\prime\prime}.$
@@
 Let $H=\langle u,i,j,k\rangle\leqslant K$.
-Then $H\cong{\sf Z}_{4}\circ{\rm Q}_{8}$ is normal in $K$, and $K/H\cong{\rm S}_{3}$.
+Then $H\cong{\sf Z}_{4}\circ{\rm Q}_{8}$ is normal in $K$ and $K/H\cong{\rm S}_{3}$.
 Further, the involutions contained in $H$ are: $\{-1\}\cup\{\pm u\lambda\}$.
@@
 $\Box$
+Lemma 6.3.
+Let $G=({\sf Z}_{3^{\ell}}\times{\sf Z}_{3}){:}{\sf Z}_{2}=\langle a,b\rangle{:}\langle c\rangle$, with $a^{c}=a^{-1}$, $b^{c}=b^{-1}$. Then $G$ has no rotary pair.
+Proof.
+Let $(\alpha,z)$ be a rotary pair for $G$.
+Since $G$ is not dihedral, $|\alpha|\neq 2$, so $\alpha\in\langle a,b\rangle$ and $\langle\alpha\rangle$ is normal in $G$.
+It follows that $\overline{G}=G/\langle\alpha\rangle=\langle\overline{z}\rangle={\sf Z}_{2}$, not possible.
+$\Box$
 Now we are ready to prove Theorem 1.2.
 Proof of Theorem 1.2:
-Let $\mathcal{M}$} be a regular map with square-free Euler characteristic, and let $G={\sf Aut}{\mathcal{M}}$.
-By Lemma 2.1, each Sylow subgroup of $G$ has a cyclic or dihedral subgroup of prime index, and so G=(A:B):KG=(A{:}B){:}K as described in Theorem 1.1.
-Note that $G$ has a regular triple.
-Then it is of even order, and so $K\neq 1$.
-Further, if $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}G_{7}$ or $K=G_{\{2,3,7\}}={\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3})$, then $K$ has $G_{2}={\sf Z}_{2}^{3}$ as a minimal normal subgroup.
-By checking the $p$-groups appearing in Lemma 4.3, meanwhile noticing that ${\sf Z}_{2}^{3}$ cannot be embedded into ${\rm GL}(2,p)$ for all primes $p$ (as a faithful representation of ${\sf Z}_{2}^{3}$ on $\mathbb{F}_{p}$ is of degree at least $3$), we deduce that $G_{2}={\sf Z}_{2}^{3}$ centralizes each Sylow subgroup of A:BA{:}B.
-Hence $G_{2}\lhd G$, and so $G$ has an odd factor group $G/G_{2}$, leading to a contradiction by Lemma 6.1.
-Therefore, $K=G_{2}$ or $G_{\{2,3\}}$.
-Suppose $K=G_{2}$.
-Since K=G/(A:B)K=G/(A{:}B) is a factor group of $G$, by Lemma 6.1, either $K$ has a regular triple and then it is one of the groups given in Proposition 4.2; or $K$ is ${\sf Z}_{2}$ or a dihedral group.
-In conclusion, $K$ is ${\sf Z}_{2}$, ${\rm D}_{2^{\ell+1}}$, ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$ or ${\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}$, where $\ell\geqslant 1$.
-Suppose $K=G_{\{2,3\}}$, which is one of the groups given in Lemma 5.6 (ii) - (iii).
-Since K=G/(A:B)K=G/(A{:}B) is a factor group of $G$, by Lemma 6.1, either $K$ has a regular triple, or $K$ is a dihedral group.
-Further, note that a factor group of a dihedral group is also dihedral (or ${\sf Z}_{2}$).
-Thus, $K$ has no odd factor groups, and has no factor groups isomorphic to ${\rm Q}_{8}{:}{\rm S}_{3}$ or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ either, because by Lemma 6.2 neither ${\rm Q}_{8}{:}{\rm S}_{3}$ nor ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ has a regular triple, and meanwhile, neither of them is a dihedral group.
-Then, by checking the groups appearing in Lemma 5.6 (ii) - (iii), we find that the only candidates for $K=G_{\{2,3\}}$ are the following groups in Table 1: $K=X$ or $X\times{\sf Z}_{2}$, where $X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\};$ and the two groups in Table 2: $K={\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$.
-$\Box$
+Let $\mathcal{M}$} be a map with square-free Euler characteristic, and let $G\leqslant{\sf Aut}{\mathcal{M}}$ be a solvable arc-transitive automorphism group of $\mathcal{M}$}.
+By Lemmas 2.1 and 2.3, the group $G$ satisfies Hypothesis 2.2, so that
+G=(A:B):KG=(A{:}B){:}K
+as described in Theorem 1.1.
+Recall that $\mathcal{M}$} as a $G$-arc-transitive map satisfies one of the following properties: (i) $G$-regular; (ii) $G$-vertex-reversing; (iii) $G$-vertex-rotary; respectively in each case, the group $G$ has: (i) a regular triple; (ii) a reversing triple; (iii) a rotary pair.
+In particular, $G$ is of even order, and so $K\neq 1$.
+First, suppose $K=G_{2}$.
+By Lemma 6.1, either $K$ has a regular triple/reversing triple/rotary pair as $G$ does, and then $K$ is one of the 2-groups given in Proposition 4.2; or $K$ is ${\sf Z}_{2}$, dihedral ${\rm D}_{2^{\ell}}$ or cyclic ${\sf Z}_{2^{\ell}}$ with $\ell\geqslant 2$, while ${\rm D}_{2^{\ell}}$ and ${\sf Z}_{2^{\ell}}$ already appear in Proposition 4.2.
+Then, suppose $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}G_{7}$ or $K=G_{\{2,3,7\}}={\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3})$.
+In this case, $G$ has an odd factor group $G_{7}$ or $G_{7}{:}G_{3}$, respectively.
+By Lemma 6.1, G has neither regular triples nor reversing triples, so $G$ has a rotary pair, with each odd factor group cyclic.
+It follows that $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}{\sf Z}_{7^{\ell}}$.
+At last, suppose that $K=G_{\{2,3\}}$ is one of the groups given in 5.6 (2) - (3).
+(a) If $\mathcal{M}$} is $G$-vertex-reversing, then $G$ has a reversing triple, with each factor group of even order by Lemma 6.1.
+Hence, $K$ is one of the groups in Tables 1 - 2, excluding those corresponding to cases (1.3), (1.4), (1.7), (2.1) - (2.3) as they have an odd factor group ${\sf Z}_{3}$ or ${\sf Z}_{3^{\ell}}$.
+(b) If $\mathcal{M}$} is $G$-regular, then $G$ has a regular triple.
+In particular, $G$ has a reversing triple, and so $K$ is one of the groups in (a).
+Furthermore, by Lemma 6.2 the groups which have ${\rm Q}_{8}{:}{\rm S}_{3}$ or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ as a factor group should be excluded.
+It follows that $K$ is one of: $X$, $X\times{\sf Z}_{2}$, ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$, where
+$X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\}.$
+(c) If $\mathcal{M}$} is $G$-vertex-rotary, then $G$ has a rotary pair, with each odd factor group cyclic by Lemma 6.1.
+So, either $K=X{:}K_{3}=X{:}{\sf Z}_{3^{\ell}}$ with $X\in\{{\sf Z}_{2}^{2},\,{\sf Z}_{2}^{3},\,{\rm Q}_{8},\,{\sf Z}_{4}\circ{\rm Q}_{8}\};$ or $K$ is one of the groups in Tables 1 - 2.
+For the former, $X\neq{\rm Q}_{8}$, as the only involution of ${\rm Q}_{8}{:}{\sf Z}_{3^{\ell}}$ lies in the center and so it does not have a rotary pair.
+For the latter, by Lemma 6.3 the groups corresponding to cases (1.5) and (1.6), which have $({\sf Z}_{3^{\ell}}\times{\sf Z}_{3}){:}{\sf Z}_{2}$ as a factor group, should be excluded.
+In conclusion, $K$ is one of the groups stated in Theorem 1.2. $\Box$
 Declarations

### v3

@@
 The project was partially supported by the NNSF of China (11931005).
-December 1, 2025
+December 11, 2025
 1. Introduction
@@
 As usual, for a group $G$ and a subset $\pi$ of prime divisors of $|G|$, denote by $G_{\pi}$ a Hall $\pi$-subgroup of $G$, which is a Sylow subgroup if $\pi$ contains a single prime $p$, and denoted by $G_{p}$.
-For a positive integer $n$, denote by ${\sf Z}_{n}$ a cyclic group of order $n$, by ${\rm D}_{2n}$ or ${\rm Q}_{2n}$ a dihedral or generalized quaternion group of order $2n$, and by ${\rm A}_{n}$ or ${\rm S}_{n}$ an alternating or symmetric group of degree $n$, respectively.
+For a positive integer $n$, denote by ${\sf Z}_{n}$ the cyclic group of order $n$, by ${\rm D}_{2n}$ or ${\rm Q}_{2n}$ the dihedral or generalized quaternion group of order $2n$, and by ${\rm A}_{n}$ or ${\rm S}_{n}$ the alternating or symmetric group of degree $n$, respectively.
 A finite group $G$ is a semidirect product of $A$ by $B$ if $A\lhd G$ and $B<G$ are such that $G/A\cong B$, denoted by G=A:BG=A{:}B.
+Denote by $A\circ_{C}B$ the central product of $A$ and $B$ (find the definition in Preliminaries).
 Our first main result is the following theorem.
@@
 Let $G$ be a finite solvable group such that each Sylow subgroup of $G$ has a cyclic or dihedral subgroup of prime index. Then :
-(1) The Sylow subgroups of $G$ are characterized in Proposition 4.1 (for $G_{2}$) and Lemma 4.3 (for $G_{p}$, $p$ odd prime).
-(2) G=(A:B):KG=(A{:}B){:}K, where $\gcd(|A{:}B|,|K|)=1$, such that
+(1) The Sylow subgroups $G_{2}$ and $G_{p}\ (p>2)$ of $G$ are respectively given in Proposition 4.1 and Lemma 4.3.
+(2) Let $H\leqslant G$ be the largest normal Hall subgroup of odd order. Then, H=A:BH=A{:}B, where $A$ is abelian, $B$ is nilpotent and $A\cap{\bf Z}(H)=1$.
+(3) G=H:KG=H{:}K, where $K$ satisfies one of the following :
   • (i)
-    AA is abelian, $B$ is nilpotent, $A\cap{\bf Z}(A{:}B)=1;$
+    $K=1$ or $G_{2}\,;$
   • (ii)
-    $K$ is one of the following groups:
-    • (a)
-      $K=G_{2}$ ($K$ might be $1$ in case of $G$ of odd order);
-    • (b)
-      $K=G_{\{2,3\}}$, explicitly given in Lemma 5.6 (2) - (3);
-    • (c)
-      $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}G_{7}$, or $K=G_{\{2,3,7\}}={\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3})$.
+    $K=G_{\{2,3\}}$ or $G_{\{2,7\}}$ is one of the groups :
+    ${\sf Z}_{2}^{2}{:}G_{3},\ {\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}G_{3}),\ {\rm Q}_{8}{:}G_{3},\ {\sf Z}_{4}\circ_{{\sf Z}_{2}}({\rm Q}_{8}{:}G_{3}),\,{\sf Z}_{2}^{3}{:}G_{7}\,;$
+  • (iii)
+    $K=G_{\{2,3\}}$ is one of the groups :
+    $X$, ${\sf Z}_{2}\times X$, $Y$, ${\sf Z}_{4}\circ_{{\sf Z}_{2}}Y$, $W$,
+    where $X,Y,W$ are given in Table 1, $X$ is homomorphic to ${\sf Z}_{2}^{2}{:}{\rm S}_{3}\cong{\rm S}_{4}$, $Y$ is homomorphic to ${\rm Q}_{8}{:}{\rm S}_{3}\cong{\rm GL}_{2}(3)$, and $W$ is homomorphic to ${\sf Z}_{2}^{2}{:}{\sf Z}_{3}\cong{\rm A}_{4}$, ${\rm S}_{4}$, ${\sf Z}_{4}\circ{\rm SL}_{2}(3)$ or ${\sf Z}_{4}\circ{\rm GL}_{2}(3);$
+    Table 1. $X,Y,W$, $\ell\geqslant 1$
+    Row $X$ $Y$ $W$
+    11 ${\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}}$ ${\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}}$ ${\rm D}_{2.3^{\ell}}\times{\rm A}_{4}$, $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\sf Z}_{3})$
+    22 ${\sf Z}_{3^{\ell}}\times{\rm S}_{4}$ ${\sf Z}_{3^{\ell}}\times{\rm GL}_{2}(3)$ $({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})\times{\rm D}_{6}$, $({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$
+    33 $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\sf Z}_{3}$ $({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})\times{\sf Z}_{3}$ ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$, $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\rm S}_{3})$
+    44 $({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$ $({\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\sf Z}_{3})){:}{\sf Z}_{2}$ $({\sf Z}_{2}^{2}{:}{{\rm D}_{2.3^{\ell}}})\times{\rm D}_{6}$, $({\rm Q}_{8}{:}{{\rm D}_{2.3^{\ell}}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$
+    55 $(({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})\times{\sf Z}_{3}){:}{\sf Z}_{2}$ $(({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})\times{\sf Z}_{3}){:}{\sf Z}_{2}$ $({\sf Z}_{2}^{2}\times{\rm D}_{2.3^{\ell+1}}){:}{\sf Z}_{3}$, $({\rm Q}_{8}\circ({\sf Z}_{3^{\ell+1}}{:}{\sf Z}_{4})){:}{\sf Z}_{3}$
+    66 ${\sf Z}_{2}^{2}{:}({\sf Z}_{3^{\ell+1}}{:}{\sf Z}_{6})$ ${\rm Q}_{8}{:}({\sf Z}_{3^{\ell+1}}{:}{\sf Z}_{6})$
+  • (iv)
+    $K=G_{\{2,3,7\}}={\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3}).$
 The next theorem shows that arc-transitive automorphism groups of maps with square-free Euler characteristic are more restricted.
@@
 Let $\mathcal{M}$} be a map with square-free Euler characteristic, and let $G\leqslant{\sf Aut}{\mathcal{M}}$ be solvable and arc-transitive on $\mathcal{M}$}.
-Then G=(A:B):KG=(A{:}B){:}K such that $A$ is abelian with $A\cap{\bf Z}(A{:}B)=1$, $B$ is nilpotent, and either $K=G_{2}$ is ${\sf Z}_{2}$ or one of the groups given in Proposition 4.2, or one of the following holds:
+Then, G=(A:B):KG=(A{:}B){:}K, where $A$ is abelian, $B$ is nilpotent, $A\cap{\bf Z}(A{:}B)=1$, and one of the following holds :
   • (1)
-    $\mathcal{M}$} is $G$-regular, $K=G_{\{2,3\}}$ is one of $X$, $X\times{\sf Z}_{2}$, ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$, where $X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\};$
+    $\mathcal{M}$} is $G$-arc-transitive, $K=G_{2}$ is ${\sf Z}_{2}$ or one of the groups given in Proposition 4.2.
   • (2)
-    $\mathcal{M}$} is $G$-vertex-reversing, $K=G_{\{2,3\}}$ is one of the groups in Table 1 and Table 2, corresponding to cases (1.1), (1.2), (1.5), (1.6), (2.4), (2.5);
+    $\mathcal{M}$} is $G$-regular, $K=G_{\{2,3\}}$ is one of the groups :
+    $X$, $X\times{\sf Z}_{2}$, ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$, $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$,
+    where $X\in\{{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,(({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})\times{\sf Z}_{3}){:}{\sf Z}_{2}\}.$
   • (3)
-    $\mathcal{M}$} is $G$-vertex-rotary, $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}{\sf Z}_{7^{\ell}}$, or $K=G_{\{2,3\}}$ is one of ${\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}}$, ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})$, ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})$, or one of the groups in Table 1 and Table 2, excluding cases (1.5), (1.6).
-The proofs of the above two results are given in Section 4 - 5 and Section 6.
-Before that, we construct several interesting examples of regular maps with square-free Euler characteristic in Section 3, from which the following result follows.
+    $\mathcal{M}$} is $G$-vertex-reversing, $K=G_{\{2,3\}}$ is either one of the groups given in (2), or one of the groups :
+    $Y$, $Y\circ_{{\sf Z}_{2}}{\sf Z}_{4}$, $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\rm S}_{3})$, $({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$,
+    where $Y\in\{{\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}},\,({\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\sf Z}_{3})){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\}.$
+  • (4)
+    $\mathcal{M}$} is $G$-vertex-rotary, and either $K=G_{\{2,3\}}$ or $G_{\{2,7\}}$ is one of the groups :
+    ${\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}}$,  ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})$,  ${\sf Z}_{4}\circ_{{\sf Z}_{2}}({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})$,  ${\sf Z}_{2}^{3}{:}{\sf Z}_{7^{\ell}}$ ;
+    or $K=G_{\{2,3\}}$ is one of the groups given in Theorem 1.1 (iii), excluding the following four :
+    $({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$,  $({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$,  $({\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\sf Z}_{3})){:}{\sf Z}_{2}$,  $({\sf Z}_{3}\times({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$.
+The proofs of the above two results are respectively given in Section 4 - 5 and Section 6.
+We remark that, Theorem 1.2 does not guarantee the existence of arc-transitive maps $\mathcal{M}$} corresponding to each of these groups.
+The existence and construction of such maps is the focus of the next paper in this series.
+Here, we construct several interesting examples of regular maps with square-free Euler characteristic in Section 3, from which the following result follows.
 Theorem 1.3.
-There are infinite families of regular maps $\mathcal{M}$} with underlying graphs ${\bf C}_{n}^{(n)}$ or ${\bf C}_{n}\square{\bf C}_{n}$, and with Euler characteristic $\chi({\mathcal{M}})=-n(n-3)$ or n(n2)-n(n-2), respectively.
-In particular, there are infinitely many of them with $\chi({\mathcal{M}})$ square-free.
+There are infinite families of regular maps which have square-free Euler characteristic of form n(n2)-n(n-2) or n(n3)-n(n-3), with $n$ an positive integer.
 2. Preliminaries
-Let $G$ be a group.
-Denote by $\Phi(G)$ and $G^{\prime}$ the Frattini subgroup and the commutator subgroup of $G$.
-If $G$ is a $p$-group, then $^{\prime}\leqslant\Phi(G), G/G$$G^{\prime}$ is abelian, and $G/\Phi(G)$ is elementary abelian.
-For a subgroup $H\leqslant G$, denote by ${\bf N}_{G}(H)$ and ${\bf C}_{G}(H)$ the normalizer and the centralizer of $H$ in $G$, respectively.
+Let $G$ be a finite group, and let $H\leqslant G$ be a subgroup of $G$.
+Denote by $G^{\prime}$ the commutator subgroup of $G$, and by ${\bf N}_{G}(H)$ and ${\bf C}_{G}(H)$ the normalizer and the centralizer of $H$ in $G$, respectively.
+If $G$ is solvable, it is well-known that, for any subset $\pi$ of prime divisors of $|G|$, there exist Hall $\pi$-subgroups of $G$ with any two of them conjugate in $G$, and any $\pi$-subgroup is contained in some Hall $\pi$-subgroup.
+Let $A,B$ be finite groups whose centers ${\bf Z}(A),{\bf Z}(B)$ have isomorphic subgroups $C_{1},C_{2}$, respectively.
+Let $\varphi$ be an isomorphism from $C_{1}$ to $C_{2}$, and let $C=\{(c,c^{\varphi})\mid c\in C_{1}\}$.
+Then $C$ is a normal subgroup of $A\times B$.
+The factor group $(A\times B)/C$ is called a central product of $A$ and $B$, denoted by $A\circ_{C}B$, and sometimes simply by $A\circ B$ if $C$ is equal to ${\bf Z}(A)$ or ${\bf Z}(B)$.
 Let ${\it\Gamma},\Sigma$ be two graphs and $\lambda$ be a positive integer.
@@
 Denote by ${\it\Gamma}\square\Sigma$ the Cartesian product of $\it\Gamma$} and $\Sigma$, which has vertex set $V({\it\Gamma})\times V(\Sigma)$, where $(x_{1},y_{1})$ is adjacent to $(x_{2},y_{2})$ if and only if $x_{1}=y_{1}$ and $x_{2}$ is adjacent to $y_{2}$, or $x_{1}$ is adjacent to $y_{1}$ and $x_{2}=y_{2}$.
+It is known (refer to [13]) that an arc-transitive (multi-) graph $\it\Gamma$} can be expressed as a coset graph :
+${\it\Gamma}={\sf Cos}(G,H,J),\ \mbox{or}\ {\sf Cos}(G,H,HgH)\ \mbox{in the simple graph case},$
+which has vertex set $[G:H]$ and edge set $[G:J]$, such that
+    $Hg_{1}$ is incident with $Jg_{0}$ if and only if $g_{1}g_{0}^{-1}\in HJ$; or equivalently,
+    $Hg_{1}$, $Hg_{2}$ are adjacent if and only if $g_{2}g_{1}^{-1}\in HgH$ in the simple graph case.
+Further, $\it\Gamma$} is of valency $|H:H\cap H^{g}|$, where $g\in J\setminus(H\cap J)$, and $\it\Gamma$} is connected if and only if $G=\langle H,J\rangle$.
 A map ${\mathcal{M}}=(V,E,F)$ is a $2$-cell embedding of a graph into a closed surface, with vertex set $V$, edge set $E$ and face set $F$.
@@
 Let $G\leqslant{\sf Aut}({\mathcal{M}})$, and let $(\alpha,e,f)$ be a flag.
-It is well-known that each of the stabilizers $G_{\alpha}$ and $G_{f}$ is cyclic or dihedral, and meanwhile, $G_{e}=1$, ${\sf Z}_{2}$ or ${\rm D}_{4}$.
+It is well-known that, each of the stabilizers $G_{\alpha}$ and $G_{f}$ is cyclic or dihedral, and meanwhile, $G_{e}=1$, ${\sf Z}_{2}$ or ${\sf Z}_{2}^{2}$.
 A map $\mathcal{M}$} is called a $G$-arc-transitive map, if $G$ acts transitively on arcs (directed edges) of $\mathcal{M}$}.
 Arc-transitive maps are divided into five types in [9, 16], for which recent papers [12, 13] provide a new description: letting $\mathcal{M}$} be a $G$-arc-transitive map, then one of the following holds :
-(1) $\mathcal{M}$} is a $G$-regular map (type 1), which is determined a regular triple $(x,y,z)$ for $G$ satisfying:
-$\mbox{$G=\langle x,y,z\rangle$, $|x|=|y|=|z|=2$ and $\langle x,z\rangle\cong{\sf Z}_{2}^{2}$}.$
-In this case, we have:
-$G={\sf Aut}({\mathcal{M}}),\,G_{\alpha}=\langle x,y\rangle,\,G_{e}=\langle x,z\rangle,\,G_{f}=\langle y,z\rangle,$
-and the underlying graph of $\mathcal{M}$} can be expressed as a coset graph ${\sf Cos}(G,H,HzH)$, where $H=\langle x,y\rangle$ (refer to [12].)
+(1) $\mathcal{M}$} is a $G$-regular map (type 1), which is determined by a regular triple $(x,y,z)$ for $G$ satisfying :
+$\mbox{$G=\langle x,y,z\rangle$, $|x|=|y|=|z|=2$ and $\langle x,z\rangle={\sf Z}_{2}^{2}$}.$
+In this case, $G={\sf Aut}({\mathcal{M}})$, the stabilizers $G_{\alpha}=\langle x,y\rangle,\,G_{e}=\langle x,z\rangle,\,G_{f}=\langle y,z\rangle,$ and the underlying graph $\it\Gamma$} is expressed as the coset graph ${\sf Cos}(G,H,HJ)$, where $H=\langle x,y\rangle$, $J=\langle x,z\rangle$, so $\it\Gamma$} is connected, of valency $|H:H\cap H^{z}|$.
 (2) $\mathcal{M}$} is a $G$-vertex-reversing map (type $2^{\ast}$ or $2^{P}$), which is determined by a reversing triple $(x,y,z)$ for $G$ satisfying : $G=\langle x,y,z\rangle$ and $|x|=|y|=|z|=2$.
@@
 The following lemma established in [11] characterizes Sylow subgroups of ${\sf Aut}({\mathcal{M}})$ in the case where $\mathcal{M}$} is with square-free characteristic.
-We note that if $\chi({\mathcal{M}})$ is square-free, then of course so is $\gcd(\chi({\mathcal{M}}),|{\sf Aut}{\mathcal{M}}|)$.
+Note that, if $\chi({\mathcal{M}})$ is square-free, then of course so is $\gcd(\chi({\mathcal{M}}),|{\sf Aut}{\mathcal{M}}|)$.
 Lemma 2.1.
@@
 3. Examples of Square-free Euler Characteristic Regular Maps
-We construct several infinite families of regular maps with square-free Euler characteristic in this section. The constructions are based on the following groups:
+We construct three infinite families of regular maps with square-free Euler characteristic in this section. The constructions are based on the following groups:
 $\begin{array}[]{lllll}X&=&(\langle a,s\rangle\times\langle b,t\rangle){:}\langle\sigma\rangle\cong{\rm D}_{2n}\wr{\rm S}_{2}&&\\ &=&\langle a,b\rangle{:}\langle s,\sigma\rangle\cong{\sf Z}_{n}^{2}{:}{\rm D}_{8},&&(\ast)\end{array}$
 where $\langle a,s\rangle=\langle a\rangle{:}\langle s\rangle={\rm D}_{2n}$, $\langle b,t\rangle=\langle b\rangle{:}\langle t\rangle={\rm D}_{2n}$, $(a,s)^{\sigma}=(b,t)$ and $|\sigma|=2$.
-The first family is given now.
+The first family is given below.
 Construction 3.1.
-Let $G=\langle a,s\rangle\times\langle b,t\rangle={\rm D}_{2n}\times{\rm D}_{2n}\leqslant X$ as given in ($\ast$), with $n$ odd. Let
+Let $X$ be the group given in ($\ast$), and let $G=\langle a,s\rangle\times\langle b,t\rangle={\rm D}_{2n}\times{\rm D}_{2n}\leqslant X$. Assume that $n$ is odd. Define $x,y,z$ as follows :
 x=s,y=abst,z=st.x=s,\ y=abst,\ z=st.
-Then $(x,y,z)$ is a triple of involutions with $xz=zx$. Further,
-$(yx)^{2}=(abt)^{2}=a^{2}$ and $(yxz)^{2}=(abs)^{2}=b^{2}$.
-As $n$ is odd, $\langle a\rangle=\langle a^{2}\rangle$ and $\langle b\rangle=\langle b^{2}\rangle$, so that $G=\langle x,y,z\rangle$. Thus, $(x,y,z)$ is a regular triple for $G$.
+Then, $(x,y,z)$ is a triple of involutions such that $xz=zx$. Note that,
+$(yx)^{2}=(abt)^{2}=a^{2}$, $(yxz)^{2}=(abs)^{2}=b^{2}$,
+and meanwhile, $\langle a\rangle=\langle a^{2}\rangle$, $\langle b\rangle=\langle b^{2}\rangle$ since $n$ is odd.
+Thus, $a,b,s$ and $t=xz$ are all contained in $\langle x,y,z\rangle,$ so $G=\langle x,y,z\rangle$, and $(x,y,z)$ is a regular triple for $G$.
 Lemma 3.2.
@@
 The vertex stabilizer $H$ is
-$H=\langle x,y\rangle=\langle s,abst\rangle=\langle abt\rangle{:}\langle s\rangle=\langle a,s\rangle\times\langle bt\rangle={\rm D}_{4n}.$
-Meanwhile, the face stabilizer is
+$H=\langle x,y\rangle=\langle s,abst\rangle=\langle abt\rangle{:}\langle s\rangle=\langle a,s\rangle\times\langle bt\rangle={\rm D}_{4n},$
+and the face stabilizer $L$ is
 $\langle y,z\rangle=\langle abst,st\rangle=\langle ab\rangle{:}\langle st\rangle={\rm D}_{2n}.$
-Thus the map ${\mathcal{M}}_{n}$ has valency $2n$ and face length $n$.
-Further, $H\cap H^{z}=\langle a,s\rangle\cong{\rm D}_{2n}$ is of index $2$ in $H$, so the underlying graph ${\it\Gamma}={\sf Cos}(G,H,HzH)\cong{\bf C}_{n}^{(n)}$.
-The Euler characteristic
-$\chi({\mathcal{M}}_{n})=|G|\left({1\over|\langle x,y\rangle|}-{1\over 4}+{1\over|\langle y,z\rangle|}\right)=4n^{2}\left({1\over 4n}-{1\over 4}+{1\over 2n}\right)=-n(n-3).$
+Thus, $(|V|,|E|,|F|)=(\frac{|G|}{|H|},\frac{|G|}{4},\frac{|G|}{|L|})=(n,n^{2},2n)$, and the Euler characteristic
+$\chi({\mathcal{M}}_{n})=n-n^{2}+2n=-n(n-3)$.
 Let $n=4x+1$.
@@
 Ricci in 1933 (which is improved by $\rm P.Erd\ddot{o}s$} in [8]), there are infinitely many positive integers $x$ for which f(x)=(4x+1)(2x1)f(x)=(4x+1)(2x-1) is square-free, so there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free.
+Here are some examples :
+$\begin{array}[]{|c|c|c|c|c|c|c|c|}\hline\cr n&5&13&17&29&37&41&\cdots\\ \hline\cr-n(n-3)&-2.5&-2.5.13&-2.7.17&-2.13.29&-2.17.37&-2.19.41$$\cdots\\ \hline\cr\end{array}$
+At last, the valency of ${\mathcal{M}}_{n}$ is $\frac{2|E|}{|V|}=2n$, and the face length is $\frac{2|E|}{|F|}=n$.
+Note that, $H\cap H^{z}=\langle a,s\rangle\cong{\rm D}_{2n}$ is of index $2$ in $H$.
+Then, the underlying graph ${\it\Gamma}={\sf Cos}(G,H,HJ)$ is connected of valency 2, so it is the multiple cycle ${\bf C}_{n}^{(n)}$.
 $\Box$
-Here is some data such that n(n3)-n(n-3) is square-free, with $n$ odd.
-$\begin{array}[]{c|ccccccc}\hline\cr n&5&13&17&29&37&41&\cdots\\ -n(n-3)&-2.5&-2.5.13&-2.7.17&-2.13.29&-2.17.37&-2.19.41$$\cdots\\ \hline\cr\end{array}$
-Next we construct the second family.
+The second family is given below.
 Construction 3.3.
-Let $G=\langle a,b\rangle{:}\langle st,\sigma\rangle={\sf Z}_{n}^{2}{:}{\sf Z}_{2}^{2}\leqslant X$ as given in ($\ast$). Let
+Let $X$ be the group given in ($\ast$), and let $G=\langle a,b\rangle{:}\langle st,\sigma\rangle={\sf Z}_{n}^{2}{:}{\sf Z}_{2}^{2}\leqslant X$. Define $x,y,z$ as follows :
 $x=\sigma,\ y=ast,\ z=abst.$
-Then $(x,y,z)$ is a triple of involutions with $xz=zx$. Further,
-$yz=b^{-1}$ and $(yz)^{x}=a^{-1}$,
-so that $G=\langle x,y,z\rangle$. Thus, $(x,y,z)$ is a regular triple for $G$.
+Then, $(x,y,z)$ is a triple of involutions such that $xz=zx$. Note that,
+$zy=b$, $(zy)^{x}=b^{\sigma}=a$.
+Thus, $a,b,\sigma$ and $st=a^{-1}y$ are all contained in $\langle x,y,z\rangle$, so $G=\langle x,y,z\rangle$, and $(x,y,z)$ is a regular triple for $G$.
 Lemma 3.4.
@@
 The vertex stabilizer $H$ is
-$H=\langle x,y\rangle=\langle ast\sigma\rangle{:}\langle\sigma\rangle={\rm D}_{4n}.$
-Meanwhile, the face stabilizer is
+$H=\langle x,y\rangle=\langle ast\sigma\rangle{:}\langle\sigma\rangle={\rm D}_{4n},$
+and the face stabilizer $L$ is
 $\langle y,z\rangle=\langle abst,ast\rangle=\langle b\rangle{:}\langle ast\rangle={\rm D}_{2n}.$
-Thus the map ${\mathcal{M}}_{n}$ has valency $2n$ and face length $n$.
-Further, $H\cap H^{z}=\langle a^{-1}b\rangle{:}\langle\sigma\rangle={\rm D}_{2n}$ is of index $2$ in $H$, so the underlying graph ${\it\Gamma}={\sf Cos}(G,H,HzH)\cong{\bf C}_{n}^{(n)}$.
-The Euler characteristic
-$\chi({\mathcal{M}})=|G|\left({1\over|\langle x,y\rangle|}-{1\over 4}+{1\over|\langle y,z\rangle|}\right)=4n^{2}\left({1\over 4n}-{1\over 4}+{1\over 2n}\right)=-n(n-3).$
+Thus, $(|V|,|E|,|F|)=(\frac{|G|}{|H|},\frac{|G|}{4},\frac{|G|}{|L|})=(n,n^{2},2n)$, and the Euler characteristic
+$\chi({\mathcal{M}}_{n})=n-n^{2}+2n=-n(n-3)$.
 By Lemma 3.2, there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free.
-Further, let $n=4x+2$.
-Then according to [8], there are infinitely many positive integers $x$ for which f(x)=(2x+1)(4x1)f(x)=(2x+1)(4x-1) is square-free, so there are infinitely many even values of $n$ such that $\chi({\mathcal{M}}_{n})=-2(2x+1)(4x-1)$ is square-free.
+In a similar way, we can prove that there are infinitely many even values of $n$ with $n=4x+2$, such that $\chi({\mathcal{M}}_{n})=-2(2x+1)(4x-1)$ is square-free.
+Here are some examples :
+$\begin{array}[]{|c|c|c|c|c|c|c|c|}\hline\cr n&2&10&14&22&26&34&\cdots\\ \hline\cr-n(n-3)&2&-2.5.7&-2.7.11&-2.11.19&-2.13.23&-2.17.31$$\cdots\\ \hline\cr\end{array}$
+At last, the valency of ${\mathcal{M}}_{n}$ is $\frac{2|E|}{|V|}=2n$, and the face length is $\frac{2|E|}{|F|}=n$.
+Note that, $H\cap H^{z}=\langle a^{-1}b\rangle{:}\langle\sigma\rangle={\rm D}_{2n}$ is of index $2$ in $H$.
+Then, the underlying graph ${\it\Gamma}={\sf Cos}(G,H,HJ)$ is connected of valency 2, so it is the multiple cycle ${\bf C}_{n}^{(n)}$.
 $\Box$
-Here is some data such that n(n3)-n(n-3) is square-free, with $n$ even.
-$\begin{array}[]{c|ccccccc}\hline\cr n&2&10&14&22&26&34&\cdots\\ -n(n-3)&2&-2.5.7&-2.7.11&-2.11.19&-2.13.23&-2.17.31$$\cdots\\ \hline\cr\end{array}$
 The third family is given below.
 Construction 3.5.
-Let $G=\langle a,b\rangle{:}\langle s,\sigma\rangle={\sf Z}_{n}^{2}{:}{\rm D}_{8}=X$ as given in ($\ast$), with $n$ odd. Let
+Let $G=X=\langle a,b\rangle{:}\langle s,\sigma\rangle={\sf Z}_{n}^{2}{:}{\rm D}_{8}$ be the group given in ($\ast$). Assume that $n$ is odd. Define $x,y,z$ as follows :
 $x=\sigma,\ y=s,\ z=abst.$
-Then $(x,y,z)$ is a triple of involutions with $xz=zx$. Further,
-$(zy)^{2}=(abt)^{2}=a^{2}$ and $((zy)^{2})^{x}=(a^{2})^{x}=b^{2}$.
-As $n$ is odd, $\langle a\rangle=\langle a^{2}\rangle$ and $\langle b\rangle=\langle b^{2}\rangle$, so that $G=\langle x,y,z\rangle$. Thus, $(x,y,z)$ is a regular triple for $G$.
+Then, $(x,y,z)$ is a triple of involutions such that $xz=zx$. Note that,
+$(zy)^{2}=(abt)^{2}=a^{2}$, $((zy)^{2})^{x}=(a^{2})^{x}=b^{2}$,
+and meanwhile, $\langle a\rangle=\langle a^{2}\rangle$, $\langle b\rangle=\langle b^{2}\rangle$ since $n$ is odd.
+Thus, $a,b,s,\sigma$ are all contained in $\langle x,y,z\rangle$, so $G=\langle x,y,z\rangle$, and $(x,y,z)$ is a regular triple for $G$.
 Lemma 3.6.
@@
 The vertex stabilizer $H$ is
-$H=\langle x,y\rangle=\langle\sigma,s\rangle={\rm D}_{8}.$
-Meanwhile, the face stabilizer
+$H=\langle x,y\rangle=\langle\sigma,s\rangle={\rm D}_{8},$
+and the face stabilizer $L$ is
 $\langle y,z\rangle=\langle s,abst\rangle=\langle abt\rangle{:}\langle s\rangle={\rm D}_{4n}.$
-Thus the map ${\mathcal{M}}_{n}$ has valency $4$ and face length $2n$.
-Let ${\it\Gamma}={\sf Cos}(G,H,HzH)$ be the underlying graph of ${\mathcal{M}}_{n}$.
-Then $G=\langle a,b\rangle{:}H$, and so $\langle a,b\rangle$ is regular on the vertex set of $\it\Gamma$}.
-Hence $\it\Gamma$} is a (connected and undirected) Cayley graph ${\sf Cay}(K,S)$ of valency 4, where $K=\langle a,b\rangle={\sf Z}_{n}^{2}$.
-It follows that $S=\{x,x^{-1},y,y^{-1}\}$ with $\langle x,y\rangle=\langle a,b\rangle$, and so ${\it\Gamma}={\bf C}_{n}\square{\bf C}_{n}$.
-The Euler characteristic
-$\chi({\mathcal{M}}_{n})=|G|\left({1\over|\langle x,y\rangle|}-{1\over 4}+{1\over|\langle y,z\rangle|}\right)=8n^{2}\left({1\over 8}-{1\over 4}+{1\over 4n}\right)={-n(n-2).}$
-According to [8], there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free.
+Thus, $(|V|,|E|,|F|)=(\frac{|G|}{|H|},\frac{|G|}{4},\frac{|G|}{|L|})=(n^{2},2n^{2},2n)$, and the Euler characteristic
+$\chi({\mathcal{M}}_{n})=n^{2}-2n^{2}+2n=-n(n-2).$
+It is similar to Lemma 3.2 to prove that, there are infinitely many odd values of $n$ such that $\chi({\mathcal{M}}_{n})$ is square-free.
 Further, by a theorem of Chen [3], there indeed exist infinitely many primes $p$ such that p2p-2 is square-free, then so is p(p2)-p(p-2).
+Here are some examples :
+$\begin{array}[]{|c|c|c|c|c|c|c|c|c|}\hline\cr p&3&5&7&13&17&19&23&\cdots\\ \hline\cr-p(p-2)&-3&-3.5&-5.7&-11.13&-3.5.17&-17.19&-3.7.23$$\cdots\\ \hline\cr\end{array}$
+At last, the valency of ${\mathcal{M}}_{n}$ is $\frac{2|E|}{|V|}=4$, and the face length is $\frac{2|E|}{|F|}=2n$.
+Let ${\it\Gamma}={\sf Cos}(G,H,HJ)$ be the underlying graph of ${\mathcal{M}}_{n}$, and let $K=\langle a,b\rangle={\sf Z}_{n}^{2}\leqslant G$.
+Note that, G=K:HG=K{:}H, so $K$ is regular on the vertex set of $\it\Gamma$}, and $\it\Gamma$} is thus a connected Cayley graph ${\sf Cay}(K,S)$ of valency 4.
+It follows that $S=\{x,x^{-1},y,y^{-1}\}$ with $\langle x,y\rangle=\langle a,b\rangle$, and ${\it\Gamma}={\sf Cay}(K,S)\cong{\bf C}_{n}\square{\bf C}_{n}$.
 $\Box$
-Here is some data such that p(p2)-p(p-2) is square-free, with $p$ prime.
-$\begin{array}[]{c|cccccccc}\hline\cr p&3&5&7&13&17&19&23&\cdots\\ -p(p-2)&-3&-3.5&-5.7&-11.13&-3.5.17&-17.19&-3.7.23$$\cdots\\ \hline\cr\end{array}$
 4. $p$-Groups
 In this section, we study $p$-groups satisfying Hypothesis 2.2, i.e., $p$-groups with a cyclic or dihedral subgroup of index $p$.
-Note that the case for $p=2$ is dealt with in the previous work [10], which is introduced below.
-Let $A,B$ be finite groups whose centers ${\bf Z}(A),{\bf Z}(B)$ have isomorphic subgroups $C_{1},C_{2}$, respectively.
-Let $\varphi$ be an isomorphism from $C_{1}$ to $C_{2}$, and let $C=\{(c,c^{\varphi})\mid c\in C_{1}\}$.
-Then $C$ is a normal subgroup of $A\times B$.
-The factor group $(A\times B)/C$ is called a central product of $A$ and $B$, denoted by $A\circ_{C}B$, and sometimes simply by $A\circ B$ if $C$ is equal to ${\bf Z}(A)$ or ${\bf Z}(B)$.
+Note that, the case for $p=2$ is dealt with in the previous work [10], which is introduced below.
 Let ${\rm Q}_{2^{\ell+1}}=\langle u,v\rangle$ be a generalized quaternion group of order $2^{\ell+1}$, where
@@
 1.1) Let $G$ be a $2$-group which has a cyclic or dihedral subgroup of index $2$.
-Then $G$ is one of the following groups.
+Then one of the following holds.
 Further, either ${\sf Aut}(G)$ is a $2$-group, or $(G,{\sf Aut}(G))$ is one of $({\sf Z}_{2}^{2},{\rm S}_{3}),\,({\sf Z}_{2}^{3},{\rm GL}_{3}(2)),\,({\rm Q}_{8},{\rm S}_{4}),\,({\rm Q}_{8}\circ{\sf Z}_{4},{\rm S}_{4}\times{\sf Z}_{2})$.
@@
   • (1)
-    If $G$ has a reversing triple, then $G={\rm D}_{2^{\ell+1}}$, ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$, ${\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}$ or ${\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4}$, where $\ell\geqslant 1;$ further, if $G$ has a regular triple, then $G\neq{\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4};$
+    If $G$ has a reversing triple, then $G={\rm D}_{2^{\ell+1}}$, ${\rm D}_{2^{\ell+1}}\times{\sf Z}_{2}$, ${\rm D}_{2^{\ell+3}}{:}{\sf Z}_{2}$ or ${\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4}$, where $\ell\geqslant 1;$ further, if $G$ has a regular triple, then $G\neq{\rm Q}_{2^{\ell+2}}\circ{\sf Z}_{4}.$
   • (2)
     If $G$ has a rotary pair, then $G={\sf Z}_{2^{\ell+1}}$, ${\sf Z}_{2^{\ell}}\times{\sf Z}_{2}$, ${\rm D}_{2^{\ell+2}}$ or ${\sf Z}_{2^{\ell+2}}{:}{\sf Z}_{2}$, $\ell\geqslant 1$.
-For the odd primes $p$, we have the following characterization.
+The case for $p>2$ is dealt with in the following lemma.
 Lemma 4.3.
@@
   • (2)
-    $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$ is abelian, and
+    $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}=\langle a,b\rangle$ is abelian, and
     1. (2.1)
-      if $\ell>1$, then ${\sf Aut}(G)$ is solvable, and letting $G=\langle a,b\rangle$, then any Hall $p^{\prime}$-subgroup of ${\sf Aut}(G)$ is conjugate to $({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{p^{\prime}}={\sf Z}_{p-1}\times{\sf Z}_{p-1};$
+      if $\ell>1$, then ${\sf Aut}(G)$ is solvable of order $p^{\ell+1}(p-1)^{2}$, with an abelian Hall $p^{\prime}$-subgroup ${\sf Aut}(G)_{p^{\prime}}=({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{p^{\prime}}={\sf Z}_{p-1}\times{\sf Z}_{p-1};$
     2. (2.2)
-      if $\ell=1$, then ${\sf Aut}(G)={\rm GL}(2,p)$, and further, any subgroup $M\leqslant{\rm GL}(2,p)$ of order coprime to $2p$ is an abelian group.
+      if $\ell=1$, then ${\sf Aut}(G)={\rm GL}(2,p)$ is of order $p(p+1)(p-1)^{2}$, with any subgroup $M\leqslant{\rm GL}(2,p)$ of order coprime to $2p$ being abelian.
   • (3)
-    $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ is non-abelian with $a^{b}=a^{p^{\ell-1}+1}$, $\ell\geqslant 2$, and ${\sf Aut}(G)$ is solvable such that any Hall $p^{\prime}$-subgroups of ${\sf Aut}(G)$ is conjugate to $\langle\rho\rangle={\sf Z}_{p-1}$, where $a^{\rho}=a^{\xi}\not=a$ with $\xi\,\,\big|\,\,p-1$, and $b^{\rho}=b$.
+    $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ is non-abelian with $a^{b}=a^{p^{\ell-1}+1}$, $\ell\geqslant 2$, and ${\sf Aut}(G)$ is solvable of order $(p-1)p^{\ell+1}$, with a cyclic Hall $p^{\prime}$-subgroup ${\sf Aut}(G)_{p^{\prime}}=\langle\rho\rangle={\sf Z}_{p-1}$, where $a^{\rho}=a^{\xi}\not=a$ and $b^{\rho}=b$.
 Proof.
@@
 (i) The automorphism groups of cyclic groups are well-known; refer to [15, Thm. 7.3].
-(ii) Suppose $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$ with $\ell>1$. Then by [1, Thm. 3.2], ${\sf Aut}(G)$ is of order
-$\begin{array}[]{ccl}|{\sf Aut}(G)|&=&|{\sf Aut}({\sf Z}_{p^{\ell}})|\cdot|{\sf Aut}({\sf Z}_{p})|\cdot|{\sf Hom}({\sf Z}_{p^{\ell}},{\sf Z}_{p})|\cdot|{\sf Hom}({\sf Z}_{p},{\sf Z}_{p^{\ell}})|\\ &=&(p-1)p^{\ell-1}\cdot(p-1)\cdot p\cdot p=p^{\ell+1}(p-1)^{2}.\end{array}$
-More explicitly, let $G=\langle a\rangle\times\langle b\rangle$ with $|a|=p^{\ell}$, $|b|=p$. Then $\Phi(G)=\langle a^{p}\rangle$. Let $\overline{G}=G/\Phi(G)$, and consider the homomorphism
+(ii) Suppose $G={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}=\langle a,b\rangle$, with $|a|=p^{\ell}>p$, $|b|=p$. By [1, Thm. 3.2],
+$\begin{array}[]{ccl}|{\sf Aut}(G)|&=&|{\sf Aut}({\sf Z}_{p^{\ell}})|\cdot|{\sf Aut}({\sf Z}_{p})|\cdot|{\sf Hom}({\sf Z}_{p^{\ell}},{\sf Z}_{p})|\cdot|{\sf Hom}({\sf Z}_{p},{\sf Z}_{p^{\ell}})|\\ &=&(p-1)p^{\ell-1}\cdot(p-1)\cdot p\cdot p\\ &=&p^{\ell+1}(p-1)^{2}.\end{array}$
+In particular, ${\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle)\leqslant{\sf Aut}(G)$, so the subgroup
+$({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{p^{\prime}}\cong({\sf Z}_{(p-1)p^{\ell-1}}\times{\sf Z}_{p-1})_{p^{\prime}}={\sf Z}_{p-1}\times{\sf Z}_{p-1}$
+is an abelian Hall $p^{\prime}$-subgroup of ${\sf Aut}(G)$ of order $(p-1)^{2}$.
+Let $\Phi(G)=\langle a^{p}\rangle$ be the Frattini subgroup of $G$, and let $\overline{G}=G/\Phi(G)$. Consider the homomorphism
 $\sigma:{\sf Aut}(G)\rightarrow{\sf Aut}(\overline{G})={\rm GL}(2,p).$
@@
 $\langle\overline{b}\rangle=\{\overline{1},\overline{b},\cdots,\overline{b}^{p-1}\}=\{\Phi(G),b\Phi(G),\cdots,b^{p-1}\Phi(G)\}$
-exactly consists of those elements of $G$ of order less than $p^{\ell}$.
-Moreover, it follows from ${\sf Aut}(G)/{\rm Ker}(\sigma)\cong{\rm Im}(\sigma)$ that $|{\rm Im}(\sigma)|=p(p-1)^{2}=|P|$, and so ${\rm Im}(\sigma)=P$.
-Now both ${\rm Ker}(\sigma)$ and ${\rm Im}(\sigma)$ are solvable, so is ${\sf Aut}(G)$. Note that the following subgroup
-$X=({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{p^{\prime}}={\sf Z}_{p-1}\times{\sf Z}_{p-1}$
-is a Hall $p^{\prime}$-subgroup of $G$. Then any other Hall $p^{\prime}$-subgroup is conjugate to $X$.
+exactly consists of those elements of $G$ of order less than $p^{\ell}$. Moreover, ${\rm Im}(\sigma)=P$, since $|{\rm Im}(\sigma)|=|{\sf Aut}(G)|/|{\rm Ker}(\sigma)|=p(p-1)^{2}=|P|$.
+Now, both ${\rm Ker}(\sigma)$ and ${\rm Im}(\sigma)$ are solvable, so is ${\sf Aut}(G)$.
 (iii) Suppose $G={\sf Z}_{p}\times{\sf Z}_{p}$.
-Then ${\sf Aut}(G)={\rm GL}(2,p)$ is of order $p(p+1)(p-1)^{2}$, with prime divisors less than or equal to $p$.
-Further, let $M\leqslant{\rm GL}(2,p)$ be a subgroup of order coprime to $2p$, and let $N=M/{\bf Z}(M)$.
-Then
-$N=M/{\bf Z}(M)\leqslant M/({\bf Z}({\rm GL}(2,p))\cap M)\lesssim{\rm PGL}(2,p).$
-Further, since ${\rm PSL}(2,p)N\leqslant{\rm PGL}(2,p)$, we have
-$|N/({\rm PSL}(2,p)\cap N)|\leqslant|{\rm PGL}(2,p)|/|{\rm PSL}(2,p)|=2.$
-It then follows by $2\nmid|N|$ that $N$ is indeed embedded into ${\rm PSL}(2,p)$.
-Note that the subgroups of ${\rm PSL}(2,p)$ are all explicitly known, by check which a subgroup of order coprime to $2p$ must be a cyclic group.
-Then $M$ is an abelian group, since $N=M/{\bf Z}(M)$ is cyclic.
-(iv) Suppose that $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ with $a^{b}=a^{p^{\ell-1}+1}$ is non-abelian, $\ell\geqslant 2$.
-The automorphism group ${\sf Aut}(G)$ is described in [2, Thm.
-3.1 & 3.2], by which we have: $|{\sf Aut}(G)|=(p-1)p^{\ell+1}$, ${\sf Aut}(G)$ is solvable with the Sylow $p$-subgroup normal, and $X=\langle\rho\rangle={\sf Z}_{p-1}$ is a Hall $p^{\prime}$-subgroup of $G$, where $\rho$ stabilizes $\langle a\rangle$ and centralizes $\langle b\rangle$, that is, $a^{\rho}=a^{\xi}\not=a$ with $\xi\,\,\big|\,\,p-1$, and $b^{\rho}=b$.
+Then, ${\sf Aut}(G)={\rm GL}(2,p)$ is of order $(p^{2}-1)(p^{2}-p)=p(p+1)(p-1)^{2}$, with prime divisors less than or equal to $p$.
+Let $M\leqslant{\rm GL}(2,p)$ be a subgroup of order coprime to $2p$, and let $N=M/{\bf Z}(M)$.
+Then,
+$N=M/{\bf Z}(M)\leqslant M/({\bf Z}({\rm GL}(2,p))\cap M)\lesssim{\rm PGL}(2,p),$
+and, in fact, $N$ is embedded into ${\rm PSL}(2,p)$, because $2\nmid|N|$, and meanwhile, ${\rm PSL}(2,p)$ is a index 2 normal subgroup of ${\rm PGL}(2,p)$.
+The subgroups of ${\rm PSL}(2,p)$ are all explicitly known, by check which a subgroup of order coprime to $2p$ must be a cyclic group.
+Thus, $N=M/{\bf Z}(M)$ is cyclic, so that, $M$ is abelian.
+(iv) Suppose $G=\langle a\rangle{:}\langle b\rangle={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ with $a^{b}=a^{p^{\ell-1}+1}$, $\ell\geqslant 2$.
+By [2, Thm.
+3.1], $|{\sf Aut}(G)|=(p-1)p^{\ell+1}$, and hence the Sylow subgroup ${\sf Aut}(G)_{p}$ is normal in ${\sf Aut}(G)$.
+By [2, Thm.
+3.2], there exists a cyclic Hall $p^{\prime}$-subgroup ${\sf Aut}(G)_{p^{\prime}}=\langle\rho\rangle={\sf Z}_{p-1}$, where $\rho$ stabilizes $\langle a\rangle$ and centralizes $\langle b\rangle$ as $a^{\rho}=a^{\xi}\not=a$ and $b^{\rho}=b$.
+Then, ${\sf Aut}(G)={\sf Aut}(G)_{p}{:}{\sf Aut}(G)_{p^{\prime}}$ is solvable.
+$\Box$
+The following corollary will be used in the next Section.
+Corollary 4.4.
+Let $H$ be a $3$-group satisfying Hypothesis 2.2. Then the following statements hold.
+  • (1)
+    Let $\sigma\in{\sf Aut}(H)$ be an involution.
+Then $H{:}\langle\sigma\rangle$ is isomorphic to one of ${\rm D}_{2.3^{\ell}}$, ${\rm D}_{2.3^{\ell}}\times{\sf Z}_{3}$, ${\sf Z}_{3^{\ell}}\times{\rm D}_{6}$, $({\sf Z}_{3^{\ell}}\times{\sf Z}_{3}){:}{\sf Z}_{2}$ and ${\sf Z}_{3^{\ell}}{:}{\sf Z}_{6}$.
+  • (2)
+    If ${\sf Aut}(H)$ contains a subgroup $X$ such that $X\cong{\sf Z}_{2}^{2}$, then $H={\sf Z}_{3^{\ell}}\times{\sf Z}_{3}$, and $H{:}X={\rm D}_{2.3^{\ell}}\times{\rm D}_{6}$.
+  • (3)
+    The group ${\sf Out}(H)$ contains no subgroups isomorphic to ${\rm A}_{4}$.
+Proof.
+By Lemma 4.3, $H$ is one of ${\sf Z}_{3^{\ell}}$, ${\sf Z}_{3^{\ell}}\times{\sf Z}_{3}$ and ${\sf Z}_{3^{\ell}}{:}{\sf Z}_{3}$. Let $\sigma\in{\sf Aut}(H)$ be an involution.
+(i) If $H={\sf Z}_{3^{\ell}}$, then ${\sf Aut}(H)={\sf Z}_{2.3^{\ell-1}}$ contains only one involution, $H{:}\langle\sigma\rangle={\rm D}_{2.3^{\ell}}$, and ${\sf Out}(H)={\sf Aut}(H)$ contains no subgroups isomorphic to ${\sf Z}_{2}^{2}$ or ${\rm A}_{4}$.
+(ii) If $H={\sf Z}_{3^{\ell}}{:}{\sf Z}_{3}=\langle a\rangle{:}\langle b\rangle$ is non-abelian as given in Lemma 4.3 (3), then ${\sf Aut}(H)$ is solvable of order $2.3^{\ell+1}$, so $\langle\sigma\rangle$ is a Hall $3^{\prime}$-subgroup of ${\sf Aut}(H)$ of order 2, and it is conjugate to the following one:
+$\langle\rho\rangle={\sf Z}_{2}$, with $(a,b)^{\rho}=(a^{-1},b)$.
+Thus, $H{:}\langle\sigma\rangle\cong H{:}\langle\rho\rangle={\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3}={\sf Z}_{3^{\ell}}{:}{\sf Z}_{6}$.
+Meanwhile, since $|{\sf Aut}(H)|$ is not divisible by 4, ${\sf Aut}(H)$ has no subgroups isomorphic to ${\sf Z}_{2}^{2}$, and ${\sf Out}(H)$ has no subgroups isomorphic to ${\rm A}_{4}$.
+(iii) If $H={\sf Z}_{3}\times{\sf Z}_{3}$, then ${\sf Aut}(H)={\sf Out}(H)$, and ${\sf Aut}(H)={\rm GL}_{2}(3)\cong{\rm Q}_{8}{:}{\rm S}_{3}$ is solvable, containing one central involution and twelve non-central involutions, with the non-central ones conjugate to the element with diagonal matrix (1,1)(-1,1).
+Hence, $H{:}\langle\sigma\rangle$ is isomorphic to $({\sf Z}_{3}\times{\sf Z}_{3}){:}{\sf Z}_{2}$ or ${\rm D}_{6}\times{\sf Z}_{3}$, where in the former case, ${\bf Z}({\rm GL}_{2}(3))={\sf Z}_{2}$ acts on ${\sf Z}_{3}\times{\sf Z}_{3}$ by mapping each element to its inverse.
+Let $X\leqslant{\sf Aut}(H)$ be a subgroup with $X\cong{\sf Z}_{2}^{2}$.
+Then, $X\cap{\rm Q}_{8}>1$, so $X$ contains the central involution with diagonal matrix (1,1)(-1,-1), and $X$ contains other two non-central involutions conjugate to (1,1)(-1,1) and (1,1)(1,-1).
+Hence $H{:}X={\rm D}_{6}\times{\rm D}_{6}$.
+Moreover, if ${\sf Aut}(H)$ contains a subgroup isomorphic to ${\rm A}_{4}\cong{\sf Z}_{2}^{2}{:}{\sf Z}_{3}$, then ${\sf Z}_{3}$ acts on ${\sf Z}_{2}^{2}$ by moving the central involution, not possible.
+(iv) At last, suppose $H={\sf Z}_{3^{\ell}}\times{\sf Z}_{3}=\langle a,b\rangle$, where $|a|=3^{\ell}>3$, $|b|=3$.
+Then, ${\sf Aut}(H)={\sf Out}(H)$.
+Let $X\leqslant{\sf Aut}(H)$ be a subgroup with $X\cong{\sf Z}_{2}^{2}$.
+By Lemma 4.3 (2), ${\sf Aut}(H)$ is solvable of order $4.3^{\ell+1}$, so $X$ is a Hall $3^{\prime}$-subgroup of ${\sf Aut}(H)$ of order $4$, and it is conjugate to the following one:
+$M=({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{3^{\prime}}={\sf Z}_{2}\times{\sf Z}_{2}=\langle x,y\rangle,$
+where $(a,b)^{x}=(a^{-1},b)$ and $(a,b)^{y}=(a,b^{-1})$.
+Thus, $H{:}X\cong H{:}M={\rm D}_{2.3^{\ell}}\times{\rm D}_{6}$.
+In particular, the involution $\sigma\in{\sf Aut}(H)$ is contained in some Hall $3^{\prime}$-subgroup, thus conjugate to $x$, $y$ or $xy$; respectively, $H{:}\langle\sigma\rangle$ is isomorphic to ${\rm D}_{2.3^{\ell}}\times{\sf Z}_{2}$, ${\sf Z}_{3^{\ell}}\times{\rm D}_{6}$ or $({\sf Z}_{3^{\ell}}\times{\sf Z}_{3}){:}{\sf Z}_{2}$.
+At last, let $g\in{\bf N}_{{\sf Aut}(H)}(M)$. Note that, $M$ stabilizes subgroups $\langle a\rangle$ and $\langle b\rangle$, and meanwhile it cannot stabilize
+    any of the other two subgroups of order $3^{\ell}$, i.e., $\langle ab\rangle$, $\langle ab^{-1}\rangle$, and
+    any of the other two non-characteristic subgroups of $3$, i.e., $\langle a^{3^{\ell-1}}b\rangle$, $\langle a^{-3^{\ell-1}}b\rangle$.
+Now, $M=M^{g}$ stabilizes $\langle a\rangle^{g}$ and $\langle b\rangle^{g}$, which implies $\langle a\rangle^{g}=\langle a\rangle$, $\langle b\rangle^{g}=\langle b\rangle$, so that, ${\bf N}_{{\sf Aut}(H)}(M)\leqslant{\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle)$ is abelian.
+Since any subgroup $Y\leqslant{\sf Aut}(H)$ with $Y\cong{\sf Z}_{2}^{2}$ is conjugate to $M$, the normalizer ${\bf N}_{{\sf Aut}(H)}(Y)$ is conjugate to ${\bf N}_{{\sf Aut}(H)}(M)$, so it is abelian, too.
+Consequently, ${\sf Aut}(H)$ has no subgroup isomorphic to ${\rm A}_{4}$.
 $\Box$
 5. Proof of Theorem 1.1
-In this section, we study the way how the $p$-groups described in Section 4 compose a larger solvable group, and give a proof of Theorem 1.1
+In this section, we give a proof of Theorem 1.1, by studying the way how the $p$-groups compose a larger solvable group.
 Let $G$ be a finite group, with $\{p_{1},\ldots,p_{s}\}$ the prime divisors of $|G|$.
 For $p=p_{i}$, denote by ${\bf O}_{p}(G)$ the maximal normal $p$-subgroup of $G$, or equivalently, the intersection of all Sylow $p$-subgroups of $G$.
-Let $F={\rm Fit}(G)$ be the Fitting subgroup of $G$, which is the largest nilpotent normal subgroup.
-Then
+Let $F={\rm Fit}(G)$ be the Fitting subgroup of $G$, which is the unique largest nilpotent normal subgroup.
+Then,
 $F=F_{p_{1}}\times\cdots\times F_{p_{s}},\mbox{\ with\ each\ }F_{p_{i}}={\bf O}_{p_{i}}(G).$
-Note that both ${\bf O}_{p}(G)$ and $F$ are characteristic in $G$. Further, if $G$ is solvable, then ${\bf C}_{G}(F)\leqslant F$, and so
+Note that, both ${\bf O}_{p}(G)$ and ${\rm Fit}(G)$ are characteristic in $G$. Further, if $G$ is solvable, then ${\bf C}_{G}(F)\leqslant F$, and so
 $G/F={\bf N}_{G}(F)/\left(F{\bf C}_{G}(F)\right)\lesssim{\sf Out}(F).$
@@
 Lemma 5.1.
-Let $G$ be a solvable group of odd order satisfying Hypothesis 2.2.
-Let $p$ be the largest prime divisor of $|G|$.
-Then a Sylow $p$-subgroup $G_{p}$ is normal in $G$, and so $G=G_{p}{:}G_{p^{\prime}}$.
+Let $G$ be a solvable group of odd order satisfying Hypothesis 2.2. Let $p$ be the largest prime divisor of $|G|$. Then:
+  • (1)
+    a Sylow p-subgroup $G_{p}$ is normal in $G$, and hence $G=G_{p}{:}G_{p^{\prime}}\,;$
+  • (2)
+    if $G$ is a $\{p,q\}$-group, with q a prime such that $q<p$ and $q\nmid p^{2}-1$, then $G=G_{p}\times G_{q}$.
 Proof.
-Let $\overline{G}=G/{\bf O}_{p}(G)$, and let $F={\rm Fit}(\overline{G})$. Notice that ${\bf O}_{p}(\overline{G})=1$. We have $F=F_{p_{1}}\times\cdots\times F_{p_{s}}$ with all $p_{i}<p$. Now,
-$\overline{G}/F\lesssim{\sf Out}(F)={\sf Out}(F_{p_{1}})\times\cdots\times{\sf Out}(F_{p_{s}}),$
-and so $p\nmid|\overline{G}/F|$ by Lemma 4.3. Thus, $p\nmid|\overline{G}|$, so that, $G_{p}={\bf O}_{p}(G)\lhd G$. $\Box$
+(1) Let $\overline{G}=G/{\bf O}_{p}(G)$. Note that, $\overline{G}$ has no normal $p$-subgroup. Thus, $F:={\rm Fit}(\overline{G})=F_{p_{1}}\times\cdots\times F_{p_{s}}$ with all $p_{i}<p$, and
+$\overline{G}/F\lesssim{\sf Out}(F)={\sf Out}(F_{p_{1}})\times\cdots\times{\sf Out}(F_{p_{s}}).$
+By Lemma 4.3, $p\nmid|{\sf Out}(F_{p_{i}})|$ as $p>p_{i}$. Then, $p\nmid|\overline{G}/F|$. It follows that $p\nmid|\overline{G}|$, so $G_{p}={\bf O}_{p}(G)\lhd G$ and $G=G_{p}{:}G_{p^{\prime}}$.
+(2) Suppose that $G$ is a $\{p,q\}$}-group with $p>q$ and $q\nmid p^{2}-1$.
+By (1), we have $G=G_{p}{:}G_{q}$.
+Meanwhile, $H:=G/{\bf O}_{q}(G)$ has no normal $q$-subgroup, so ${\rm Fit}(H)$ is a $p$-group, and $H/{\rm Fit}(H)\lesssim{\sf Out}({\rm Fit}(H))$.
+By Lemma 4.3, any prime divisor of $|{\sf Out}({\rm Fit}(H))|$ is a divisor of $p(p^{2}-1)$.
+Then, $q\nmid|H/{\rm Fit}(H)|$ as $q\nmid p(p^{2}-1)$.
+It follows that $q\nmid|H|$, so $G_{q}={\bf O}_{q}(G)\lhd G$ and $G=G_{p}\times G_{q}$.
+$\Box$
+Before giving a general characterization of solvable groups of odd order, we need the following fact.
 Lemma 5.2.
-Let $G$ be a $\{p,q\}$-group satisfying Hypotheses 2.2, where $p>q$ are both odd primes, and $q\nmid p^{2}-1$. Then $G=G_{p}\times G_{q}$.
+Let $H={\sf Z}_{p^{\ell}}$, $1\neq x\in{\sf Aut}(H)$ be an automorphism with $\gcd(|x|,p)=1$, and $X=H{:}\langle x\rangle$. Then, (i) $H\leqslant X^{\prime}\,;$ (ii) $H\cap{\bf Z}(X)=1$.
 Proof.
-By Lemma 5.1, $G=G_{p}{:}G_{q}$.
-Since $G$ is solvable, we obtain that $G_{q}/{\bf C}_{G_{q}}(G_{p})\cong G/(G_{p}{\bf C}_{G_{q}}(G_{p}))=G/(G_{p}{\bf C}_{G}(G_{p}))\leq{\sf Out}(G_{p})$.
-By Lemma 4.3, any prime divisor of $|{\sf Out}(G_{p})|$ divides $p(p^{2}-1)$.
-However, $q\nmid p(p^{2}-1)$ by our assumption, and we conclude that $G_{q}/{\bf C}_{G_{q}}(G_{p})=1$, so that $G_{q}={\bf C}_{G_{q}}(G_{p})$ and $G=G_{p}\times G_{q}$.
-$\Box$
-Before giving a general characterization of solvable groups of odd order, we need the following technical lemma.
-Lemma 5.3.
-Let $H={\sf Z}_{p^{\ell}}$ be a cyclic group and $x$ be an automorphism of $H$ with order coprime to $p$. Let $X=H{:}\langle x\rangle$. Then, (i) $H\leqslant X^{\prime};$ (ii) $H\cap{\bf Z}(X)=1$.
-Proof.
-By Lemma 4.3 (1), the order of $x$ divides p1p-1. Let $H=\langle a\rangle$ and suppose
+Let $H=\langle a\rangle$. By Lemma 4.3 (1), $|x|$ divides p1p-1, and thus we can suppose
 $a^{x}=a^{m}$, where $1<m\leqslant p^{\ell}-1$,  $p\nmid m$  and  $p^{\ell}\mid m^{p-1}-1$.
-If $p\mid m-1$, then by LTE (Lift the Exponent Lemma) we have $p^{\ell}\mid m-1$, a contradiction.
+If $p\mid m-1$, then by LTE (Lift the Exponent Lemma), we have $p^{\ell}\mid m-1$, so $a^{x}=a^{m}=a$ and $x=1$, a contradiction.
 Hence $p\nmid m-1$.
@@
 Lemma 5.3.
-Let $G$ be a solvable group of odd order satisfying Hypothesis 2.2. Then G=C:DG=C{:}D, where $C$ is abelian, $D$ is nilpotent, and $C\cap{\bf Z}(G)=1$.
+Let $G$ be a solvable group of odd order satisfying Hypothesis 2.2. Then G=A:BG=A{:}B, where $A$ is abelian, $B$ is nilpotent, and $A\cap{\bf Z}(G)=1$.
 Proof.
-Let $n$ be the number of prime divisors of $|G|$.
-The proof proceeds by induction on $n$.
-For $n=1$, the lemma holds with $C=1$, $D=G$.
+The proof proceeds by induction on $n$, the number of prime divisors of $|G|$.
+For $n=1$, the lemma holds with $A=1$, $B=G$.
 Suppose $n\geqslant 2$, and assume inductively that the lemma holds for groups of order with less prime divisors.
 Let $p$ be the largest prime divisor of $|G|$. By Lemma 5.1 and the induction assumption, we have
-$G=G_{p}{:}G_{p^{\prime}}=G_{p}{:}(A{:}B),$
-with $A$ abelian, $B$ nilpotent, and $A\cap{\bf Z}(A{:}B)=1$, so that, $B$ does not centralize any Sylow subgroup of $A$. Further, if $G=G_{p}\times G_{p^{\prime}}$, then
-$G=G_{p}{\times}(A{:}B)=(G_{p}{\times}A){:}B=A{:}(G_{p}\times B),$
-which gives an expected factorization of $G$. We thus assume that $G\not=G_{p}\times G_{p^{\prime}}$, namely, $G_{p^{\prime}}$ acts by conjugation on $G_{p}$ non-trivially. We shall confirm
-Claim: $G_{p}$ is centralized by each Sylow $r$-subgroup $A_{r}$ of $A$, so that, $G_{p}$ is centralized by $A$.
-By Lemma 4.3, we have two cases to treat: $A_{r}={\sf Z}_{r^{k}}$ or ${\sf Z}_{r^{k}}\times{\sf Z}_{r}$.
-Case 1. Let $A_{r}={\sf Z}_{r^{k}}$. If $A_{r},B_{r^{\prime}}]=1$, then by Lemma 4.3, $(A{:}B)_{r}=A_{r}{:}B_{r}={\sf Z}_{r^{k}}{:}{\sf Z}_{r}$ is non-abelian, and
-$1\not=A_{r}\cap{\bf Z}(A_{r}{:}B_{r})=A_{r}\cap{\bf Z}(A{:}B)\leqslant A\cap{\bf Z}(A{:}B),$
-a contradiction.
-Thus there exists a Sylow $q$-subgroup $B_{q}$ with $q\not=r$ which does not centralize $A_{r}$.
-Take $x\in B_{q}$ to be an element acting non-trivially on $A_{r}$ and let $X=A_{r}{:}\langle x\rangle$.
-Then $X$ acts on $G_{p}$ and induces an abelian automorphism group of $G_{p}$ by Lemma 4.3.
-It then follows by Lemma 5.3 (i) that $A_{r}\leqslant X^{\prime}\leqslant{\bf C}_{X}(G_{p})$.
-Case 2. Let $A_{r}={\sf Z}_{r^{k}}\times{\sf Z}_{r}$, where $k\geqslant 1$. Then $r$ does not divide $|B|$. By Lemma 4.3 (2), it implies that either
-  • (i)
-    $A_{r}={\sf Z}_{r}^{2}$ and $B$ is irreducible on $A_{r}$, or
-  • (ii)
-    there exists $a,b\in A_{r}$ such that $A_{r}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$ and $\langle b\rangle$ are normalized by $B$.
-For the former, $A_{r}$ centralizes $G_{p}$ since $A_{r}{:}B$ induces an abelian automorphism group of $G_{p}$ by Lemma 4.3.
-For the latter, there exist elements $u,v\in B$ such that $u$ acts on $\langle a\rangle$ non-trivially, and $v$ acts on $\langle b\rangle$ non-trivially.
-Then similar to Case 1 we have $G_{p},a]=[G_{p},b]=1$, so that, $G_{p},A_{r}]=[G_{p},\langle a,b\rangle]=1$.
-We have therefore confirmed the claim. Consequently we have
-$G=G_{p}{:}(A{:}B)=(G_{p}\times A){:}B.$
-Meanwhile, by Lemma 4.3, $G_{p}$ is one of the following groups:
-$G_{p}={\sf Z}_{p^{\ell}}$,  ${\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$  or  ${\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$.
-Note that by assumption, $G_{p^{\prime}}=A{:}B$ acts by conjugation on $G_{p}$ non-trivially, which implies that $B$ acts on $G_{p}$ non-trivially.
-Suppose $G_{p}={\sf Z}_{p^{\ell}}$.
-As $B$ acts on $G_{p}$ non-trivially, it follows from Lemma 5.3 (ii) that $G_{p}\cap{\bf Z}(G)=1$.
-Thus G=C:DG=C{:}D satisfies the theorem, with $C=G_{p}\times A$, $D=B$.
-Suppose $G_{p}={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}=\langle a\rangle{:}\langle b\rangle$.
-By Lemma 4.3 (3), $B$ acts on $\langle a\rangle$ non-trivially and acts on $\langle b\rangle$ trivially.
+$G=G_{p}{:}G_{p^{\prime}}=G_{p}{:}(C{:}D),$
+with $C$ abelian, $D$ nilpotent, and $C\cap{\bf Z}(C{:}D)=1$, so that, $D$ does not centralize any Sylow subgroup of $C$.
+If $G=G_{p}\times G_{p^{\prime}}$, then
+$G=G_{p}{\times}(C{:}D)=(G_{p}{\times}C){:}D=C{:}(G_{p}\times D),$
+which gives an expected factorization of $G$, with $A=C$, $B=G_{p}\times D$.
+We thus assume that $G\not=G_{p}\times G_{p^{\prime}}$, namely, $G_{p^{\prime}}$ acts by conjugation on $G_{p}$ non-trivially. We shall confirm the following claim.
+Claim: $G_{p}$ is centralized by each Sylow $r$-subgroup $C_{r}$ of $C$, so that, $G_{p}$ is centralized by $C$.
+By Lemma 4.3, we have two cases to treat: $C_{r}={\sf Z}_{r^{k}}$ or ${\sf Z}_{r^{k}}\times{\sf Z}_{r}$.
+Case 1.
+Let $C_{r}={\sf Z}_{r^{k}}$.
+If $C_{r},D_{r^{\prime}}]=1$, since $C_{r},D]\neq 1$, then $(C{:}D)_{r}=C_{r}{:}D_{r}={\sf Z}_{r^{k}}{:}{\sf Z}_{r}$ is non-abelian as given in Lemma 4.3 (3), and so
+$1\not=C_{r}\cap{\bf Z}(C_{r}{:}D_{r})=C_{r}\cap{\bf Z}(C{:}D)\leqslant C\cap{\bf Z}(C{:}D),$
+a contradiction. Thus, there exists a Sylow $q$-subgroup $D_{q}$ with $q\not=r$ which does not centralize $C_{r}$. Let $x\in D_{q}$ act non-trivially on $C_{r}$, and let
+$X=C_{r}{:}\langle x\rangle.$
+Then, $X$ acts on $G_{p}$ with $\overline{X}=X/{\bf C}_{X}(G_{p})\lesssim{\sf Aut}(G_{p})$.
+Note that, $|X|$ is coprime to $2p$, so is $|\overline{X}|$.
+By Lemma 4.3, either ${\sf Aut}(G_{p})={\rm GL}_{2}(p)$, or ${\sf Aut}(G_{p})$ is solvable with an abelian Hall $p^{\prime}$-subgroup, while both cases imply that $\overline{X}$ is abelian, that is, $X^{\prime}\leqslant{\bf C}_{X}(G_{p})$.
+At last, by Lemma 5.2 (i), we have $C_{r}\leqslant X^{\prime}$, and so $C_{r}\leqslant{\bf C}_{X}(G_{p})$.
+Case 2. Let $C_{r}={\sf Z}_{r^{k}}\times{\sf Z}_{r}$, where $k\geqslant 1$. In this case, $C_{r}=(C{:}D)_{r}$, namely, $r$ does not divide $|D|$. It follows by Lemma 4.3 (2) that, either
+    $C_{r}={\sf Z}_{r}^{2}$ and $D$ is irreducible on $C_{r}$; or
+    there exists $a,b\in C_{r}$ such that $C_{r}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$ and $\langle b\rangle$ are normalized by $D$.
+For the former, let $x\in D$ act non-trivially on ${\sf Z}_{r}^{2}$, and let $X={\sf Z}_{r}^{2}{:}\langle x\rangle$, which has ${\sf Z}_{r}^{2}$ as a minimal normal subgroup; for the latter, there exist $u,v\in D$ respectively acts on $\langle a\rangle$, $\langle b\rangle$ non-trivially, so let $Y=\langle a\rangle{:}\langle u\rangle$, $Z=\langle b\rangle{:}\langle v\rangle$.
+Note that, each of $X,Y,Z$ induces some automorphism group of $G_{p}$ of order coprime to $2p$.
+It is similar to Case 1 to show $G_{p},C_{r}]=1$.
+We have therefore confirmed the claim. Consequently, we have
+$G=G_{p}{:}(C{:}D)=(G_{p}\times C){:}D.$
+We then find an expected factorization of $G$.
+By assumption, $G_{p^{\prime}}=C{:}D$ acts by conjugation on $G_{p}$ non-trivially.
+Now, as $C$ acts on $G_{p}$ trivially, $D$ acts on $G_{p}$ non-trivially, so it induces some automorphism of $G_{p}$ of order coprime to $p$.
+By Lemma 4.3, we have three cases to treat:
+(1) Suppose $G_{p}={\sf Z}_{p^{\ell}}$. By Lemma 5.2 (ii), any non-identity element of $G_{p}$ cannot fixed by $D$. Thus G=A:BG=A{:}B, with $A=G_{p}\times C$, $B=D$, satisfies the theorem.
+(2) Suppose $G_{p}={\sf Z}_{p^{\ell}}{:}{\sf Z}_{p}$ as given in Lemma 4.3 (3).
+Then there exist $a,b\in G_{p}$ such that $G_{p}=\langle a\rangle{:}\langle b\rangle$, $\langle a\rangle$ is normalized by $D$, and $\langle b\rangle$ is centralized by $D$.
 Rewrite $G$ as
-$G=(A\times(\langle a\rangle{:}\langle b\rangle)){:}B=(A\times\langle a\rangle){:}(B\times\langle b\rangle).$
-Then by Lemma 5.3 (ii), G=C:DG=C{:}D satisfies the theorem, with $C=A\times\langle a\rangle$, $D=B\times\langle b\rangle$.
-Suppose $G_{p}={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$, where $\ell\geqslant 1$.
-By Lemma 4.3 (2), it implies that either (i) $G_{p}={\sf Z}_{p}^{2}$ and $B$ is irreducible on $G_{p}$; or (ii) there exists $a,b\in G_{p}$ such that $G_{p}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$ and $\langle b\rangle$ are normalized by $B$.
-For the former, G=C:DG=C{:}D satisfies the theorem, with $C=G_{p}\times A$, $D=B$.
-For the latter, if $B$ acts non-trivially on both $\langle a\rangle$ and $\langle b\rangle$, then also let $C=G_{p}\times A$, $D=B$.
-Otherwise, $B$ acting non-trivially on $G_{p}$ implies $B$ acting non-trivially on some one of $\langle a\rangle$ and $\langle b\rangle$, say $\langle a\rangle$.
-Then similar to the last paragraph, let $C=A\times\langle a\rangle$, $D=B\times\langle b\rangle$.
+$G=(C\times(\langle a\rangle{:}\langle b\rangle)){:}D=(C\times\langle a\rangle){:}(D\times\langle b\rangle).$
+By Lemma 5.2 (ii), G=A:BG=A{:}B, with $A=C\times\langle a\rangle$, $B=D\times\langle b\rangle$, satisfies the theorem.
+(3) Suppose $G_{p}={\sf Z}_{p^{\ell}}\times{\sf Z}_{p}$.
+It follows by Lemma 4.3 (2) that, either $G_{p}={\sf Z}_{p}^{2}$ and $D$ is irreducible on $G_{p}$; or there exist $a,b\in G_{p}$ such that $G_{p}=\langle a\rangle\times\langle b\rangle$, and both $\langle a\rangle$, $\langle b\rangle$ are normalized by $D$.
+For the former, G=A:BG=A{:}B satisfies the theorem, with $A=G_{p}\times C$, $B=D$.
+For the latter, if $D$ acts non-trivially on both $\langle a\rangle$ and $\langle b\rangle$, then also let $A=G_{p}\times C$, $B=D$; otherwise, $D$ acts non-trivially on some one of $\langle a\rangle$ and $\langle b\rangle$, say $\langle a\rangle$, so let $A=C\times\langle a\rangle$, $B=D\times\langle b\rangle$.
 $\Box$
 We next deal with solvable groups of even order.
-Lemma 5.5.
-Let $G$ be a solvable group satisfying Hypothesis 2.2. If $F:={\rm Fit}(G)={\bf O}_{2}(G)$, then either $G$ is a $2$-group, or $G$ is one of the following:
-  • (1)
-    $G={\sf Z}_{2}^{3}{:}{\sf Z}_{7}$ or ${\sf Z}_{2}^{3}{:}{\sf Z}_{7}{:}{\sf Z}_{3};$
-  • (2)
-    $G={\sf Z}_{2}^{2}{:}H$, ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}H)$, ${\rm Q}_{8}{:}H$ or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}H)$, where $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$, ${\sf Z}_{2}^{2}{:}H\cong{\rm A}_{4}$ or ${\rm S}_{4}$, and ${\rm Q}_{8}{:}H\cong{\rm SL}(2,3)$ or ${\rm GL}(2,3)$.
-Proof.
-Assume that $G$ is not a 2-group. Then $G/F\leqslant{\sf Out}(F)$ is not a 2-group either. Thus by Lemma 4.1, $F$ is one of the following 2-groups:
-$F\in\{{\sf Z}_{2}^{2},{\sf Z}_{2}^{3},{\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8}\}.$
-Moreover, ${\bf O}_{2}(G/F)$ is trivial, namely, $G/F$ has no normal 2-subgroups.
-Then by checking ${\sf Out}(F)$, either $|G/F|$ is divisible by 7, and $F={\sf Z}_{2}^{3}$, $G/F={\sf Z}_{7}$ or ${\sf Z}_{7}{:}{\sf Z}_{3}$; or $G/F={\sf Z}_{3}$ or ${\rm S}_{3}$.
-It follows that $G$ is one of the groups listed in the lemma.
-In particular, if $G=F{:}{\rm S}_{3}$, then $G_{2}=F{:}{\sf Z}_{2}>F$, and respectively for $F={\sf Z}_{2}^{2}$, ${\sf Z}_{2}^{3}$, ${\rm Q}_{8}$ or ${\sf Z}_{4}\circ{\rm Q}_{8}$, it is easy to verify that $G_{2}$ has a subgroup $K$ of index $2$ with $K={\sf Z}_{4}$, ${\rm D}_{8}$, ${\sf Z}_{8}$ or ${\rm D}_{16}$.
-That is, $G=F{:}{\rm S}_{3}$ surely satisfies Hypothesis 2.2.
-$\Box$
-Lemma 5.6.
-Let $K$ be a $\{2,p\}$-group with $p$ odd prime satisfying Hypotheses 2.2. If $K\neq K_{p}{:}K_{2}$, then one of the following holds.
+Proposition 5.4.
+Let $K$ be a $\{2,p\}$-group with $p$ odd prime satisfying Hypotheses 2.2. Assume that $K\neq K_{p}{:}K_{2}$. Then $p\in\{3,7\}$ and one of the following holds.
   1. (1)
-    $p=7$, $K={\sf Z}_{2}^{3}{:}K_{7};$
+    $K={\sf Z}_{2}^{3}{:}K_{7}$,  ${\sf Z}_{2}^{2}{:}K_{3}$,  ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}K_{3})$,  ${\rm Q}_{8}{:}K_{3}$ or ${\sf Z}_{4}\circ_{{\sf Z}_{2}}({\rm Q}_{8}{:}K_{3})\,;$
   2. (2)
-    $p=3$, $K=X{:}K_{3}$ with $X\in\{{\sf Z}_{2}^{2},\,{\sf Z}_{2}^{3},\,{\rm Q}_{8},\,{\sf Z}_{4}\circ{\rm Q}_{8}\};$
-  3. (3)
-    $p=3$, $K$ is one of the groups in Tables 1 - 2
-Proof.
-Suppose that $K_{2}$ is normal in $K$, that is, $K=K_{2}{:}K_{p}$. Then $K_{2}$ has an automorphism of order $p$. By Lemma 4.1, $K_{2}$ is one of the following groups:
-${\sf Z}_{2}^{2},\ {\rm Q}_{8},\ {\sf Z}_{2}^{3}\ \mbox{and}\ {\sf Z}_{4}\circ{\rm Q}_{8},$
-and either $p=3$, $K=K_{2}{:}K_{3}$; or $K_{2}={\sf Z}_{2}^{3}$, $p=7$ and $K={\sf Z}_{2}^{3}{:}K_{7}$, as in (1) - (2).
-We then assume that neither $K_{p}$ nor $K_{2}$ is normal in $K$. Let
-$F={\rm Fit}(K)=F_{p}\times F_{2}$, with $F_{p}<K_{p}$, $F_{2}<K_{2}$.
-Consider the quotient group $\overline{K}=K/F_{p}$, whose Fitting group is a 2-group, namely ${\rm Fit}(\overline{K})={\bf O}_{2}(\overline{K})$.
-Thus, by Lemma 5.5, $\overline{K}={\bf O}_{2}(\overline{K}){:}H$, where either
-(i) ${\bf O}_{2}(\overline{K})\in\{{\sf Z}_{2}^{2},{\sf Z}_{2}^{3},{\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8}\}$, $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$; or (ii) ${\bf O}_{2}(\overline{K})={\sf Z}_{2}^{3}$, $H={\sf Z}_{7}$ or ${\sf Z}_{3}$.
-Further, if (ii) holds, $\overline{K}={\sf Z}_{2}^{3}{:}{\sf Z}_{p}$ with $p=3$ or $7$, then ${\sf Z}_{2}^{3}$ is a minimal normal subgroup of $\overline{K}$, and so $\overline{F_{2}}=F_{2}$ as a normal subgroup of $\overline{K}$ must be identity.
-Consequently, we have $F=F_{p}$, ${\sf Z}_{2}^{3}{:}{\sf Z}_{p}=K/F_{p}\lesssim{\sf Out}(F_{p})$, which does not hold for the $p$-groups appearing in Lemma 4.3.
-Therefore, $\overline{K}={\bf O}_{2}(\overline{K}){:}H$ is given in (i). We then analyze all the possibilities for $F_{2}=\overline{F_{2}}$, which a normal 2-subgroup of $\overline{K}$.
-Case 1. $F_{2}={\bf O}_{2}(\overline{K})$. In this case, we have $H={\rm S}_{3}$ since $F_{2}<K_{2}$. Then
-$K=F_{3}.(F_{2}{:}{\rm S}_{3})=(F_{3}\times F_{2}).{\rm S}_{3}=F_{2}{:}(F_{3}.{\rm S}_{3})=F_{2}{:}(K_{3}{:}{\sf Z}_{2}).$
-Note that $K_{3}{:}{\sf Z}_{2}/F_{3}\cong{\rm S}_{3}$.
-By checking the $3$-groups appearing in Lemma 4.3, we conclude that there are seven possibilities for the triple $T:=(K_{3},\,K_{3}{:}{\sf Z}_{2},\,F_{3})$ as follows (also note that ${\sf Aut}({\sf Z}_{3}\times{\sf Z}_{3})={\rm GL}_{2}(3)$ is solvable):
-[…]
-Consequently, Case 1 gives the groups listed in Table 1.
-Table 1. Case 1: $F_{2}={\bf O}_{2}(\overline{K})$
-case $F_{2}$ ${\sf Z}_{2}^{2}$ ${\sf Z}_{2}^{3}$ ${\rm Q}_{8}$ ${\sf Z}_{4}\circ{\rm Q}_{8}$
-(1.1) ${\rm S}_{4}$ ${\sf Z}_{2}\times{\rm S}_{4}$ ${\rm GL}_{2}(3)$ ${\sf Z}_{4}\circ{\rm GL}_{2}(3)$
-(1.2) ${\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}}$ ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}})$ ${\rm Q}_{8}{:}{\rm D}_{2.3^{\ell+1}}$ ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell+1}})$
-(1.3) ${\sf Z}_{3^{\ell}}\times{\rm S}_{4}$ ${\sf Z}_{2.3^{\ell}}\times{\rm S}_{4}$ ${\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\rm S}_{3})$ ${\sf Z}_{4.3^{\ell}}\circ({\rm Q}_{8}{:}{\rm S}_{3})$
-(1.4) ${\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{6}\times({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{3}\times({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$ ${\sf Z}_{12}\circ({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})$
-(1.5) $({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$ $({\sf Z}_{2.3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$ $({\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\rm S}_{3})){:}{\sf Z}_{2}$ $({\sf Z}_{4.3^{\ell}}\circ({\rm Q}_{8}{:}{\rm S}_{3})){:}{\sf Z}_{2}$
-(1.6) $({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$ $({\sf Z}_{6}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$ $({\sf Z}_{3}\times({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$ $({\sf Z}_{12}\circ({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$
-(1.7) ${\sf Z}_{2}^{2}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3})$ ${\sf Z}_{2}\times{\sf Z}_{2}^{2}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3})$ ${\rm Q}_{8}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3})$ ${\sf Z}_{4}\circ({\rm Q}_{8}{:}({\rm D}_{2.3^{\ell}}{:}{\sf Z}_{3}))$
-Case 2. $(F_{2},{\bf O}_{2}(\overline{K}))=({\sf Z}_{2}^{2},{\sf Z}_{2}^{3})$ or $({\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8})$. Then
-$K=F_{3}.({\bf O}_{2}(\overline{K}){:}H)=(F_{3}\times F_{2}).({\sf Z}_{2}\times H)=F_{2}{:}(F_{3}.({\sf Z}_{2}\times H)).$
-Further, if $H={\sf Z}_{3}$, then $K/F_{2}=K_{3}{:}{\sf Z}_{2}$.
-Note that ${\sf Z}_{2}$ acts non-trivially on $K_{3}$, and meanwhile it acts trivially on $K_{3}/F_{3}=H$.
-Then by checking the $3$-groups appearing in Lemma 4.3, we conclude that there are three possibilities:
-[…]
-Otherwise, $H={\rm S}_{3}$, and so $K/F_{2}=K_{3}{:}({\sf Z}_{2}\times{\sf Z}_{2})$.
-By Lemma 4.3, ${\sf Z}_{2}\times{\sf Z}_{2}\leqslant{\sf Aut}(K_{3})$ implies $K_{3}={\sf Z}_{3^{\ell}}\times{\sf Z}_{3}$ and $K_{3}{:}({\sf Z}_{2}\times{\sf Z}_{2})={\rm D}_{2.3^{\ell}}\times{\rm D}_{6}$.
-We then get two possibilities:
-[…]
-Consequently, Case 2 gives the groups listed in Table 2.
-Table 2. Case 2: $F_{2}<{\bf O}_{2}(\overline{K})$
-case $(F_{2},{\bf O}_{2}(\overline{K}))$ $({\sf Z}_{2}^{2},{\sf Z}_{2}^{3})$ $({\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8})$
-(2.1) ${\rm D}_{2.3^{\ell}}\times{\rm A}_{4}$ $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\sf Z}_{3})$
-(2.2) $({\sf Z}_{2}^{2}{:}{{\sf Z}_{3^{\ell}}})\times{\rm D}_{6}$ $({\rm Q}_{8}{:}{{\sf Z}_{3^{\ell}}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$
-(2.3) $({\rm D}_{2.3^{\ell}}\times{\sf Z}_{2}^{2}){:}{\sf Z}_{3}$ $(({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ{\rm Q}_{8}){:}{\sf Z}_{3}$
-(2.4) ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\rm S}_{3})$
-(2.5) $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$ $({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$
-Case 3.
-Recall that $F_{2}$ is a normal 2-subgroup of ${\bf O}_{2}(\overline{K})$.
-Notice that ${\rm Q}_{8}{:}{\sf Z}_{3}$ or ${\rm Q}_{8}{:}{\rm S}_{3}$ only has the center ${\sf Z}_{2}$ and ${\rm Q}_{8}$ as normal $2$-subgroups.
-Then all the remaining cases are collected as follows :
-[…]
-For case (3.1), $(F_{2},{\bf O}_{2}(\overline{K}))=(1,{\rm Q}_{8})$, we have
-$F=F_{3}\mbox{\ and\ }K/F_{3}={\rm Q}_{8}{:}H\lesssim{\sf Aut}(F_{3}).$
-By lemma 4.3, it implies that $F_{3}={\sf Z}_{3}\times{\sf Z}_{3}$ and $K_{3}={\sf Z}_{9}{:}{\sf Z}_{3}$ is non-abelian. Let
-$K_{3}=\langle a\rangle{:}\langle b\rangle={\sf Z}_{9}{:}{\sf Z}_{3}$, with $a^{b}=a^{4}$, $b^{a}=a^{6}b$,
-where $F_{3}=\langle a^{3}\rangle\times\langle b\rangle$.
-Note that $K_{2}=({\rm Q}_{8}{:}H)_{2}={\rm Q}_{8}$ or ${\rm Q}_{8}{:}{\sf Z}_{2}$.
-Consider the central involution $c$ of $K_{2}$.
-Then $\overline{c}$ is the central involution of $\overline{K}$.
-On the one hand, $\overline{c}$ acts on $F_{3}$ by mapping each $x$ to $x^{-1}$.
-That is,
-$x^{c}=x^{-1}$, so that, $c^{x}=x^{-2}c$ for any $x\in F_{3}$.
-On the other hand, $\overline{c}^{\overline{a}}=\overline{c}$, and so $c^{a}=cy$ where $y\in F_{3}$. Let $y=a^{3i}b^{j}$. Then
-$c^{a^{3}}=(ca^{3i}b^{j})^{a^{2}}=(ca^{(6i+6j)}b^{2j})^{a}=ca^{(9i+18j)}b^{3j}=c,\mbox{\ where\ }a^{3}\in F_{3},$
-a contradiction. Hence case (3.1) cannot occur.
-For case (3.2), $(F_{2},{\bf O}_{2}(\overline{K}))=(1,{\sf Z}_{4}\circ{\rm Q}_{8})$, we have ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\sf Z}_{3})\lesssim{\sf Aut}(F_{3})$, which is impossible by checking the 3-groups appearing in Lemma 4.3.
-For case (3.3), since ${\bf O}_{2}(\overline{K})$ acts on $F_{3}$ with the centralizer ${\bf C}_{{\bf O}_{2}(\overline{K})}(F_{3})=F_{2}$, we deduce that in this case ${\sf Aut}(F_{3})$ contains a subgroup isomorphic to ${\rm A}_{4}\cong{\sf Z}_{2}^{2}{:}{\sf Z}_{3}$.
-By Lemma 4.3, $|{\sf Aut}(F_{3})|$ being divisible by $4$ implies $F_{3}={\sf Z}_{3^{\ell}}\times{\sf Z}_{3}$.
-For $\ell=1$, it is clear that ${\sf Aut}({\sf Z}_{3}^{2})={\rm GL}(3,2)\cong{\rm Q}_{8}{:}{\rm S}_{3}$ contains no subgroups isomorphic to ${\rm A}_{4}$.
-We thus assume $\ell>1$ and let $F_{3}=\langle a\rangle\times\langle b\rangle$ with $|a|=3^{\ell}$, $|b|=3$.
-By Lemma 4.3, ${\sf Aut}(F_{3})$ is solvable, and it contains a Hall $3^{\prime}$-subgroup
-$H_{0}=({\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle))_{3^{\prime}}\cong{\sf Z}_{2}^{2},$
-which stabilizes subgroups $\langle a\rangle$ and $\langle b\rangle$. Meanwhile, by simple calculation, $H_{0}$ cannot stabilize
-    any of the other two subgroups of order $3^{\ell}$, i.e., $\langle ab\rangle$, $\langle ab^{-1}\rangle$, and
-    any of the other two non-characteristic subgroups of $3$, i.e., $\langle a^{3^{\ell-1}}b\rangle$, $\langle a^{-3^{\ell-1}}b\rangle$.
-Consider the normalizer $N={\bf N}_{{\sf Aut}(F_{3})}(H_{0})$.
-For any $g\in N$, $H_{0}=H_{0}^{g}$ stabilizes subgroups $\langle a\rangle^{g}$ and $\langle b\rangle^{g}$, and so $\langle a\rangle^{g}=\langle a\rangle$, $\langle b\rangle^{g}=\langle b\rangle$, which implies that $N\leqslant{\sf Aut}(\langle a\rangle)\times{\sf Aut}(\langle b\rangle)$ is abelian.
-Note that any subgroup $H\leqslant{\sf Aut}(F_{3})$ with $H\cong{\sf Z}_{2}^{2}$ is a Hall $3^{\prime}$-subgroup, then conjugate to $H_{0}$.
-Then the normalizer ${\bf N}_{{\sf Aut}(F_{3})}(H)$, which is conjugate to $N$, is also abelian.
-Thus, ${\sf Aut}(F_{3})$ has no subgroup isomorphic to ${\rm A}_{4}$.
-Finally, with Case 3 excluded, the proof is complete. $\Box$
+    $K$ is one of the $\{2,3\}$-groups: $X$, ${\sf Z}_{2}\times X$, $Y$, ${\sf Z}_{4}\circ_{{\sf Z}_{2}}Y$ and $W$, where $X,Y,W$ are given in Table 1, $X$ is homomorphic to ${\sf Z}_{2}^{2}{:}{\rm S}_{3}\cong{\rm S}_{4}$, $Y$ is homomorphic to ${\rm Q}_{8}{:}{\rm S}_{3}\cong{\rm GL}_{2}(3)$, and $W$ is homomorphic to ${\sf Z}_{2}^{2}{:}{\sf Z}_{3}\cong{\rm A}_{4}$, ${\rm S}_{4}$, ${\sf Z}_{4}\circ{\rm SL}_{2}(3)$ or ${\sf Z}_{4}\circ{\rm GL}_{2}(3)$.
+We remark that, the proof of Proposition 5.4 has been moved to the end of this Section, as it is rather lengthy.
 Lemma 5.5.
@@
 Proof.
-By Lemma 5.1, $K_{\{3,7\}}=K_{7}{:}K_{3}$. Thus $K_{\{2,7\}}\neq K_{7}{:}K_{2}$, and by Lemma 5.6,
-$K_{\{2,7\}}=K_{2}{:}K_{7}={\sf Z}_{2}^{3}{:}K_{7},$
-Let $F={\rm Fit}(K)=F_{2}\times F_{3}\times F_{7}$.
-As $K_{2}$ is normalized by $K_{7}$, and meanwhile it normalizes $F_{7}$, so $K_{2}$ centralizes $F_{7}$.
-Moreover, $K_{2}$ also centralizes $F_{3}$.
-If not, consider the subgroup $F_{3}{:}(K_{2}{:}K_{7})\leqslant K$.
-Then ${\sf Z}_{2}^{3}$ as a minimal normal subgroup of $K_{2}{:}K_{7}$ is embedded into ${\sf Aut}(F_{3})$, which is impossible due to Lemma 4.3 and the fact ${\sf Z}_{2}^{3}\nleqslant{\rm GL}(2,3)\cong{\rm Q}_{8}{:}{\rm S}_{3}$.
-Hence $K_{2}$ centralizes $F$, so $K_{2}\leqslant C_{K}(F)\leqslant F$, and then $K_{2}=F_{2}\lhd K$, and $K=K_{2}{:}K_{\{3,7\}}=K_{2}{:}(K_{7}{:}K_{3})$.
+By Lemma 5.1, $K_{\{3,7\}}=K_{7}{:}K_{3}$. If $K_{\{2,7\}}=K_{7}{:}K_{2}$, then $K_{7}$ is normal in $K$, not in our case. Thus, $K_{\{2,7\}}\neq K_{7}{:}K_{2}$, and by Proposition 5.4, we have
+$K_{\{2,7\}}=K_{2}{:}K_{7}={\sf Z}_{2}^{3}{:}K_{7}.$
+We claim that, $K_{2}={\sf Z}_{2}^{3}$ centralizes the Fitting group $F={\rm Fit}(K)=F_{2}\times F_{3}\times F_{7}$.
+If this is true, then we are done, since $K_{2}\leqslant C_{K}(F)\leqslant F$, so $K_{2}=F_{2}\lhd K$, and
+$K=K_{2}{:}K_{\{3,7\}}=K_{2}{:}(K_{7}{:}K_{3})={\sf Z}_{2}^{3}{:}(K_{7}{:}K_{3}).$
+It follows by $F_{2}\leqslant K_{2}={\sf Z}_{2}^{3}$ that $K_{2}$ centralizes $F_{2}$.
+Meanwhile, $K_{2}$ is normalized by $K_{7}$, so $K_{2}$ and $F_{7}$ normalizes, and hence centralizes, each other.
+The claim is then proved if $K_{2}$ centralizes $F_{3}$.
+Suppose that $K_{2}$ doe not centralize $F_{3}$.
+Consider the subgroup $F_{3}{:}(K_{2}{:}K_{7})\leqslant K$.
+Then $K_{2}={\sf Z}_{2}^{3}$ as a minimal normal subgroup of $K_{2}{:}K_{7}$ is embedded into ${\sf Aut}(F_{3})$.
+By Lemma 4.3, $|{\sf Aut}(F_{3})|$ being divisible by 8 only happens when $F_{3}={\sf Z}_{3}^{2}$, ${\sf Aut}(F_{3})={\rm GL}_{2}(3)\cong{\rm Q}_{8}{:}{\rm S}_{3}$.
+However, ${\rm GL}_{2}(3)$ has no subgroup isomorphic to ${\sf Z}_{2}^{3}$, a contradiction.
 $\Box$
-Now we are ready to prove Theorem 1.1.
+Now, we are ready to prove Theorem 1.1.
 Proof of Theorem 1.1 :
 Let $G$ be a finite solvable group of which each Sylow subgroup has a cyclic or dihedral subgroup of prime index.
-Let $H$ be the largest normal Hall subgroup of odd order, with $\pi=\pi(H)$.
-Let $K$ be a Hall $\pi^{\prime}$-subgroup of $G$, where $\pi^{\prime}=\pi(G)\setminus\pi(H)$.
-Then we have factorization
+Then, the Sylow subgroups $G_{2}$ and $G_{p}$ ($p>2$) are respectively given in Proposition 4.1 and Lemma 4.3.
+Let $H\leqslant G$ be the largest normal Hall subgroup of odd order, with $\pi=\pi(H)$. Let $K$ be a Hall $\pi^{\prime}$-subgroup of $G$, where $\pi^{\prime}=\pi(G)\setminus\pi(H)$. Then, we have :
 G=H:K.G=H{:}K.
-Since $H$ is of odd order, by Lemma 5.4, H=A:BH=A{:}B such that $A$ is abelian, $B$ is nilpotent, and $A\cap{\bf Z}(H)=1$.
-We then consider the subgroup $K$.
-Let $p$ be the largest prime divisor of $|G|$.
-If $p>7$, then by Lemmas 5.1 and 5.6, $G_{p}$ is normal in both $G_{2^{\prime}}$ and $G_{\{2,p\}}$, and so $G_{p}$ is normal in $G$, and $p\in\pi$.
-Then, by inductively considering $G/G_{p}$ we deduce that any prime divisor $r$ of $|G|$ with $r>7$ lies in $\pi$.
-That is, $\pi^{\prime}\subseteq\{2,3,5,7\}$.
-Further, by Lemmas 5.1, 5.2 and 5.6, $G_{5}$ is normalized by $G_{3}$, $G_{7}$ and $G_{2}$, respectively.
-Hence $5\notin\pi^{\prime}$.
-Thus, either
+Since $H$ is of odd order, by Lemma 5.3, H=A:BH=A{:}B such that $A$ is abelian, $B$ is nilpotent, and $A\cap{\bf Z}(H)=1$. In particular, if $G$ is of odd order, the $G=H$, $K=1$.
+We then assume that $G$ is of even order, so $2\in\pi^{\prime}$.
+For any prime divisor $p$ of $|G|$ with $p>7$ or $p=5$, by Lemma 5.1, $G_{p}$ is normalized by $G_{3}$ and $G_{7}$, while by Proposition 5.4, $G_{p}$ is normalized by $G_{2}$.
+Thus, $G=G_{\{2,3,7\}^{\prime}}{:}G_{\{2,3,7\}}$.
+Note that, $HG_{\{2,3,7\}^{\prime}}=G_{\pi\,\cup\,\{2,3,7\}^{\prime}}$ is also a normal Hall subgroup of odd order, so by our assumption, $\{2,3,7\}^{\prime}\subseteq\pi$.
+That is, $\pi^{\prime}\subseteq\{2,3,7\}$, and one of the following holds :
   • (i)
-    $\pi^{\prime}=\emptyset$, $G=H$ is of odd order; or
+    $\pi^{\prime}=\{2\}$, $K=G_{2}$ ;
   • (ii)
-    $\pi^{\prime}=\{2\}$, $\{2,3\}$, $\{2,7\}$ or $\{2,3,7\}$.
-At last, respectively for $\pi^{\prime}=\{2\}$, $\pi^{\prime}=\{2,3\}$ or $\{2,7\}$, and $\pi^{\prime}=\{2,3,7\}$, the group $K=G_{\pi^{\prime}}$ is described in Proposition 4.1, Lemma 5.6 and Lemma 5.7.
+    $\pi^{\prime}=\{2,3\}$ or $\{2,7\}$, where, respectively, $K=G_{\{2,3\}}\neq G_{3}{:}G_{2}$ or $K=G_{\{2,7\}}\neq G_{7}{:}G_{2}$, so $K$ is given in Proposition 5.4 ;
+  • (iii)
+    $\pi^{\prime}=\{2,3,7\}$, $K=G_{\{2,3,7\}}$ with neither $G_{3}$ nor $G_{7}$ normal in $K$, so $K$ is given in Lemma 5.5.
+Then, the proof is complete. $\Box$
+A proof of Proposition 5.4 is given in the remaining part of this section :
+Let $K$ be a $\{2,p\}$-group with $p$ odd prime satisfying Hypotheses 2.2.
+Let $F={\rm Fit}(K)=F_{2}\times F_{p}$.
+Assume that $K_{p}$ is not normal in $K$, so $K\neq K_{p}{:}K_{2}$, and $F_{p}<K_{p}$.
+We first prove $p\in\{3,7\}$ and determine the group $\overline{K}:=K/F_{p}$.
+The group $\overline{K}$ is also a $\{2,p\}$-group as $F_{p}<K_{p}$, and meanwhile, $\overline{K}$ has no normal $p$-subgroups, so ${\rm Fit}(\overline{K})={\bf O}_{2}(\overline{K})$, and
+$H:=\overline{K}/{\rm Fit}(\overline{K})\lesssim{\sf Out}({\rm Fit}(\overline{K}))={\sf Out}({\bf O}_{2}(\overline{K})).$
+Now, $|\overline{K}|$ is divisible by $p$, so are $|H|$ and $|{\sf Out}({\bf O}_{2}(\overline{K}))|$.
+Thus, by Proposition 4.1, the triple $({\bf O}_{2}(\overline{K}),{\sf Aut}({\bf O}_{2}(\overline{K})),{\sf Out}({\bf O}_{2}(\overline{K})))$ is one of the following :
+$({\sf Z}_{2}^{2},{\rm S}_{3},{\rm S}_{3}),\,({\sf Z}_{2}^{3},{\rm GL}_{3}(2),{\rm GL}_{3}(2)),\,({\rm Q}_{8},{\rm S}_{4},{\rm S}_{3}),\,({\rm Q}_{8}\circ{\sf Z}_{4},{\rm S}_{4}\times{\sf Z}_{2},{\rm S}_{3}\times{\sf Z}_{2})$,
+where ${\sf Out}({\bf O}_{2}(\overline{K}))={\sf Aut}({\bf O}_{2}(\overline{K}))/{\sf Inn}({\bf O}_{2}(\overline{K}))\lesssim{\sf Aut}({\bf O}_{2}(\overline{K}))$ can be regarded as a subgroup.
+Note that, $H$ is a subgroup of ${\sf Out}({\bf O}_{2}(\overline{K}))$, and meanwhile, $H=\overline{K}/{\bf O}_{2}(\overline{K})$, so it has no normal 2-subgroups.
+By checking the subgroups of ${\sf Out}({\bf O}_{2}(\overline{K}))$, we conclude that, either $p=7$, $H={\sf Z}_{7}$, $\overline{K}={\sf Z}_{2}^{3}{:}{\sf Z}_{7}$; or $p=3$, $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$, and one of the cases in Table 2 holds.
+Table 2. $\overline{K}=K/F_{3}$
+${\bf O}_{2}(\overline{K})$ ${\sf Z}_{2}^{2}$ ${\sf Z}_{2}^{3}$ ${\rm Q}_{8}$ ${\sf Z}_{4}\circ{\rm Q}_{8}$
+$\overline{K}={\bf O}_{2}(\overline{K}){:}{\sf Z}_{3}$ ${\rm A}_{4}$ ${\sf Z}_{2}\times{\rm A}_{4}$ ${\rm SL}_{2}(3)$ ${\sf Z}_{4}\circ{\rm SL}_{2}(3)$
+$\overline{K}={\bf O}_{2}(\overline{K}){:}{\rm S}_{3}$ ${\rm S}_{4}$ ${\sf Z}_{2}\times{\rm S}_{4}$ ${\rm GL}_{2}(3)$ ${\sf Z}_{4}\circ{\rm GL}_{2}(3)$
+We remark that, the groups in Table 2 satisfy Hypothesis 2.2.
+In the case $H={\sf Z}_{3}$, this is natural as the Sylow 2-subgroup $\overline{K}_{2}={\bf O}_{2}(\overline{K})$ comes from Proposition 4.1.
+Meanwhile, in the case $H={\rm S}_{3}$, we have $\overline{K}_{2}={\bf O}_{2}(\overline{K}){:}{\sf Z}_{2}$, and respectively for ${\bf O}_{2}(\overline{K})={\sf Z}_{2}^{2}$, ${\sf Z}_{2}^{3}$, ${\rm Q}_{8}$ or ${\sf Z}_{4}\circ{\rm Q}_{8}$, it is verified by simple calculation that $\overline{K}_{2}$ has a index 2 subgroup isomorphic to ${\sf Z}_{4}$, ${\rm D}_{8}$, ${\sf Z}_{8}$ or ${\rm D}_{16}$.
+For convenience, we then consider $K$ in three separate cases.
+Note that, $F_{2}\lhd K$, $F_{2}\cong\overline{F_{2}}\lhd\overline{K}$, so $F_{2}\lesssim{\bf O}_{2}(\overline{K})$.
+We thus have cases as follows :
+(I) $F_{2}={\bf O}_{2}(\overline{K})$, $H\in\{{\sf Z}_{7},{\sf Z}_{3}\}$ ; (II) $F_{2}={\bf O}_{2}(\overline{K})$, $H={\rm S}_{3}$ ; (III) $F_{2}<{\bf O}_{2}(\overline{K}).$
+Note that, (I) exactly corresponds to the case $F_{2}=K_{2}\lhd K$.
+In this case, either $p=7$, $K={\sf Z}_{2}^{3}{:}K_{7}$ is with $K/F_{7}\cong{\sf Z}_{2}^{3}{:}{\sf Z}_{7}$; or $p=3$, $K_{2}\in\{{\sf Z}_{2}^{2},{\sf Z}_{2}^{3},{\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8}\}$, and $K=K_{2}{:}K_{3}$ is with $K/F_{3}$ equal to one of the groups in Table 2, so $K$ is one of
+${\sf Z}_{2}^{2}{:}K_{3},\ {\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}K_{3}),\ {\rm Q}_{8}{:}K_{3},\ {\sf Z}_{4}\circ_{{\sf Z}_{2}}({\rm Q}_{8}{:}K_{3}),$
+where ${\sf Z}_{4}\circ_{{\sf Z}_{2}}({\rm Q}_{8}{:}K_{3})$ cannot be simply written as ${\sf Z}_{4}\circ({\rm Q}_{8}{:}K_{3})$, as we might have ${\bf Z}({\rm Q}_{8}{:}K_{3})>{\sf Z}_{2}$.
+These groups are given in Proposition 5.4 (1).
+We then dealt with (II) and (III) in Lemmas 5.6 and 5.7 - 5.9, respectively.
+Lemma 5.6.
+If it is in case (II), then $K$ is one of the groups : $X$, ${\sf Z}_{2}\times X$, $Y$, ${\sf Z}_{4}\circ_{{\sf Z}_{2}}Y$, where $X,Y$ are given in Table 1, $X$ is homomorphic to ${\rm S}_{4}$, and $Y$ is homomorphic to ${\rm GL}_{2}(3)$.
+Proof.
+In case (II), we have :
+$F_{2}={\bf O}_{2}(\overline{K})\in\{{\sf Z}_{2}^{2},\,{\sf Z}_{2}^{3},\,{\rm Q}_{8},\,{\sf Z}_{4}\circ{\rm Q}_{8}\}$, $H={\rm S}_{3}$, and $K=F_{3}.\overline{K}=F_{3}.(F_{2}{:}{\rm S}_{3})$.
+Let $L\leqslant K$ be the full preimage of ${\rm S}_{3}\leqslant\overline{K}$.
+Then, $\langle F_{2},L\rangle=F_{2}L$ is the full preimage of $\overline{K}$, so $F_{2}L=K$.
+Meanwhile, $|F_{2}||L|=|F_{2}||F_{3}||{\rm S}_{3}|=|K|$, so $F_{2}\cap L=1$.
+Thus, $K=F_{2}{:}L$ is a semidirect product, which satisfies :
+${\bf C}_{L}(F_{2})=F_{3}$, and $K/{\bf C}_{L}(F_{2})=\overline{K}=F_{2}{:}{\rm S}_{3}$ as given in Table 2,
+so that, $K$ is determined once $F_{2}$, $L$ and $F_{3}$ are all determined.
+Note that, $L$ has a normal subgroup $F_{3}$ such that $L/F_{3}\cong{\rm S}_{3}$.
+It follows that the Sylow subgroup $L_{3}$ of $L$ is normal, and $L=L_{3}{:}{\sf Z}_{2}$.
+Thus, $L$ is one of the groups given in Corollary 4.4 (1), by checking which we conclude that, there are six possibilities for the pair $(L,\,F_{3})$ as follows, where $\ell\geqslant 1$ :
+$({\rm D}_{2.3^{\ell}},\,{\sf Z}_{3^{\ell-1}})$, $({\sf Z}_{3^{\ell}}\times{\rm D}_{6},\,{\sf Z}_{3^{\ell}}),\ \ ({\rm D}_{2.3^{\ell}}\times{\sf Z}_{3},\,{\sf Z}_{3^{\ell-1}}\times{\sf Z}_{3})$,
+$(({\sf Z}_{3^{\ell}}\times{\sf Z}_{3}){:}{\sf Z}_{2},\,{\sf Z}_{3^{\ell}}),\ \ (({\sf Z}_{3^{\ell}}\times{\sf Z}_{3}){:}{\sf Z}_{2},\,{\sf Z}_{3^{\ell-1}}\times{\sf Z}_{3})$  and  $({\sf Z}_{3^{\ell+1}}{:}{\sf Z}_{6},\,{\sf Z}_{3^{\ell}}\times{\sf Z}_{3})$.
+(1) Suppose $F_{2}={\sf Z}_{2}^{2}$, so $K={\sf Z}_{2}^{2}{:}L$ satisfies that $K/F_{3}={\sf Z}_{2}^{2}{:}{\rm S}_{3}\cong{\rm S}_{4}$.
+Then, respectively for the six possibilities, we obtain the following groups, which are collected in Table 1 as $X$ :
+${\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}}$, ${\sf Z}_{3^{\ell}}\times{\rm S}_{4},\ \ ({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\sf Z}_{3},$
+$({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\ \ (({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})\times{\sf Z}_{3}){:}{\sf Z}_{2}$  and  ${\sf Z}_{2}^{2}{:}({\sf Z}_{3^{\ell+1}}{:}{\sf Z}_{6})$.
+(2) Suppose $F_{2}={\sf Z}_{2}^{3}$, so $K={\sf Z}_{2}^{3}{:}L$ satisfies that $K/F_{3}={\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\rm S}_{3})$.
+That is, $L$ centralizes some ${\sf Z}_{2}$, and normalizes its complement ${\sf Z}_{2}^{2}$.
+Thus, $K={\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}L)$, where ${\sf Z}_{2}^{2}{:}L=X$ is given above.
+(3) Suppose $F_{2}={\rm Q}_{8}$, so $K={\rm Q}_{8}{:}L$ satisfies that $K/F_{3}={\rm Q}_{8}{:}{\rm S}_{3}\cong{\rm GL}_{2}(3)$.
+Then, similar to (1), we obtain the following groups , which are collected in Table 1 as $Y$ :
+${\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}}$, ${\sf Z}_{3^{\ell}}\times{\rm GL}_{2}(3)$, $({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})\times{\sf Z}_{3}$,
+$\mbox{$({\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\sf Z}_{3})){:}{\sf Z}_{2}$,\ \ $(({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})\times{\sf Z}_{3}){:}{\sf Z}_{2}$ \,and\, ${\rm Q}_{8}{:}({\sf Z}_{3^{\ell+1}}{:}{\sf Z}_{6})$}.$
+(4) Suppose $F_{2}={\sf Z}_{4}\circ{\rm Q}_{8}$, so $K=({\sf Z}_{4}\circ{\rm Q}_{8}){:}L$ satisfies that $K/F_{3}={\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$.
+That is, $L$ centralizes ${\sf Z}_{4}$, and normalizes ${\rm Q}_{8}$.
+Thus, ${\sf Z}_{4}$ and ${\rm Q}_{8}{:}L=Y$ as given above are two normal subgroups of $K$ commuting with each other.
+Further, ${\sf Z}_{4}Y=K$, so $|{\sf Z}_{4}\cap Y|=2$, and $K={\sf Z}_{4}\circ_{{\sf Z}_{2}}Y$.
+Note that, ${\sf Z}_{4}\circ_{{\sf Z}_{2}}Y$ cannot be simply written as ${\sf Z}_{4}\circ Y$, as we might have ${\bf Z}(Y)>{\sf Z}_{2}$.
+$\Box$
+Lemma 5.7.
+If it is in case (III), then $p=3$, and $(F_{2},{\bf O}_{2}(\overline{K}))=({\sf Z}_{2}^{2},{\sf Z}_{2}^{3})$ or $({\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8})$.
+Proof.
+In case (III), we have $F_{2}\lhd\overline{K}={\bf O}_{2}(\overline{K}){:}H$ and $F_{2}<{\bf O}_{2}(\overline{K})$.
+If $p=7$, then $\overline{K}={\sf Z}_{2}^{3}{:}{\sf Z}_{7}$ has ${\bf O}_{2}(\overline{K})={\sf Z}_{2}^{3}$ as a minimal normal subgroup, and so $F_{2}=1$. That is, $F=F_{7}$, and
+${\sf Z}_{2}^{3}{:}{\sf Z}_{7}=\overline{K}=K/F_{7}\lesssim{\sf Out}(F_{7}).$
+It is impossible, as by Lemma 4.3, $|{\sf Aut}(F_{7})|$ being divisible by 8 only happens when $F_{7}={\sf Z}_{7}^{2}$, ${\sf Out}(F_{7})={\rm GL}_{2}(7)$.
+However, ${\rm GL}_{2}(7)$ has no subgroup isomorphic to ${\sf Z}_{2}^{3}$.
+Thus, $p=3$, $H\in\{{\sf Z}_{3},{\rm S}_{3}\}$, and $\overline{K}={\bf O}_{2}(\overline{K}){:}H$ is one of the groups in Table 2 :
+${\sf Z}_{2}^{2}{:}H,\ {\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}H),\ {\rm Q}_{8}{:}H,\ {\sf Z}_{4}\circ({\rm Q}_{8}{:}H).$
+Note that, ${\sf Z}_{2}^{2}{:}H$ has ${\sf Z}_{2}^{2}$ as a minimal normal subgroup. Thus,
+    if $\overline{K}={\sf Z}_{2}^{2}{:}H$, then $F_{2}=1$ ;
+    if $\overline{K}={\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}H)=\langle a\rangle\times(\langle b,c\rangle{:}H)$, then $F_{2}$ is one of $1$, $\langle a\rangle={\sf Z}_{2}$, $\langle b,c\rangle={\sf Z}_{2}^{2}$.
+Also note that, ${\rm Q}_{8}{:}H$ only has the center ${\sf Z}_{2}$ and ${\rm Q}_{8}$ as normal $2$-subgroups. Thus,
+    if $\overline{K}={\rm Q}_{8}{:}H$, then $F_{2}=1$ or ${\sf Z}_{2}$ ;
+    if $\overline{K}={\sf Z}_{4}\circ({\rm Q}_{8}{:}H)$, then either $F_{2}\leqslant{\rm Q}_{8}$, so $F_{2}=1,{\sf Z}_{2},{\rm Q}_{8}$ ; or $F_{2}{\rm Q}_{8}={\sf Z}_{4}\circ{\rm Q}_{8}$, implying that $F_{2}={\sf Z}_{4}$.
+In conclusion, $(F_{2},{\bf O}_{2}(\overline{K}))$ with $F_{2}<{\bf O}_{2}(\overline{K})$ is one of the pairs listed as follows :
+  • (a)
+    $(1,{\sf Z}_{2}^{2})$, $(1,{\sf Z}_{2}^{3})$, $({\sf Z}_{2},{\sf Z}_{2}^{3})$, $({\sf Z}_{2},{\rm Q}_{8})$, $({\sf Z}_{2},{\sf Z}_{4}\circ{\rm Q}_{8})$, $({\sf Z}_{4},{\sf Z}_{4}\circ{\rm Q}_{8})$ ;
+  • (b)
+    $(1,{\rm Q}_{8})$, $(1,{\sf Z}_{4}\circ{\rm Q}_{8})$ ;
+  • (c)
+    $({\sf Z}_{2}^{2},{\sf Z}_{2}^{3})$, $({\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8})$.
+We thus need to exclude the pairs in (a) and (b). For the pairs in (a), we have :
+$\overline{K}/F_{2}=({\bf O}_{2}(\overline{K}){:}H)/F_{2}={\sf Z}_{2}^{2}{:}H$ or ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}H),$
+which contains a subgroup ${\sf Z}_{2}^{2}{:}{\sf Z}_{3}\cong{\rm A}_{4}$. Meanwhile, $F=F_{3}\times F_{2}$, so
+$K/F\cong\overline{K}/F_{2}\lesssim{\sf Out}(F)={\sf Out}(F_{3})\times{\sf Out}(F_{2}),$
+where ${\sf Out}(F_{2})=1$ or ${\sf Z}_{2}$.
+Thus, ${\rm A}_{4}$ is embedded into either ${\sf Out}(F_{3})$ or ${\sf Out}(F_{3})\times{\sf Z}_{2}$.
+For the latter, as ${\rm A}_{4}\cap{\sf Z}_{2}\lhd{\rm A}_{4}$, ${\rm A}_{4}\cap{\sf Z}_{2}=1$, so ${\rm A}_{4}$ is also embedded into ${\sf Out}(F_{3})$.
+We thus get a contradiction by Corollary 4.4 (3).
+Hence this case cannot occur.
+(b.1) Suppose $(F_{2},{\bf O}_{2}(\overline{K}))=(1,{\rm Q}_{8})$.
+Then, $F=F_{3}$, $K/F_{3}={\rm Q}_{8}{:}H\lesssim{\sf Out}(F_{3})$.
+By lemma 4.3, $|{\sf Out}(F_{3})|$ being divisible by $8$ only happens when
+$F_{3}={\sf Z}_{3}^{2},\ \ {\sf Out}(F_{3})={\rm GL}_{2}(3)\cong{\rm Q}_{8}{:}{\rm S}_{3}.$
+Since $H={\sf Z}_{3}$ or ${\rm S}_{3}$, $K_{3}>F_{3}$, and in particular, $K_{3}$ acts on $F_{3}$ non-trivially, so $K_{3}$ is non-abelian as given in Lemma 4.3 (3), written explicitly as
+$K_{3}=\langle a\rangle{:}\langle b\rangle={\sf Z}_{9}{:}{\sf Z}_{3}$, with $a^{b}=a^{4}$, $b^{a}=a^{6}b$.
+By simple calculation, any element lying in $K_{3}\setminus\langle a^{3},b\rangle$ has order $9$, so $F_{3}=\langle a^{3}\rangle\times\langle b\rangle$.
+Let $c$ be the central involution of $K_{2}=({\rm Q}_{8}{:}H)_{2}={\rm Q}_{8}$ or ${\rm Q}_{8}{:}{\sf Z}_{2}$.
+Then $\overline{c}$ is the central involution of $\overline{K}\lesssim{\rm GL}_{2}(3)$, so $\overline{c}$ acts on $F_{3}$ by mapping each $x$ to $x^{-1}$.
+That is, for any for any $x\in F_{3}$, we have
+$x^{\overline{c}}=x^{c}=x^{-1}$, so $c^{x}=cx^{2}$.
+Meanwhile, $\overline{c}\in{\bf Z}(\overline{K})$, $\overline{c}^{\overline{a}}=\overline{c}$, and so $c^{a}=cy$ with some $y=a^{3i}b^{j}\in F_{3}$. Thus, for $a^{3}\in F_{3}$, we have
+$c^{a^{3}}=(ca^{3i}b^{j})^{a^{2}}=(ca^{(6i+6j)}b^{2j})^{a}=ca^{(9i+18j)}b^{3j}=c\neq ca^{6},$
+a contradiction. Hence this case cannot occur.
+(b.2) Suppose $(F_{2},{\bf O}_{2}(\overline{K}))=(1,{\sf Z}_{4}\circ{\rm Q}_{8})$.
+Then, $F=F_{3}$, $\overline{K}={\sf Z}_{4}\circ({\rm Q}_{8}{:}H)\lesssim{\sf Out}(F_{3})$.
+Similar to (b.1), $|{\sf Out}(F_{3})|$ being divisible by $8$ implies $F_{3}={\sf Z}_{3}^{2}$, ${\sf Out}(F_{3})={\rm GL}_{2}(3)$.
+By comparing the orders, we have $H={\sf Z}_{3}$ and ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\sf Z}_{3})={\rm GL}_{2}(3)$, which is impossible as ${\bf Z}({\rm GL}_{2}(3))={\sf Z}_{2}\neq{\sf Z}_{4}$.
+Hence this case cannot occur.
+$\Box$
+Lemma 5.8.
+If $(F_{2},{\bf O}_{2}(\overline{K}))=({\sf Z}_{2}^{2},{\sf Z}_{2}^{3})$, then $K=W$ is given in Table 1 (the left side), homomorphic to ${\rm A}_{4}$ or ${\rm S}_{4}$.
+Proof.
+The proof is similar to Lemma 5.6. In this case, we have :
+$K=F_{3}.\overline{K}=F_{3}.({\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}H))=F_{3}.({\sf Z}_{2}^{2}{:}({\sf Z}_{2}\times H)),\mbox{\ where\ }F_{2}={\sf Z}_{2}^{2}.$
+Let $L\leqslant K$ the full preimage of ${\sf Z}_{2}\times H\leqslant\overline{K}$.
+Then, $\langle F_{2},L\rangle=F_{2}L$ is the full preimage of $\overline{K}$, so $F_{2}L=K$.
+Meanwhile, $|F_{2}||L|=|F_{2}||F_{3}|\cdot 2|H|=|K|$, so $F_{2}\cap L=1$.
+Thus, $K=F_{2}{:}L={\sf Z}_{2}^{2}{:}L$ is a semidirect product, which satisfies :
+${\bf C}_{L}({\sf Z}_{2}^{2})=F_{3}{:}{\sf Z}_{2}$, and $K/{\bf C}_{L}({\sf Z}_{2}^{2})\cong\overline{K}/{\sf Z}_{2}={\sf Z}_{2}^{2}{:}H\cong{\rm A}_{4}$ or ${\rm S}_{4}$,
+so that, $K$ is determined once the pair $(L,F_{3})$ is determined.
+(1) Suppose $H={\sf Z}_{3}$.
+Then, $L$ has a normal subgroup $F_{3}$ such that $L/F_{3}={\sf Z}_{6}$.
+It follows that the Sylow subgroup $L_{3}$ of $L$ is normal, and $L=L_{3}{:}{\sf Z}_{2}$.
+Further, $L\neq L_{3}\times{\sf Z}_{2}$; otherwise we can get a normal 2-subgroup of $K$ larger than $F_{2}$.
+Thus, $L$ is one of the groups given in Corollary 4.4 (1), by checking which we conclude that, there are three possibilities for the pair $(L,F_{3})$ as follows, where $\ell\geqslant 1$ :
+$({\rm D}_{2.3^{\ell}}\times{\sf Z}_{3},\,{\sf Z}_{3^{\ell}})$, $({\sf Z}_{3^{\ell}}\times{\rm D}_{6},\,{\sf Z}_{3^{\ell-1}}\times{\sf Z}_{3})$, $({\sf Z}_{3^{\ell+1}}{:}{\sf Z}_{6},\,{\sf Z}_{3^{\ell+1}})$,
+which, respectively, provide groups $K={\sf Z}_{2}^{2}{:}L$ with $K/({F_{3}{:}{\sf Z}_{2}})\cong{\sf Z}_{2}^{2}{:}{\sf Z}_{3}={\rm A}_{4}$ as follows :
+${\rm D}_{2.3^{\ell}}\times{\rm A}_{4}$, $({\sf Z}_{2}^{2}{:}{{\sf Z}_{3^{\ell}}})\times{\rm D}_{6}$, $({\sf Z}_{2}^{2}\times{\rm D}_{2.3^{\ell+1}}){:}{\sf Z}_{3}$.
+(2) Suppose $H={\rm S}_{3}$.
+Then, $L$ has a normal subgroup $F_{3}$ such that $L/F_{3}={\sf Z}_{2}\times{\rm S}_{3}$, and so $L=L_{3}{:}({\sf Z}_{2}\times{\sf Z}_{2})$.
+Further, ${\bf C}_{{\sf Z}_{2}\times{\sf Z}_{2}}(L_{3})=1$; otherwise we can get a normal 2-subgroup of $K$ larger than $F_{2}$.
+Thus, by Corollary 4.4 (2), $L={\rm D}_{2.3^{\ell}}\times{\rm D}_{6}$ with $\ell\geqslant 1$, and $F_{3}={\sf Z}_{3^{\ell}}$ or ${\sf Z}_{3^{\ell-1}}\times{\sf Z}_{3}$.
+This provides groups $K={\sf Z}_{2}^{2}{:}L$ with $K/({F_{3}{:}{\sf Z}_{2}})\cong{\sf Z}_{2}^{2}{:}{\rm S}_{3}={\rm S}_{4}$ as follows :
+${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$  or  $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$.
+$\Box$
+Lemma 5.9.
+If $(F_{2},{\bf O}_{2}(\overline{K}))=({\rm Q}_{8},{\sf Z}_{4}\circ{\rm Q}_{8})$, then $K=W$ is given in Table 1 (the right side), homomorphic to ${\sf Z}_{4}\circ{\rm SL}_{2}(3)$ or ${\sf Z}_{4}\circ{\rm GL}_{2}(3)$.
+Proof.
+In this case, we have :
+$K=F_{3}.(({\sf Z}_{4}\circ{\rm Q}_{8}){:}H)=F_{3}.({\sf Z}_{4}\circ({\rm Q}_{8}{:}H)),$
+so $K$ is homomorphic to ${\sf Z}_{4}\circ{\rm SL}_{2}(3)$ or ${\sf Z}_{4}\circ{\rm GL}_{2}(3)$, respectively for $H={\sf Z}_{3}$ or ${\rm S}_{3}$.
+Let $L\leqslant K$ be the full preimage of ${\sf Z}_{4}\times H\leqslant\overline{K}$. Then, $\langle F_{2},L\rangle=F_{2}L$ is the full preimage of $\overline{K}$, so $K=F_{2}L$.
+(1) Suppose $H={\sf Z}_{3}$.
+Then, $L$ has a normal subgroup $F_{3}$ with $L/F_{3}={\sf Z}_{12}$.
+It follows that the Sylow subgroup $L_{3}$ is normal, and $L=L_{3}{:}{\sf Z}_{4}$.
+Let $x\in{\bf C}_{{\sf Z}_{4}}(L_{3})$, so $x\in{\bf Z}(L)$, and $F_{2}\langle x\rangle$ is a normal 2-subgroup of $K$.
+It implies $|x|<4$, and so ${\bf C}_{{\sf Z}_{4}}(L_{3})={\sf Z}_{4}\cap{\rm Q}_{8}={\sf Z}_{2}$.
+That is, $L/{\bf C}_{{\sf Z}_{4}}(L_{3})=L{:}{\sf Z}_{2}$ is one of the groups given in Corollary 4.4 (1), by checking which we conclude that, there are three possibilities for the pair $(L,F_{3})$ as follows, where $\ell\geqslant 1$.
+(1.1) $(L,\,F_{3})=(({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\times{\sf Z}_{3},\,{\sf Z}_{3^{\ell}})$.
+Let $M={\sf Z}_{3^{\ell}}{:}{\sf Z}_{4}$, $N=\langle F_{2},{\sf Z}_{3}\rangle={\rm Q}_{8}{:}{\sf Z}_{3}$.
+Then, $M$, NN are two normal subgroups of $K$, which generate $K$ and commutes with each other.
+In particular, $|M\cap N|=2$, so $M\cap N={\sf Z}_{2}={\sf Z}_{4}\cap{\rm Q}_{8}={\sf Z}(N)$, and
+$K=M\circ N=({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\sf Z}_{3}).$
+(1.2) $(L,\,F_{3})=({\sf Z}_{3^{\ell}}\times({\sf Z}_{3}{:}{\sf Z}_{4}),\,{\sf Z}_{3^{\ell-1}}\times{\sf Z}_{3})$. Similar to (1.1), we have :
+$K=({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})\circ({\sf Z}_{3}{:}{\sf Z}_{4}).$
+(1.3) $(L,F_{3})=({\sf Z}_{3^{\ell+1}}{:}{\sf Z}_{12},{\sf Z}_{3^{\ell+1}})$.
+Let $M={\sf Z}_{3^{\ell+1}}{:}{\sf Z}_{4}$.
+Then, $F_{2}={\rm Q}_{8}$ and $M$ are two normal subgroups of $K$ commuting with each other.
+Moreover, ${\rm Q}_{8}M$ is normal in $K$, with a complementary subgroup ${\sf Z}_{3}\leqslant{\sf Z}_{12}$, so $K=({\rm Q}_{8}M){:}{\sf Z}_{3}$.
+In particular, $|{\rm Q}_{8}\cap M|=2$, ${\rm Q}_{8}\cap M={\sf Z}_{2}={\rm Q}_{8}\cap{\sf Z}_{4}={\bf Z}({\rm Q}_{8})$, so ${\rm Q}_{8}M={\rm Q}_{8}\circ M$, and
+$K=({\rm Q}_{8}\circ({\sf Z}_{3^{\ell+1}}{:}{\sf Z}_{4})){:}{\sf Z}_{3}.$
+(2) Suppose $H={\rm S}_{3}$.
+Then, similar to (1), it is shown that, $L=L_{3}{:}({\sf Z}_{4}\times{\sf Z}_{2})$ with ${\bf C}_{{\sf Z}_{4}\times{\sf Z}_{2}}(L_{3})={\sf Z}_{4}\cap F_{2}={\sf Z}_{2}$, so by Corollary 4.4 (2), $L/{\bf C}_{{\sf Z}_{4}\times{\sf Z}_{2}}(L_{3})=L{:}({\sf Z}_{2}\times{\sf Z}_{2})={\rm D}_{2.3^{\ell}}\times{\rm D}_{6},$ where $\ell\geqslant 1$.
+Note that, $L/F_{3}={\sf Z}_{4}\times{\rm S}_{3}$.
+We thus have two possibilities :
+$(L,\,F_{3})=(({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\times{\rm D}_{6},\,{\sf Z}_{3^{\ell}})\mbox{\ or\ }({\rm D}_{2.3^{\ell}}\times{\rm D}_{6},\,{\sf Z}_{3^{\ell-1}}\times{\sf Z}_{3}).$
+This, similar to (1.1), (1.2), provides groups $K$ as follows :
+$({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\rm S}_{3})$  or  $({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$.
 $\Box$
 6. Proof of Theorem 1.2
-In this section we focus on the arc-transitive automorphism groups of maps. The following lemma shows that the class of such groups is closed under taking quotients.
+In this section, we focus on the arc-transitive automorphism groups of maps. As the following lemma shows, the class of such groups is nearly closed under taking quotients.
 Lemma 6.1.
@@
 Let $(x,y,z)$ be a regular triple for $G$.
-If $\overline{x},\overline{y},\overline{z}\in\overline{G}$ are all involutions, and $\overline{x}\neq\overline{z}$, then they form a regular triple for $\overline{G}$, since $\overline{G}=\langle\overline{x},\overline{y},\overline{z}\rangle$ and $\overline{x}\,\overline{z}=\overline{z}\,\overline{x}$.
-Otherwise, either some of $\overline{x},\overline{y},\overline{z}$ equals the identity-element in $\overline{G}$, or $\overline{x}=\overline{z}$.
-In both cases, $\overline{G}$ is generated by one or two involutions and then it is ${\sf Z}_{2}$ or dihedral.
+Then, $G=\langle x,y,z\rangle$ and $\langle x,z\rangle={\sf Z}_{2}^{2}$, so $\overline{G}=\langle\overline{x},\overline{y},\overline{z}\rangle$, $\overline{x}\,\overline{z}=\overline{z}\,\overline{x}$.
+Now, if $\overline{x},\overline{y},\overline{z}$ are involutions, and $\overline{x}\neq\overline{z}$, then $(\overline{x},\overline{y},\overline{z})$ is a regular triple for $\overline{G}$.
+Otherwise, either some of $\overline{x},\overline{y},\overline{z}$ equals the identity-element $\overline{1}$, or $\overline{x}=\overline{z}$, where, in both cases, $\overline{G}$ is generated by at most two involutions, so it is ${\sf Z}_{2}$ or dihedral.
+In particular, $\overline{G}$ is always of even order.
 It is similar to prove the remaining part of this lemma. $\Box$
 Lemma 6.2.
-Let $G={\rm Q}_{8}{:}{\rm S}_{3}\cong{\rm GL}(2,3)$. Then neither $G$ nor ${\sf Z}_{4}\circ G$ has a regular triple.
+Let $G={\rm GL}(2,3)$. Then neither $G$ nor ${\sf Z}_{4}\circ G$ has a regular triple.
 Proof.
-Suppose that $G$ has a regular triple $(x,y,z)$.
-Since $xz=zx$, we have $\langle x,z\rangle\cong{\sf Z}_{2}^{2}$, and so $\{x,z,xz\}\cap{\rm Q}_{8}\not$$\emptyset$.
-That is, one of $x,z,xz$ has to be the central involution 1-1 of $G$.
-It follows that $\overline{G}=G/\langle-1\rangle=\langle\overline{x},\overline{y}\rangle$ or $\langle\overline{y},\overline{z}\rangle$ is a dihedral group, which is impossible as $\overline{G}={\sf Z}_{2}^{2}{:}{\rm S}_{3}$.
-Suppose that $K={\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ has a regular triple $(x,y,z)$. Write $K$ as
+Suppose that $G\cong{\rm Q}_{8}{:}{\rm S}_{3}$ has a regular triple $(x,y,z)$.
+Then, $\langle x,z\rangle={\sf Z}_{2}^{2}$, so $\{x,z,xz\}\cap{\rm Q}_{8}\not$$\emptyset$, namely, one of $x,z,xz$ is the central involution 1-1 of $G$.
+It follows that $\overline{G}=G/\langle-1\rangle=\langle\overline{x},\overline{y}\rangle$ or $\langle\overline{y},\overline{z}\rangle$ is dihedral, and this contradicts $\overline{G}\cong{\sf Z}_{2}^{2}{:}{\rm S}_{3}={\rm S}_{4}$.
+Suppose that $K={\sf Z}_{4}\circ G\cong{\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ has a regular triple $(x,y,z)$. Write $K$ as
 $K=\langle u\rangle\circ({\rm Q}_{8}{:}(\langle\sigma\rangle{:}\langle\tau\rangle)),$
-where $|u|=4$, ${\rm Q}_{8}=\pm\{1,i,j,k\}$ with $i^{2}=j^{2}=k^{2}=ijk=-1$, and $\sigma,\tau\in{\sf Aut}({\rm Q}_{8})$ such that
-$\sigma:i\rightarrow j\rightarrow k\,(\rightarrow i),\mbox{\,\ and\ \,}\tau:i\leftrightarrow-j,j\leftrightarrow-i,k\leftrightarrow-k.$
-Then the subgroup $\langle\sigma\rangle{:}\langle\tau\rangle\cong{\rm S}_{3}$ contains three involutions, i.e., $\tau_{k}=\tau$,
-$\tau_{i}:j\leftrightarrow-k,k\leftrightarrow-j,i\leftrightarrow-i,\mbox{\ and\ }\tau_{j}:i\leftrightarrow-k,k\leftrightarrow-i,j\leftrightarrow-j.$
-By calculation, we conclude that $K$ contains the following involutions:
+where $|u|=4$, ${\rm Q}_{8}=\pm\{1,i,j,k\}$ with $i^{2}=j^{2}=k^{2}=ijk=-1$, and $\sigma,\tau\in{\sf Aut}({\rm Q}_{8})$ act on ${\rm Q}_{8}$ as follows :
+$\sigma:i\rightarrow j\rightarrow k\,(\rightarrow i)\mbox{\,\ and\,\ }\tau:i\leftrightarrow-j,j\leftrightarrow-i,k\leftrightarrow-k.$
+Then, $\langle\sigma\rangle{:}\langle\tau\rangle={\rm S}_{3}$ contains three involutions, i.e., $\tau_{k}:=\tau$, and $\tau_{i}$, $\tau_{j}$ such that
+$\tau_{i}:j\leftrightarrow-k,k\leftrightarrow-j,i\leftrightarrow-i\mbox{\,\ and\,\ }\tau_{j}:i\leftrightarrow-k,k\leftrightarrow-i,j\leftrightarrow-j.$
+By direct calculation, we conclude that the involutions contained in $K$ are as follows :
 $\{-1\}\cup\{\pm u\lambda\}\cup\{\pm\tau_{\lambda},\pm\lambda\tau_{\lambda}\},\mbox{\ where\ }\lambda\in\{i,j,k\}.$
-We shall confirm the following claim:
 Claim. The only involutions commutes with $u\lambda$ (or $u\lambda$) are $\pm u\lambda$ and 1-1.
 In fact, for any $\lambda^{\prime}\in\{i,j,k\}$ with $\lambda\neq\lambda^{\prime}$, we have :
-$\lambda^{\lambda^{\prime}}=\lambda^{\tau_{\lambda}}=\lambda^{\lambda\tau_{\lambda}}=-\lambda$, and $\lambda^{\tau_{\lambda^{\prime}}}=-\lambda^{\prime\prime},\ \lambda^{\lambda^{\prime}\tau_{\lambda^{\prime}}}=-\lambda^{\tau_{\lambda^{\prime}}}=\lambda^{\prime\prime}$,
+$\lambda^{\lambda^{\prime}}=\lambda^{\tau_{\lambda}}=\lambda^{\lambda\tau_{\lambda}}=-\lambda$ and $\lambda^{\tau_{\lambda^{\prime}}}=-\lambda^{\prime\prime},\ \lambda^{\lambda^{\prime}\tau_{\lambda^{\prime}}}=-\lambda^{\tau_{\lambda^{\prime}}}=\lambda^{\prime\prime}$,
 where $\lambda^{\prime\prime}\in\{i,j,k\}\setminus\{\lambda,\lambda^{\prime}\}.$ As $u$ and 1-1 lie in the center of $K$, we then have :
@@
 $(\pm u\lambda)^{\pm\tau_{\lambda}}=\pm u(\lambda^{\pm\tau_{\lambda}})=\mp u\lambda,\ (\pm u\lambda)^{\pm\lambda\tau_{\lambda}}=\pm u(\lambda^{\pm\lambda\tau_{\lambda}})=\mp u\lambda,$
-$(\pm u\lambda)^{\pm\tau_{\lambda^{\prime}}}=\pm u(\lambda^{\pm\tau_{\lambda^{\prime}}})=\mp u\lambda^{\prime\prime},\ (\pm u\lambda)^{\pm\lambda^{\prime}\tau_{\lambda^{\prime}}}=\pm u(\lambda^{\pm\lambda^{\prime}\tau_{\lambda^{\prime}}})=\pm u\lambda^{\prime\prime}.$
+[…]
 We have therefore confirmed the claim.
 Let $H=\langle u,i,j,k\rangle\leqslant K$.
-Then $H\cong{\sf Z}_{4}\circ{\rm Q}_{8}$ is normal in $K$ and $K/H\cong{\rm S}_{3}$.
-Further, the involutions contained in $H$ are: $\{-1\}\cup\{\pm u\lambda\}$.
-Recall that $(x,y,z)$ is a regular triple for $G$.
-Thus, $\langle x,z\rangle\cong{\sf Z}_{2}^{2}$, and so $\langle x,z\rangle\cap H>1$, that is,
+Then, $H\cong{\sf Z}_{4}\circ{\rm Q}_{8}$ is normal in $K$, $K/H\cong{\rm S}_{3}$, and the involutions contained in $H$ lie in $\{-1\}\cup\{\pm u\lambda\}$.
+Now, $(x,y,z)$ is a regular triple for $G$ such that $\langle x,z\rangle={\sf Z}_{2}^{2}$, so $\langle x,z\rangle\cap H>1$, that is,
 $\{x,z,xz\}\cap(\{-1\}\cup\{\pm u\lambda\})\not$$\emptyset$.
-By the claim given above, one of $x,z,xz$ has to be 1-1.
-Thus, $\overline{K}=K/\langle-1\rangle=\langle\overline{x},\overline{y}\rangle$ or $\langle\overline{y},\overline{z}\rangle$ is a dihedral group, which is impossible as $\overline{K}=K/\langle-1\rangle={\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\rm S}_{3})$.
+By the above claim, one of $x,z,xz$ has to be 1-1.
+It follows that $\overline{K}=K/\langle-1\rangle=\langle\overline{x},\overline{y}\rangle$ or $\langle\overline{y},\overline{z}\rangle$ is dihedral, and this contradicts $\overline{K}\cong{\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\rm S}_{3})={\sf Z}_{2}\times{\rm S}_{4}$.
 $\Box$
@@
 Let $(\alpha,z)$ be a rotary pair for $G$.
+Then $G=\langle\alpha,z\rangle$ and $|z|=2$.
 Since $G$ is not dihedral, $|\alpha|\neq 2$, so $\alpha\in\langle a,b\rangle$ and $\langle\alpha\rangle$ is normal in $G$.
@@
 $\Box$
-Now we are ready to prove Theorem 1.2.
+Now, we are ready to prove Theorem 1.2.
 Proof of Theorem 1.2 :
 Let $\mathcal{M}$} be a map with square-free Euler characteristic, and let $G\leqslant{\sf Aut}{\mathcal{M}}$ be a solvable arc-transitive automorphism group of $\mathcal{M}$}.
-By Lemmas 2.1 and 2.3, the group $G$ satisfies Hypothesis 2.2, so that
-G=(A:B):KG=(A{:}B){:}K
-as described in Theorem 1.1.
-Recall that $\mathcal{M}$} as a $G$-arc-transitive map satisfies one of the following properties: (i) $G$-regular; (ii) $G$-vertex-reversing; (iii) $G$-vertex-rotary; respectively in each case, the group $G$ has: (i) a regular triple; (ii) a reversing triple; (iii) a rotary pair.
-In particular, $G$ is of even order, and so $K\neq 1$.
-First, suppose $K=G_{2}$.
-By Lemma 6.1, either $K$ has a regular triple/reversing triple/rotary pair as $G$ does, and then $K$ is one of the 2-groups given in Proposition 4.2; or $K$ is ${\sf Z}_{2}$, dihedral ${\rm D}_{2^{\ell}}$ or cyclic ${\sf Z}_{2^{\ell}}$ with $\ell\geqslant 2$, while ${\rm D}_{2^{\ell}}$ and ${\sf Z}_{2^{\ell}}$ already appear in Proposition 4.2.
-Then, suppose $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}G_{7}$ or $K=G_{\{2,3,7\}}={\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3})$.
-In this case, $G$ has an odd factor group $G_{7}$ or $G_{7}{:}G_{3}$, respectively.
-By Lemma 6.1, G has neither regular triples nor reversing triples, so $G$ has a rotary pair, with each odd factor group cyclic.
-It follows that $K=G_{\{2,7\}}={\sf Z}_{2}^{3}{:}{\sf Z}_{7^{\ell}}$.
-At last, suppose that $K=G_{\{2,3\}}$ is one of the groups given in 5.6 (2) - (3).
-(a) If $\mathcal{M}$} is $G$-vertex-reversing, then $G$ has a reversing triple, with each factor group of even order by Lemma 6.1.
-Hence, $K$ is one of the groups in Tables 1 - 2, excluding those corresponding to cases (1.3), (1.4), (1.7), (2.1) - (2.3) as they have an odd factor group ${\sf Z}_{3}$ or ${\sf Z}_{3^{\ell}}$.
-(b) If $\mathcal{M}$} is $G$-regular, then $G$ has a regular triple.
-In particular, $G$ has a reversing triple, and so $K$ is one of the groups in (a).
-Furthermore, by Lemma 6.2 the groups which have ${\rm Q}_{8}{:}{\rm S}_{3}$ or ${\sf Z}_{4}\circ({\rm Q}_{8}{:}{\rm S}_{3})$ as a factor group should be excluded.
-It follows that $K$ is one of: $X$, $X\times{\sf Z}_{2}$, ${\rm D}_{2.3^{\ell}}\times{\rm S}_{4}$ or $({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6}$, where
-$X\in\{{\rm S}_{4},\,{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell+1}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\}.$
-(c) If $\mathcal{M}$} is $G$-vertex-rotary, then $G$ has a rotary pair, with each odd factor group cyclic by Lemma 6.1.
-So, either $K=X{:}K_{3}=X{:}{\sf Z}_{3^{\ell}}$ with $X\in\{{\sf Z}_{2}^{2},\,{\sf Z}_{2}^{3},\,{\rm Q}_{8},\,{\sf Z}_{4}\circ{\rm Q}_{8}\};$ or $K$ is one of the groups in Tables 1 - 2.
-For the former, $X\neq{\rm Q}_{8}$, as the only involution of ${\rm Q}_{8}{:}{\sf Z}_{3^{\ell}}$ lies in the center and so it does not have a rotary pair.
-For the latter, by Lemma 6.3 the groups corresponding to cases (1.5) and (1.6), which have $({\sf Z}_{3^{\ell}}\times{\sf Z}_{3}){:}{\sf Z}_{2}$ as a factor group, should be excluded.
-In conclusion, $K$ is one of the groups stated in Theorem 1.2. $\Box$
+By Lemmas 2.1 and 2.3, the group $G$ satisfies Hypothesis 2.2, so by Theorem 1.1, we have :
+G=H:K=(A:B):K,G=H{:}K=(A{:}B){:}K,
+where $A$ is abelian, $B$ is nilpotent and $A\cap{\bf Z}(A{:}B)=1$. We thus only need to present $K$.
+Recall that $\mathcal{M}$} as a $G$-arc-transitive map satisfies one of the following properties : $G$-regular ; $G$-vertex-reversing ; $G$-vertex-rotary ; respectively, in each case, the group $G$ has : a regular triple ; a reversing triple ; a rotary pair.
+In particular, $G$ is of even order, so $K\neq 1$.
+(a) Suppose $K=G_{2}$.
+By Lemma 6.1, either $K$ has a regular triple/reversing triple/rotary pair as $G$ does, and then $K$ is one of the 2-groups given in Proposition 4.2; or $K$ is ${\sf Z}_{2}$, dihedral ${\rm D}_{2^{\ell}}$ or cyclic ${\sf Z}_{2^{\ell}}$ ($\ell\geqslant 2$), while ${\rm D}_{2^{\ell}}$ and ${\sf Z}_{2^{\ell}}$ already appear in Proposition 4.2.
+This give the Theorem 1.2 (1).
+(b) Suppose that $K$ is one of the groups given in Theorem 1.1 (ii) and (iv) :
+${\sf Z}_{2}^{2}{:}G_{3},\ {\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}G_{3}),\ {\rm Q}_{8}{:}G_{3},\ {\sf Z}_{4}\circ_{{\sf Z}_{2}}({\rm Q}_{8}{:}G_{3}),\ {\sf Z}_{2}^{3}{:}G_{7},\ {\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3}).$
+Then, $G$ has an odd factor group $G_{3}$, $G_{7}$ or $G_{7}{:}G_{3}$.
+By Lemma 6.1, $G$ has neither regular triples nor reversing triples, so $G$ has a rotary pair, with each odd factor group cyclic, namely, $G_{3}={\sf Z}_{3^{\ell}}$, $G_{7}={\sf Z}_{7^{\ell}}$ and $K\neq{\sf Z}_{2}^{3}{:}(G_{7}{:}G_{3})$.
+Thus, $K$ is one of :
+${\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}}$,  ${\sf Z}_{2}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})$,  ${\sf Z}_{4}\circ_{{\sf Z}_{2}}({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})$,  ${\sf Z}_{2}^{3}{:}{\sf Z}_{7^{\ell}},$
+where ${\rm Q}_{8}{:}{\sf Z}_{3^{\ell}}$ is excluded, because it has the only involution lying in the center, and consequently, it does not have a rotary pair.
+This gives the first half of Theorem 1.2 (4).
+(c) At last, suppose that $K$ is one of the groups given in Theorem 1.1 (iii) : $X$, ${\sf Z}_{2}\times X$, $Y$, ${\sf Z}_{4}\circ_{{\sf Z}_{2}}Y$ and $W$, where $X,Y,W$ are given in Table 1.
+If $\mathcal{M}$} is $G$-vertex-reversing, then $G$ has a reversing triple, with each factor group of even order by Lemma 6.1.
+Note that, the groups $X,Y$ located in rows $2,3,6$ of Table 1 have an odd factor group ${\sf Z}_{3}$ or ${\sf Z}_{3^{\ell}}$, so does the groups $W$ located in rows $1,2,5$.
+Thus, $X,Y,W$ can only taken from the others, namely,
+$X\in\{{\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}},\,({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2},\,(({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})\times{\sf Z}_{3}){:}{\sf Z}_{2}\},$
+$Y\in\{{\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}},\,({\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\sf Z}_{3})){:}{\sf Z}_{2},\,({\sf Z}_{3}\times({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}\}$, and
+$W\in\{{\rm D}_{2.3^{\ell}}\times{\rm S}_{4},\,({\sf Z}_{2}^{2}{:}{\rm D}_{2.3^{\ell}})\times{\rm D}_{6},\,({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\rm S}_{3}),\,({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})\}$.
+If $\mathcal{M}$} is $G$-regular, then $G$ has a regular triple, so it has a reversing triple, and $K$ is one of the above groups.
+Furthermore, by Lemma 6.2, the groups which have ${\rm GL}_{2}(3)$ or ${\sf Z}_{4}\circ{\rm GL}_{2}(3)$ as a factor group should be excluded.
+Hence, $G\neq Y$, ${\sf Z}_{4}\circ_{{\sf Z}_{2}}Y$, $({\sf Z}_{3^{\ell}}{:}{\sf Z}_{4})\circ({\rm Q}_{8}{:}{\rm S}_{3})$ or $({\rm Q}_{8}{:}{\rm D}_{2.3^{\ell}})\circ({\sf Z}_{3}{:}{\sf Z}_{4})$.
+This gives Theorem 1.2 (2) and (3).
+(c.3) If $\mathcal{M}$} is $G$-vertex-rotary, then $G$ has a rotary pair, so by Lemma 6.3, the following four groups, which have $({\sf Z}_{3^{\ell}}\times{\sf Z}_{3}){:}{\sf Z}_{2}$ as a factor group, should be excluded :
+$({\sf Z}_{3^{\ell}}\times{\rm A}_{4}){:}{\sf Z}_{2}$,  $({\sf Z}_{3}\times({\sf Z}_{2}^{2}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$,  $({\sf Z}_{3^{\ell}}\times({\rm Q}_{8}{:}{\sf Z}_{3})){:}{\sf Z}_{2}$,  $({\sf Z}_{3}\times({\rm Q}_{8}{:}{\sf Z}_{3^{\ell}})){:}{\sf Z}_{2}$.
+Finally, this completes Theorem 1.2 (4). $\Box$
 Declarations